### v1

# Algorithms for Optimizing Acyclic Queries

Abstract
Most research on query optimization has centered on binary join algorithms like hash join and sort-merge join. However, recent years have seen growing interest in theoretically optimal algorithms, notably Yannakakis’ algorithm. These algorithms rely on join trees, which differ from the operator trees for binary joins and require new optimization techniques. We propose three approaches to constructing join trees for acyclic queries. First, we give an algorithm to enumerate all join trees of an $\alpha$-acyclic query by edits with amortized constant delay, which forms the basis of a cost-based optimizer for acyclic joins. Second, we show that the Maximum Cardinality Search algorithm by Tarjan and Yannakakis constructs a unique shallowest join tree, rooted at any relation, for a Berge-acyclic query; this tree enables parallel execution of large join queries. Finally, we prove that any connected left-deep linear plan for a $\gamma$-acyclic query can be converted into a join tree by a simple algorithm, allowing reuse of optimization infrastructure developed for binary joins.
1 Introduction
The query optimizer sits at the heart of a database system. It takes a query as input and generates a plan for efficient execution, allowing users to program declaratively without worrying about evaluation. Among the many relational algebra operators, join has received significant attention in optimization research. Its compositional nature allows for combining information from multiple relations, constructing complex queries from simple ones, and producing an output asymptotically larger than the inputs. The primary challenge is the join ordering problem to find the best arrangement of many join operations. Most existing research has focused on binary join algorithms such as hash join and sort-merge join, but these can produce unnecessarily large intermediates. Recent work has revived interest in optimal join algorithms, notably Yannakakis’ instance-optimal algorithm [44] for acyclic queries, which runs in linear time in the input and output size, $\mathcal{O}(|\textsf{IN}|+|\textsf{OUT}|)$. Its execution is guided by join trees whose nodes are relations, different from traditional binary join plans with relations at the leaves and join operators at the internal nodes. Although the algorithm is optimal regardless of the join tree, the choice of plan can affect practical performance. In this paper, we study the optimization problem in the context of Yannakakis-style algorithms.
A query optimizer typically has two parts: a plan generator and a cost model to assess each plan. This paper focuses on plan generation and presents three approaches:
  • An algorithm that enumerates all join trees of an acyclic query by edits, with amortized constant time delay.
  • A proof that Maximum Cardinality Search [40] yields a unique shallowest join tree for any Berge-acyclic query, enabling parallel execution for very large queries.
  • A characterization of when the algorithm by Hu et al. [20] can convert a binary join plan into a valid join tree, allowing reuse of existing optimizers.
The rest of the paper is organized as follows: section˜2 discusses related work; section˜3 introduces relevant concepts and notations; section˜4 presents the join tree enumeration algorithm; section˜5 introduces canonical join trees and their construction; section˜6 discusses the conversion from binary plans to join trees; finally, section˜7 points to avenues for future work and concludes the paper. All missing proofs can be found in the appendix, which also contains additional figures, tables and algorithms.
2 Related Work
Join order optimization is well studied, with algorithms based on dynamic programming (DP) from the bottom up [28, 34, 30], cost-based pruning from the top down [11, 15], greedy heuristics [5, 14, 38], and randomized search [36]. Since the plan space is exponential, most methods prune it: some restrict to left-deep plans [21, 23], while others avoid Cartesian products [28, 29]. Our algorithms restrict the query plans to those running in linear time for acyclic queries. In particular, while avoiding Cartesian products requires each subplan to form a spanning tree of the corresponding subquery’s join graph, [^1] The join graph of a query has a vertex for each relation and an edge for each pair of relations that join with each other. We later define this as the line graph of the query hypergraph in definition 2. our algorithms find maximum spanning trees of the weighted join graph. Several algorithms for ordering binary joins are based on dynamic programming and tabulate shared structures among different plans [28, 29]. This is desirable because subplans are grouped into equivalence classes, and the optimal plan can be constructed in a bottom-up manner. Our enumeration algorithm avoids redundant work by enumerating join trees by edits, i.e., it outputs the difference between consecutive join trees instead of the trees themselves. Future work may explore constructing compact representations of join trees, suitable for dynamic programming.
Several recent papers have proposed practical implementations of Yannakakis’ algorithm for acyclic queries [2, 20, 42, 47]. For example, Zhao et al. [47] find that different query plans perform similarly, thanks to the optimality of Yannakakis’ algorithm. They adopt a simple heuristic to construct the join tree by picking the largest input relation as the root, and then greedily attaching the remaining relations into the tree. Inspired by this algorithm, we prove in section˜5 that for Berge-acyclic queries there is a unique shallowest join tree for any given root where the depth of each node is minimized. Furthermore, this tree can be constructed in linear time by Tarjan and Yannakakis’ Maximum Cardinality Search algorithm [40]. Shallow trees are desirable for parallel execution, where the depth of the tree determines the number of sequential steps. Other practical implementations of Yannakakis’ algorithm leverage existing optimizers for binary joins and convert a binary plan into a join tree [2, 20]. In particular, Hu et al. [20] find that every left-deep linear plan encountered in practice can be converted into a join tree by a simple algorithm. This is not surprising, as we will prove in section˜6 that every connected left-deep linear plan of a $\gamma$-acyclic query must traverse some join tree from root to leaves.
On the theoretical side, attention has been focused on finding (hyper-)tree decompositions to improve the asymptotic complexity of query processing [37, 19, 16]. The general goal is to find a decomposition with small width which can be used to guide the execution of join algorithms. Most algorithms find a single decomposition with minimum width to achieve the optimal asymptotic complexity [37, 19, 16]. Nevertheless, different decompositions with the same width may still lead to different performance in practice, and cost-based optimization remains crucial. For this, Carmeli et al. [6] propose an algorithm to enumerate tree decompositions with polynomial delay. In this paper, we focus on acyclic queries and their join trees, which are precisely the decompositions with width 1. Our enumeration algorithm can generate all join trees by edits with amortized constant delay.
3 Preliminaries
We focus on full conjunctive queries [1] in this paper and identify each query with its hypergraph, where there is a vertex for each variable and a hyperedge for each relation.
Definition 1.
A hypergraph is a tuple $(X,R,\chi)$ where $X$ is a set of vertices, $R$ a set of hyperedges, and $\chi:R\to 2^{X}$ is a function that assigns to each hyperedge a set of vertices.
Without loss of generality, we limit the scope of our discussion to queries which admit connected hypergraphs. For queries with disconnected hypergraphs, we treat each component separately. We do not consider isolated vertices, empty hyperedges or duplicated hyperedges containing the same set of vertices. Therefore, each hyperedge can be identified with its vertex set. We will use $x\in r$ interchangeably with $x\in\chi(r)$ and apply the common set operations to hyperedges. The size of a hypergraph is the sum of all hyperedges $|H|=\sum_{r\in R}|r|$. We let $H|_{x}$ denote the neighborhood of $x$ in $H$, consisting of all hyperedges containing vertex $x$.
Input: Hypergraph $H$, root $r$
Output: Join tree $T_{r}$ rooted at $r$
1
[…]
4while $R\neq\emptyset$ do
5 $i\leftarrow i+1$
6 $r_{i}\leftarrow r$
7 for $x\in r_{i}\cap X$ do
8    for $r^{\prime}\in R:x\in r^{\prime}$ do
9       $p(r^{\prime})\leftarrow r_{i}$
10
11
12 $X\leftarrow X\setminus r_{i}$
13 $R\leftarrow R\setminus\{r_{i}\}$
14 $r\leftarrow\operatorname*{arg\,max}_{r\in R}|r\setminus X|$
Algorithm 1 MCS from $r\in R(H)$
(a) Hypergraph $H_{6}$
(b) Line graph $L_{6}$
(c) $T^{G}$ by GYO
(d) $T^{M}$ by MCS
Figure 1: A hypergraph $H_{6}$, its line graph $L_{6}$, and two join trees $T^{G}$ and $T^{M}$.
Definition 2.
The line graph $L(H)$ of a hypergraph $H=(X,R,\chi)$ is an undirected simple graph $(R,E)$ whose vertices $R$ are the hyperedges of $H$, and there is an edge $(r_{1},r_{2})\in E$ whenever $r_{1}\cap r_{2}\neq\emptyset$. We extend $\chi$ such that $\chi(e)=\chi(r_{1},r_{2}):=r_{1}\cap r_{2}$ for each edge $e=(r_{1},r_{2})\in E$, and define the edge-weight function $\omega(r_{1},r_{2}):=|r_{1}\cap r_{2}|$.
When there is no ambiguity from the context, we simply write $L(H)$ as $L$. The size of a line graph is the sum of all edge weights $|L|=\sum_{e\in E(L)}\omega(e)$. It reflects the number of join clauses in a SQL query with one clause per variable shared between two relations. We let $L|_{x}$ denote the subgraph of $L(H)$ induced by $H|_{x}$.
Example 3.
fig.˜2(a) shows an example hypergraph of size 14. The line graph $L_{6}=L(H_{6})$ is shown in fig.˜2(b). For instance, hyperedges $Y$ and $U$ share two vertices $a$ and $d$, so they are connected by an edge in the line graph. The edge weight is given by $\omega(Y,U)=|\{a,d\}|=2$. Each unannotated edge in $L_{6}$ connects two hyperedges sharing only the vertex $a$. Otherwise, the common vertices shared by a pair of hyperedges are annotated next to the corresponding edge. The size of the line graph is $|L_{6}|=19$.
Definition 4.
A join tree $T$ of hypergraph $H$ is a spanning tree of $L(H)$ such that $T|_{x}$ is a connected subtree for each $x\in X(H)$. If a certain vertex is specified as the root, $T$ becomes a rooted join tree where the edges are oriented away from the root.
A join tree $T(L)$ is a subgraph of $L$, so we use $R(T)$ to denote the set of tree nodes and $E(T)$ to denote the set of tree edges. Because $T$ is spanning and acyclic, we always have $R(T)\equiv R(L)$ and $E(T)\subseteq E(L)$. The requirement that $T|_{x\in X(H)}$ is connected is also known as the running intersection property [13]. We write $\mathcal{T}(L(H))$ or $\mathcal{T}(H)$ to denote the set of unrooted join trees. The union join graph $\mathcal{U}(L(H))$ or $\mathcal{U}(H)$ is a spanning subgraph of $L(H)$ where the edge set is the union of all join trees $E(\mathcal{U}(H))=\bigcup_{T\in\mathcal{T}(H)}E(T)$. We use $T_{r}$ to denote a join tree rooted at $r\in R(T)$. When there is no ambiguity, we simply write $T$. The depth of a node $r_{i}$ in the rooted tree $T_{r}$, denoted $0pt(T_{r},r_{i})$, is defined as its distance from the root. Join trees can be constructed by a procedure called GYO reduction.
Definition 5.
A GYO reduction order is a sequence of hyperedges $r_{1},r_{2},\ldots,r_{k}$ such that for each $r_{i<k}$, there is some $r_{p>i}$, called the parent of $r_{i}$, such that $\forall r_{j>i}:r_{i}\cap r_{j}\subseteq r_{p}$.
The GYO reduction algorithm finds such an order iteratively, and attaches each hyperedge to its parent to form a join tree. It generates the join tree $T^{G}$ as shown in fig.˜2(c). Readers may refer to the work of Yu and Ozsoyoglu [45] for details.
Four common notions of hypergraph acyclicity are defined in decreasing order of strictness, following the seminal work by Fagin [13], namely Berge-acyclic $\Rightarrow$ $\gamma$-acyclic $\Rightarrow$ $\beta$-acyclic $\Rightarrow$ $\alpha$-acyclic. The detailed definitions are given by definition˜41 in appendix˜A.
Another algorithm for checking acyclicity of hypergraphs and constructing join trees is the Maximum Cardinality Search (MCS) algorithm by Tarjan and Yannakakis [40]. algorithm˜1 shows a simplified pseudocode of MCS. We discuss the properties of MCS and how it constructs the join trees $T^{M}$ in fig.˜2(d) with example˜42 in appendix˜A.
Computation model.
Throughout the paper we assume the Random Access Machine model of computation, where one can allocate an array of size $n$ in $\mathcal{O}(n)$ time. Constant-time operations include accessing and updating an array element, adding or deleting an element in a linked list, and the common arithmetic operations on integers.
4 Enumerating Join Trees
Our strategy for enumerating join trees is based on the following classic result relating join trees of a hypergraph to maximum spanning trees (MSTs) of its line graph:
Proposition 6 (Maier [27]).
A subgraph of $L(H)$ is a join tree of $H$ if and only if it is a maximum spanning tree of $L(H)$ according to the weight function $\omega$.
A naïve approach is to apply an off-the-shelf algorithm for enumerating the maximum spanning trees of $L(H)$. The best known algorithm for this purpose is due to Eppstein [12], and it works by deriving from the input graph $G$ a so-called equivalent graph $G^{\equiv}$. Every spanning tree of $G^{\equiv}$ corresponds to an MST of $G$ and vice versa. Eppstein proves a lower bound of $\Omega(m+n\log n)$ on constructing the equivalent graph from an arbitrary weighted graph with $m$ edges and $n$ vertices. Then to enumerate all $k$ MSTs of $G$, Eppstein applies existing algorithms to enumerate the spanning trees of $G^{\equiv}$. Since there are optimal spanning tree enumeration algorithms that run in $\mathcal{O}(m+n+k)$ [22, 35], the overall time complexity to enumerate MSTs is $\mathcal{O}(m+n\log n+k)$.
The main result of this section is an algorithm for enumerating join trees leveraging the structure of acyclic hypergraphs and their line graphs. In particular:
  • for any $\alpha$-acyclic hypergraph $H$, we can construct an equivalent graph in $\mathcal{O}(|L|)=\mathcal{O}(m+n)$, where $m$ and $n$ are the numbers of edges and vertices in $L(H)$, thus enumerating the join trees in $\mathcal{O}(|L|+k)$.
  • for any $\gamma$-acyclic hypergraph $H$, we can construct an equivalent graph from the union join graph $\mathcal{U}(H)$ which can be built in $\mathcal{O}(|H|+|\mathcal{U}(H)|)$, lowering the overall time complexity of enumeration to $\mathcal{O}(|H|+k)$.
In the rest of this section, we formally introduce the notion of the equivalent graph $G^{\equiv}$. Thereafter, we outline the main algorithm for enumerating join trees, prove its correctness and analyze its complexity. Finally, we consider the case of $\gamma$-acyclic queries.
4.1 Equivalent Graph
The construction of the equivalent graph $G^{\equiv}$ for a given edge-weighted graph $G$, is based on the following local graph modification, sliding transformation.
Definition 7 (Sliding Transformation [12]).
Let $G$ be an edge-weighted graph with weight function $\omega$. Given edges $e^{*}=(u,v)$ and $e=(v,w)$ where $\omega(e)<\omega(e^{*})$, sliding $e$ along $e^{*}$ moves $e$’s end node $v$ to $u$ such that $e=(u,w)$.
(a) $T_{W}$
(b) 3 Edges Slid
(c) $G^{\equiv}$
(d) $T^{\prime}_{W}$
Figure 2: $T_{W}$ is an MST of the line graph in fig.˜2(b) where the black solid lines stand for the tree edges, the dashed line for an MST edge $(S,U)$, the dotted line for a non-MST edge $(T,Y)$ and gray solid lines for the remaining non-tree edges. $G^{\equiv}$ is the resulting equivalent graph where the thick lines highlight the parallel edges.
We can construct an equivalent graph by repeatedly performing sliding transformations. Sliding may create parallel edges, cycles, and self-loops, so the result is a multigraph.
Definition 8 (Equivalent Graph [12]).
Given weighted graph $G$ with MST $T$ rooted, an equivalent graph $G^{\equiv}$ is obtained from $G$ by exhaustively sliding every edge $e=(v,w)\in E(G)$ along another $e^{*}=(u,v)\in T$ as long as $0pt(T,u)<0pt(T,v)$.
Sliding transformation preserves the identity of each edge $e\in E(G)$ as explained in example˜10. Every spanning tree of $G^{\equiv}$ is an MST of $G$ and vice versa.
Theorem 9 (by Eppstein [12]).
The maximum spanning trees of a weighted graph $G$ are in one-to-one correspondence with the spanning trees of its equivalent graph $G^{\equiv}$. [^2] Minimum spanning trees are considered in the original literature [12].
In the following, we use $p(r)$ to denote the parent of $r$ in a given rooted tree, and $c(r)$ to denote its children. We extend the notation of parent to edges as follows: for any edge $e\in E(T)$, its parent $p(e)\in E(T)$ is another tree edge incident to $e$ and closer to the root $r$. The edges incident to the root do not have a parent and are all siblings. Otherwise, the siblings of $e$ are incident to the same parent, denoted as $s(e)=\left\{e^{\prime}\in E(T)\setminus\{e\}\mid p(e^{\prime})=p(e)\right\}$.
Example 10.
Given the 6-clique line graph $L_{6}$ in fig.˜2(b), we find a rooted MST $T_{W}$ as shown in fig.˜2(a). Among all the tree edges shown as black solid lines, we can only apply the sliding transformation to the edge $(T,P)$ whose weight $\omega(T,P)=|\{a\}|=1$ is lighter than its parent tree edge $(W,T)$ with $\omega(W,T)=|\{a,c\}|=2$. We slide along the tree edge $(W,T)$ to the root so that the edge $(T,P)$ becomes $(W,P)$ as shown by the solid curve in fig.˜2(b). Non-tree edges can slide similarly. We consider two examples, $(S,U)$ illustrated with a dashed line and $(Y,T)$ with a dotted line in fig.˜2(a). All other non-tree edges are shown in light gray. We can slide along tree edge $(U,W)$ to the root so that the edge $(S,U)$ becomes $(S,W)$ as shown by the dashed curve in fig.˜2(b). Both ends of $(Y,T)$ can slide along the tree edges to the root so that the edge becomes a self-loop as shown by the dotted loop in fig.˜2(b). This edge will not appear in any spanning tree of $G^{\equiv}$, and therefore not a part of any MST of $L_{6}$. We refer to such an edge as a non-MST edge, as opposed to an MST edge. By applying sliding transformations to a fixpoint, we obtain an equivalent graph $G^{\equiv}$, where there are two sets of parallel edges highlighted by thick lines in fig.˜2(c). For example, the tree edge $(T,P)$ in $T_{W}$ and non-tree edges $(P,U),(P,T),(P,Y)$ are parallel in $G^{\equiv}$ between $P$ and $W$. We can easily verify that each spanning tree of $G^{\equiv}$ corresponds to an MST of $L_{6}$, such as $T^{\prime}_{W}$ in fig.˜2(d).
The choice of the initial spanning tree can affect the structure of the equivalent graph, but the order of sliding transformations performed has no impact [12].
4.2 Enumerating Join Trees of $\alpha$-Acyclic Hypergraphs
The bottleneck of Eppstein’s algorithm for constructing equivalent graphs of general graphs comes from a subroutine that identifies where each edge will eventually slide to. Because each edge can only slide along a heavier edge, it will eventually be “blocked” by a lighter or equally weighted edge along its path to the root. The subroutine essentially performs binary search to find the blocking edge, leading to the $\log n$ factor in the overall complexity. The key to our improvement is to show that for every acyclic hypergraph $H$, we can construct an “equivalent hypergraph” $H^{*}$ that shares the same structure with $H$ and has the same set of join trees, but $H^{*}$ admits a special join tree with monotonically increasing weight from root to leaf, which enables constant-time identification of the blocking edge.
Definition 11.
A monotonic weight join tree $\hat{T}\in\mathcal{T}(H)$ is a rooted join tree of $H$ such that for any $e\in\hat{T}$ that has a parent $p(e)$, $\omega(e)>\omega(p(e))$.
Example 12.
$H^{*}_{6}$ in fig.˜4(a) is a hypergraph with similar structure to $H_{6}$ in fig.˜2(a). $H^{*}_{6}$ differs from $H_{6}$ by one vertex $d^{\prime}$ and admits the same set of join trees. Therefore, finding an equivalent graph $G^{\equiv}(H^{*}_{6})$ is sufficient for enumerating the join trees of $H_{6}$. $H^{*}_{6}$ also admits a monotonic weight join tree $\hat{T}_{P}$ in fig.˜4(b). This means each tree edge in $\hat{T}_{P}$ is already in place, because it cannot slide along the lighter parent. For each non-tree edge, we can identify its destination in constant time using algorithms for finding lowest common ancestors [3] and level ancestors [4], as detailed in algorithm˜3.
The example shows we can enumerate join trees of $H$ by constructing $H^{*}$ and its equivalent graph $G^{\equiv}(H^{*})$, then enumerate spanning trees of $G^{\equiv}(H^{*})$. However, constructing $H^{*}$ has a cost that may exceed our target time complexity. We show that, perhaps surprisingly, running algorithm˜3 directly on $H$ produces the same equivalent graph $G^{\equiv}(H^{*})$, so we can skip the costly construction of $H^{*}$.
Before presenting the algorithm, we first formalize the relationship between $H$ and $H^{*}$ using the concept of hypergraph homomorphisms. Note the definition of homomorphisms is not standardized in the literature, and we choose ours carefully to simplify the proofs. In particular, our definition is equivalent to the homomorphism of incidence graphs [32].
Definition 13 (Hypergraph Homomorphism).
For two hypergraphs $H_{1}$, $H_{2}$, a homomorphism $\mathsf{f}$ is a pair of functions $(\mathsf{f}_{X}:X(H_{1})\rightarrow X(H_{2}),\mathsf{f}_{R}:R(H_{1})\rightarrow R(H_{2}))$ such that:
$\forall r_{1}\in R(H_{1}),x_{1}\in r_{1}:\mathsf{f}_{X}(x_{1})\in\mathsf{f}_{R}(r_{1})$
We write $H_{1}\rightarrow H_{2}$ to denote that there is a homomorphism from $H_{1}$ to $H_{2}$.
When there is no ambiguity from the context, we simply write $\mathsf{f}$ for $\mathsf{f}_{X}$ or $\mathsf{f}_{R}$. Intuitively, homomorphisms preserve relations. A strong homomorphism also reflects relations:
Definition 14 (Strong Homomorphism).
A homomorphism $\mathsf{f}:H_{1}\rightarrow H_{2}$ is called strong if it also satisfies:
$\forall r_{1}\in R(H_{1}),x_{1}^{\prime}\not\in r_{1}:\mathsf{f}_{X}(x_{1}^{\prime})\not\in\mathsf{f}_{R}(r_{1})$
We write $H_{1}\twoheadrightarrow H_{2}$ to denote that there is a strong homomorphism from $H_{1}$ to $H_{2}$.
In this paper we only consider hypergraphs over the same set of edges, [^3] The hypergraphs share the same hyperedge set $R$ but differ in the hyperedge-to-vertices functions $\chi$, i.e., an edge may have different sets of vertices in different hypergraphs. and from now on we assume $\mathsf{f}_{R}$ is the identity function.
We now show that if there is a strong homomorphism $\mathsf{f}:H^{\prime}\twoheadrightarrow H$, then every join tree of $H$ is also a join tree of $H^{\prime}$.
Lemma 15.
If $H^{\prime}\twoheadrightarrow H$ then $\mathcal{T}(H)\subseteq\mathcal{T}(H^{\prime})$.
We construct $H^{*}$ such that $H^{*}\twoheadrightarrow H$ and $H^{*}$ admits a monotonic weight join tree $\hat{T}$, by first using the MCS algorithm to construct a join tree $T$ of $H$ called the MCS tree. Then, we perturb the weights of $T$ so that they increase monotonically from the root to leaves. Finally, we construct $H^{*}$ such that the perturbed tree is a join tree of $H^{*}$.
The next result shows every MCS tree is already “somewhat monotonic”, in that every edge must contain some variable not in its parent. With abuse of notation we write $e_{1}\subseteq e_{2}$ for $\chi(e_{1})\subseteq\chi(e_{2})$ and use the common set operations directly on edges to reduce clutter.
Lemma 16.
Let $T$ be an MCS tree. For any edge $e\in T$ that has a parent, then $e\not\subseteq p(e)$.
Furthermore, if two edges share any variable not in their parents, they must be siblings.
Lemma 17.
Let $T$ be an MCS tree. For two edges $e,e^{\prime}\in T$ that have parents,
$(e\setminus p(e))\ \cap\ (e^{\prime}\setminus p(e^{\prime}))\ \neq\ \emptyset\quad\implies\quad p(e)=p(e^{\prime}).$
We now show how to derive from an $\alpha$-acyclic hypergraph $H$ an equivalent hypergraph $H^{\prime}$ that preserves $\mathcal{T}(H)$ and admits a monotonic weight join tree $\hat{T}$.
Definition 18 (Vertex Duplication).
Given a hypergraph $H$ and a vertex $x\in X(H)$, duplicating $x$ produces a new hypergraph $H^{\prime}$ that differs from $H$ by only one vertex $\{x^{\prime}\}=X(H^{\prime})\setminus X(H)$ which is added to each $r\in H|_{x}$.
We will refer to such a hypergraph $H^{\prime}$ as a vertex duplication of $H$ or simply a duplicated hypergraph of $H$. We generalize duplication to a set of vertices in the natural way. Importantly, vertex duplication preserves the set of join trees of the original hypergraph.
Lemma 19.
Let $H^{\prime}$ be a duplicated hypergraph of $H$. Then,
  1. 1.
    $H^{\prime}\twoheadrightarrow H$ and $H^{\prime}\twoheadleftarrow H$
  2. 2.
    $\mathcal{T}(H^{\prime})=\mathcal{T}(H)$
  3. 3.
    $L(H^{\prime})=L(H)$
We are now ready to construct $H^{*}$ by vertex duplication.
Input: MCS tree $T_{r}(H)$
1
2for $r^{\prime}\in c(r)$ do
3 $e\leftarrow(r,r^{\prime});\quad q$.enqueue($e$)
4
5while $q\neq\emptyset$ do
6 $e\leftarrow q.$dequeue()
7 $\Delta\leftarrow\omega(p(e))-\omega(e)+1$
8 if $\Delta>0$ then
9    $x\leftarrow\text{a vertex in }\chi(e)\setminus\chi(p(e))$
10     make $\Delta$ duplicates of $x$
11
12 for $e^{\prime}\in c(e)$ do
13    $q$.enqueue($e^{\prime}$)
14
15
Algorithm 2 Construction of $H^{*}$
(a) $H^{*}_{6}$
(b) $\hat{T}_{p}$
Figure 3: An equivalent hypergraph $H^{*}_{6}$ and its monotonic weight join tree $\hat{T}_{P}$
Theorem 20.
Given an MCS tree $T$ of an $\alpha$-acyclic hypergraph $H$, there exists a hypergraph $H^{*}$ that
  • admits a monotonic weight join tree $\hat{T}\in\mathcal{T}(H^{*})$ whose edges are in a one-to-one correspondence to those of $T$
  • satisfies $\mathcal{T}(H^{*})=\mathcal{T}(H)$.
Proof.
To prove the existence, we construct such a hypergraph $H^{*}$ by vertex duplication. lemma˜19 guarantees that $\mathcal{T}(H^{*})=\mathcal{T}(H)$. We only need to show the existence of $\hat{T}\in\mathcal{T}(H^{*})$. algorithm˜2 constructs $H^{*}$ by vertex duplication in a breadth-first manner from the root of $T$ to make it a monotonic weight join tree in $H^{*}$.
algorithm˜2 starts with enqueuing each edge incident to the root. Once a tree edge $e\in E(T_{r})$ is dequeued, we check for the number of duplications needed to make it heavier than its parent edge $p(e)$, namely $\Delta=\omega(p(e))-\omega(e)+1$. The duplication is performed only if $\Delta>0$. lemma˜16 guarantees that $\chi(e)\setminus\chi(p(e))\neq\emptyset$. Therefore, we can always duplicate a vertex $x\in\chi(e)\setminus\chi(p(e))$ for $\Delta$ times such that $\omega(e)>\omega(p(e))$.
Let $e=(r_{u},r_{d})$ where $0pt(T,r_{u})<0pt(T,r_{d})$, $x$ only occurs in the subtree rooted at $r_{u}$. Otherwise, it violates the running intersection property. Therefore, the duplication of $x$ does not affect the weight of any edge closer to the root than $e$ in $T$. By lemma˜17, duplication in a sibling edge $e^{\prime}\in s(e)$ would not affect the weight of $e$ either, and $\omega(e)>\omega(p(e))$ will continue to hold. The tree has monotonic weights when algorithm˜2 terminates. ∎
Given a rooted join tree $T$ of a hypergraph $H$, a non-tree edge is $e=(r_{i},r_{j})\in E(L(H))\setminus E(T)$. There is a path in $T$ between $r_{i}$ and $r_{j}$ via their lowest common ancestor $\mathsf{LCA}(r_{i},r_{j})$. We define the LCA edges $\lambda(e)=\lambda(r_{i},r_{j})$ as a set of at most two tree edges on the path and incident to $\mathsf{LCA}(r_{i},r_{j})$. If $r_{i}$, $r_{j}$ are ancestor and child, then $|\lambda(e)|=1$, otherwise, $|\lambda(e)|=2$.
In the process of building $G^{\equiv}(H^{*})$ from a monotonic weight join tree $\hat{T}$ of $H^{*}$, definition˜7 and definition˜11 guarantee that:
  • no tree edge in $\hat{T}$ needs to be slid;
  • each non-tree edge $e$ can be removed if it is lighter than all $e^{\prime}\in\lambda(e)$;
  • otherwise, the non-tree edge $e$ can be slid directly to be incident to each $e^{\prime}\in\lambda(e)$.
algorithm˜3 constructs $G^{\equiv}(H^{*})$ from the line graph $L(H)$ weighted by $\omega$ and an MCS tree $T_{r}$. Full details on the constant-time operations on graphs and data structures to support operations performed in algorithm˜3 can be found in appendix˜B. As a part of pre-processing, we first conduct a breadth-first search on $\hat{T}_{r}$ to obtain the depth table $0pt$ of each tree node in $\mathcal{O}(|R(\hat{T}_{r})|)$. We will also use the algorithms proposed by Michael Bender [3, 4] to build the data structures in $\mathcal{O}(|R(\hat{T}_{r})|)$ to facilitate the constant-time query of lowest common ancestors (LCAs) and level ancestors (LAs). The constant-time query of depth, LCA and LA allows for finding the $\lambda(e)$ of any non-tree edge $e$ in constant time as shown from Line 3 to Line 3 in algorithm˜3.
Input: MCS tree $T_{r}(H)$ of line graph $L(H)$ with weight function $\omega$
1 $G^{\equiv}\leftarrow L(H)$
2for $e^{*}=(r_{i},r_{j})\in R(L(H))\setminus R(T_{r}(H))$ do
 // where $0pt(T_{r}(H),r_{i})\leq 0pt(T_{r}(H),r_{j})$
3 $w^{*}=\omega(e^{*})$
4 $l\leftarrow\mathsf{LCA}(r_{i},r_{j});d\leftarrow 0pt(l)+1$
5 $e_{1}=(l,r_{1}\leftarrow\mathsf{LA}(r_{i},d));w_{1}=\omega(e_{1})$
6 $e_{2}=(l,r_{2}\leftarrow\mathsf{LA}(r_{j},d));w_{2}=\omega(e_{2})$
7 if $w^{*}<w_{1}\wedge w^{*}<w_{2}$ then
8      delete $e^{*}$ from $R(G^{\equiv})$
9
10 if $e_{1}=e_{2}\wedge w^{*}=w_{1}$ then
11      slide $r_{j}$ to $r_{1}$
12
13 else if $e_{1}\neq e_{2}\wedge w^{*}=w_{1}=w_{2}$ then
14      slide $r_{i}$ to $r_{1}$ and $r_{j}$ to $r_{2}$
15
16 else if $e_{1}\neq e_{2}\wedge w^{*}=w_{1}<w_{2}$ then
17      slide $r_{i}$ to $l$ and $r_{j}$ to $r_{1}$
18
19 else if $e_{1}\neq e_{2}\wedge w^{*}=w_{2}<w_{1}$ then
20      slide $r_{i}$ to $l$ and $r_{j}$ to $r_{2}$
21
22return $G^{\equiv}$
Algorithm 3 buildEGbuildEG
(a) Monotonic Weight Join Tree $\hat{T}_{P}(H^{*}_{6})$
(b) 3 Non-tree Edges Processed
(c) Equivalent Graph $G^{\equiv}(H^{*}_{6})$
Figure 4: algorithm˜3 on $H^{*}_{6}$ of fig.˜4(a)
Example 21.
We consider $\hat{T}_{P}$ in fig.˜4(b) of an equivalent hypergraph $H^{*}_{6}$ in fig.˜4(a). Three of $\hat{T}_{P}$’s non-tree edges are shown in fig.˜5(a) where the non-MST edges $(Y,W),(Y,T)$ are dotted and the MST edge $(Y,S)$ is dashed. We can identify $\lambda(Y,W)=\{(T,W),(T,U)\}$ in constant time. Because $\omega(Y,W)=|\{a\}|=1$ is lighter than $\omega(T,W)=\omega(T,U)=|\{a,c\}|=2$, it is removed. $(Y,T)$ is also removed, for it is lighter than its LCA edge $(T,U)$. We compare $(Y,S)$ with $\lambda(Y,S)=\{(P,T),(P,S)\}$. Because they have the same weight with $\omega(Y,S)=\omega(P,T)=\omega(P,S)=|\{a\}|=1$, we slide $(Y,S)$ directly to $(T,S)$. The intermediate graph with those three non-tree edges processed is shown in fig.˜5(b). The resulting equivalent graph $G^{\equiv}(H^{*}_{6})$ is shown in fig.˜5(c).
The next result follows immediately from the construction of the equivalent hypergraph.
Lemma 22.
Let $H^{*}=(X^{*},R,\chi^{*})$ be an equivalent hypergraph derived from $H=(X,R,\chi)$ by vertex duplication. For any non-tree edge $e$ of an MCS tree $T$ of $H$,
  • If $\chi(e)=\chi(e^{\prime})$ for some $e^{\prime}\in\lambda(e)$, then $\chi^{*}(e)=\chi^{*}(e^{\prime})$
  • If $\chi(e)\subset\chi(e^{\prime})$ for some $e^{\prime}\in\lambda(e)$, then $\chi^{*}(e)\subset\chi^{*}(e^{\prime})$
Proof.
The duplication of a vertex $x\in\chi(e)$ adds $x^{\prime}$ to both $\chi(e)$ and $\chi(e^{\prime})$. If $\chi(e)=\chi(e^{\prime})$, $\chi^{*}(e)=\chi(e)\cup\{x^{\prime}\}=\chi(e^{\prime})\cup\{x^{\prime}\}=\chi^{*}(e^{\prime})$. A similar argument applies if $\chi(e)\subset\chi(e^{\prime})$. ∎
lemma˜22 allows us to produce $G^{\equiv}(H^{*})$ by keeping all tree edges of the MCS tree $T(H)$ in place and sliding the non-tree edges directly to their LCA edges. For example, running algorithm˜1 on the MCS tree $T(H_{6})$ in fig.˜2(d) produces an identical $G^{\equiv}(H^{*}_{6})$ in fig.˜5(c).
Theorem 23.
Let $H$ be an $\alpha$-acyclic hypergraph with weight function $\omega$, line graph $L$ and MCS tree $T$. Let $H^{*}$ be the equivalent hypergraph of $H$ with weight function $\omega^{*}$. Applying algorithm˜3 to inputs $L$, $T$ and $\omega$ produces the same equivalent graph as applying algorithm˜3 to inputs $L$, $T$ and $\omega^{*}$.
Proof.
One can readily check that for any non-tree edge $e$ of $T$, the conditions of each if statement in algorithm˜3 hold under $\omega$ if and only if they hold under $\omega^{*}$ by lemma˜22. ∎
Theorem 24.
Given line graph $L$ of $\alpha$-acyclic hypergraph $H$ with weight function $\omega$ and an MCS tree $T$ of $H$, algorithm˜3 returns equivalent graph $G^{\equiv}$ of equivalent hypergraph $H^{*}$ in $\mathcal{O}(m+n)$ where $n$ is the number of vertices and $m$ is the number of edges in $L$.
Proof.
The construction of each data structure $0pt(\cdot)$, $\mathsf{LA}(\cdot)$ and $\mathsf{LCA}(\cdot)$ can be done in $\mathcal{O}(m+n)$ [3, 4]. Execution of the rest of the algorithm also takes time $\mathcal{O}(m+n)$, as it requires examining each non-tree edge to either delete it or slide it directly to its LCA edges in the MCS tree. For each non-tree edge, the lookup of at most two LCA edges takes time $\mathcal{O}(1)$, as does the comparison of weights and the direct sliding transformation. ∎
Once we have the equivalent graph $G^{\equiv}$, the existing algorithm allows for enumerating all its spanning trees with an amortized constant delay. The total time complexity of enumerating $|\mathcal{T}(H)|$ given the line graph $L(H)$ as input is $\mathcal{O}(|L(H)|+|\mathcal{T}(H)|)$.
Starting from an $\alpha$-acyclic query $\mathcal{Q}$, section˜C.1 describes how to convert it into a hypergraph $H$. section˜C.2 describes how to convert it into a weighted line graph $L(H)$. Both conversions take linear time $\mathcal{O}(|\mathcal{Q}|)$ with respect to the query size as defined in appendix˜C. algorithm˜1 constructs an MCS tree $T$ in $\mathcal{O}(|H|)\leq\mathcal{O}(|\mathcal{Q}|)$. With the inputs $L(H)$ and $T$, we can run algorithm˜3 to construct $G^{\equiv}(H^{*})$ in $\mathcal{O}(m+n)\leq\mathcal{O}(|\mathcal{Q}|)$ (where $n$ is the number of vertices and $m$ is the number of edges in $L(H)$). Finally, we apply the spanning tree enumeration algorithm by Kapoor and Ramesh [22] on $G^{\equiv}(H^{*})$ to enumerate all join trees $|\mathcal{T}(H)|$ in $\mathcal{O}(m+n+|\mathcal{T}(H)|)\leq\mathcal{O}(|\mathcal{Q}|+|\mathcal{T}(H)|)=\mathcal{O}(|L(H)|+|\mathcal{T}(H)|)$.
4.3 $\gamma$-Acyclic Queries
The run time of algorithm˜3 depends on the size of the line graph which can be quadratically larger than the input hypergraph. But if $H$ is $\gamma$-acyclic, we can bring the total time complexity of enumeration down to $\mathcal{O}(|H|+|\mathcal{T}(H)|)$. The line graph of a $\gamma$-acyclic hypergraph $H$ is its union join graph [25] $L(H)=\mathcal{U}(H)$, i.e., every edge $e\in E(L(H))$ in the line graph is an MST edge. By modifying the MCS algorithm to algorithm˜7 in appendix˜F, we can construct a weighted MCS tree $T$ and an unweighted line graph $L(H)$, namely the union join graph, in $\mathcal{O}(|H|+|E(\mathcal{U}(H))|)$. We simplify algorithm˜3 to algorithm˜8 in appendix˜F, which compares the LCA edges $\lambda(e)$ of each non-tree edge $e$ and slides it to construct the equivalent graph $G^{\equiv}(H)$ in $\mathcal{O}(|E(\mathcal{U}(H))|)$. As $\gamma$-acyclicity guarantees that every non-tree edge $e$ is an MST edge, we can use the weights of the LCA edges to determine how to slide $e$. Let $\lambda(e)=\{e_{l},e_{r}\}$. If $e_{l}=e_{r}$, we slide $e$ to be parallel to $e_{l}$. Otherwise, if $e_{l}$ and $e_{r}$ have different weights, we slide $e$ to be parallel to the heavier one. If they have the same weight, we slide $e$ to form a triangle with them.
The number of join trees is lower bounded by the number of non-tree edges in its line graph, namely $|\mathcal{T}(H)|=\Omega(|E(\mathcal{U}(H))|-|E(T)|)$. The total time complexity of enumerating all join trees is $\mathcal{O}(|H|+|E(\mathcal{U}(H))|+|\mathcal{T}(H)|)=\mathcal{O}(|H|+|\mathcal{T}(H)|)$.
5 The Canonical Join Tree of a Berge-Acyclic Query
An acyclic query can have exponentially many join trees with respect to its size. For example, the line graph of a clique query with $n$ relations is an $n$-clique $K_{n}$ with $n^{n-2}$ join trees by Cayley’s formula [7]. Enumerating all join trees can be prohibitive for large queries. On the other hand, the query optimizer does not need to consider all possible join trees to achieve good performance. For example, the implementation of Yannakakis’ algorithm by Zhao et al. [47] has similar performance on any join tree rooted at the largest relation. An alternative to enumeration is therefore to simply construct an arbitrary join tree for a given root. This can be done in linear time by the Maximum Cardinality Search (MCS) algorithm from a chosen relation as shown in algorithm˜1. In this section, we show that the MCS algorithm produces a shallowest tree for a Berge-acyclic query, in the sense that the depth of every node in the tree is minimized. We prove that this shallowest tree is unique, and therefore call it the canonical join tree:
Definition 25.
A join tree $T_{r}$ rooted at $r$ is canonical if $0pt(T_{r},r_{i})\leq 0pt(T^{\prime}_{r},r_{i})$ for any other join tree $T^{\prime}_{r}$ rooted at $r$ and any $r_{i}\in R(T_{r})=R(T^{\prime}_{r})$.
A shallow join tree has practical benefits. For example, the depth of the join tree determines the number of sequential steps required in a parallel join algorithm. A shallow join tree tends to be wide and have more leaves, allowing better utilization of indices.
Although Berge-acyclicity was thought to be too restrictive when it was first introduced to database theory [13], we found it to be general enough to cover almost all acyclic queries encountered in the wild. As shown in table˜1 of appendix˜F, among 10 45410\,454 queries from five popular benchmarks, $9285$ are $\alpha$-acyclic, and only 8 of these are not Berge-acyclic. In retrospect, this should not be surprising, as most joins in relational databases are over primary/foreign keys. Emerging workloads in graph databases usually involve simple graphs and seldom require composite key joins. A query without composite key joins admits a linear hypergraph, where each pair of hyperedges shares at most one vertex. The following result establishes an equivalence between $\alpha$-acyclicity with linearity and Berge-acyclicity.
Proposition 26.
An $\alpha$-acyclic hypergraph is Berge-acyclic if and only if it is linear.
By proposition˜26, every edge in the line graph of a Berge-acyclic hypergraph is labeled with a single variable, thus having a weight of 1. Every spanning tree is a maximum spanning tree, therefore a join tree.
Corollary 27.
For a Berge-acyclic hypergraph $H$, any spanning tree of $L(H)$ is a join tree.
In the rest of this section, we prove the existence and uniqueness of the canonical join tree rooted at any relation of a Berge-acyclic hypergraph, and show that it can be constructed by MCS as in algorithm˜1.
5.1 Existence & Uniqueness of the Canonical Join Tree
The key insight leading to the existence and uniqueness of the canonical join tree is that the line graph of a Berge-acyclic hypergraph is geodetic [31], meaning that there is a unique shortest path between any pair of vertices. To prove this, we first review some relevant concepts and properties of graphs. All graphs below are simple and undirected.
Definition 28.
A cycle in a graph is a sequence of distinct vertices $v_{0},\ldots,v_{k-1\geq 2}$ such that there is an edge between $v_{i}$ and $v_{(i+1)\bmod k}$ for all $1\leq i\leq k$.
A special class of graphs called chordal graphs are intimately related to acyclic hypergraphs:
Definition 29.
A graph is chordal if every cycle of length at least 4 has a chord, i.e., an edge that is not part of the cycle but connects two vertices of the cycle.
The key step in this section is to show the line graph of a Berge-acyclic hypergraph is a block graph, which is a special class of chordal graphs. There are many different characterizations of block graphs [18], and we present a recent one:
Definition 30 (Block Graph [10]).
A graph $G=(R,E)$ with the vertex set $R$ and the edge set $E$ is a block graph if it is chordal and diamond-free.
Removing an edge from a 4-clique $K_{4}$ results in a diamond. Diamond-free requires that no induced subgraph $G|_{R^{\prime}}$ is a diamond, where $R^{\prime}\subseteq R$ is any subset of vertices. We now present the main result relating block graphs to Berge-acyclic hypergraphs:
Lemma 31.
The line graph $L$ of a Berge-acyclic hypergraph $H$ is a block graph.
Together with the fact that every block graph is geodetic (it has a unique shortest path between any two vertices) [31], lemma˜31 implies the following corollary.
Corollary 32.
Let $L$ be the line graph of a Berge-acyclic hypergraph $H$. There is a unique shortest path between any two vertices in $L$.
We are now ready to prove the existence and uniqueness of the canonical join tree.
Theorem 33.
Let $L$ be the line graph of a Berge-acyclic hypergraph $H$, and $\mathsf{P}(r,r^{\prime})$ be the shortest path in $L$ between $r,r^{\prime}\in R(L)$, $T_{r}=\bigcup_{r^{\prime}\in R(L)}\mathsf{P}(r,r^{\prime})$ is the unique canonical join tree for $H$ rooted at $r$.
Proof.
A spanning tree of $L$ is its join tree by corollary˜27. To show $T_{r}$ is a join tree, it is sufficient to show $T_{r}$ is spanning, connected and acyclic.
$T_{r}$ is spanning and connected because it contains a path from $r$ to each $r^{\prime}\in R(L)$.
We prove $T_{r}$ is acyclic by induction on the distance $\mathsf{dist}(r,r^{\prime})$ between $r$ and $r^{\prime}$. Let $R_{d}=\left\{r^{\prime}\in R(L)\mid\mathsf{dist}(r,r^{\prime})\leq d\right\}$. $R_{0}=\{r\}$ contains only the root. The subgraph $T_{r}|_{R_{0}}$ is trivially acyclic. Assuming that $T_{r}|_{R_{d>0}}$ is acyclic, we consider a vertex $r^{\prime}\in R_{d+1}\setminus R_{d}$. corollary˜32 guarantees a unique shortest path between each pair of vertices $r,r^{\prime}\in R(L)$. Each $r^{\prime}$ is connected to a unique neighbor $r^{\prime\prime}\in R_{d}\setminus R_{d-1}$ that is at distance $d$ from $r$. Otherwise, there are at least two distinct shortest paths from $r$ to $r^{\prime}$. Therefore $T_{r}|_{R_{d+1}}$ is acyclic, and $T_{r}$ is a join tree.
The join tree $T_{r}$ is canonical, because the path from $r$ to each $r^{\prime}\in R(L)$ is the shortest and therefore minimizing the depth $0pt(T_{r},r^{\prime})$. The canonical tree is unique by corollary˜32. ∎
5.2 Construction of the Canonical Join Tree
Given a Berge-acyclic hypergraph $H$ and $r\in R(H)$ as the root, we can construct the canonical tree $T_{r}(H)$ by running algorithm˜1. We first present a useful lemma on the structure of the canonical join tree:
Lemma 34.
Let $T_{r}$ be the canonical join tree rooted at $r$ of a Berge-acyclic hypergraph. Along its root-to-leaf path $r,r_{1},\ldots,r_{k}$, each pair of adjacent vertices shares a distinct variable.
We now show that the MCS algorithm constructs the canonical join tree.
Theorem 35.
Given a Berge-acyclic hypergraph $H$, algorithm˜1 from $r\in R(H)$ constructs the canonical tree $T_{r}(H)$.
Proof.
algorithm˜1 labels each hyperedge in $R(H)$ in ascending order from the root $r_{1}=r$ to $r_{n=|R(H)|}$. When algorithm˜1 has labeled $i\in[n]$ hyperedges, we consider a graph $G_{i}$ whose vertices are the labeled hyperedges $\{r_{1},\ldots,r_{i}\}$ and edges are the parent-child relationships $\{(r_{2},p(r_{2})),\ldots,(r_{i},p(r_{i}))\}$.
To prove the claim by induction, we show a loop invariant that $G_{i}$ is a subtree of $T_{r}$ for all $i\in[n]$. An empty graph is trivially a subtree of any graph. When the root $r$ is labeled as $r_{1}$, this graph of a single vertex is also trivially a subtree of $T_{r}$. Assuming $G_{k}$ is a subtree of $T_{r}$, algorithm˜1 proceeds to label $r_{k+1}$ and assigns its parent $p(r_{k+1})=r_{p\leq k}$. Suppose for the sake of contradiction that $r_{k+1}$ has a different parent $r_{q}$ in $T_{r}$. Let us consider the possible relationships between $r_{p}$ and $r_{q}$. First, $r_{q}$ cannot be a descendant of $r_{p}$, otherwise there would be a shorter path from $r_{k+1}$ to the root $r$ going through $r_{p}$ instead of $r_{q}$, violating theorem˜33. Suppose $r_{q}$ is an ancestor of $r_{p}$. By the running intersection property, $r_{q}\cap r_{k+1}$ must contain $r_{p}\cap r_{k+1}$, but that would imply every vertex in $r_{p}\cap r_{k+1}$ is already marked by $r_{q}$ and the algorithm would not have assigned $r_{p}$ as the parent of $r_{k+1}$. Therefore, $r_{p}$ and $r_{q}$ are not descendants of each other. Denote their lowest common ancestor in $T_{r}$ as $r_{a}=\mathsf{LCA}(r_{p},r_{q})$. Let $\lambda(r_{p},r_{q})=\{e_{l},e_{r}\}$ as shown by the solid lines in fig.˜5 of appendix˜F, $\mathsf{P}_{p}$ be the path from $r_{a}$ to $r_{p}$ excluding $e_{l}$ and $\mathsf{P}_{q}$ be the path from $r_{a}$ to $r_{q}$ excluding $e_{r}$ as shown by the dashed lines in fig.˜5. By lemma˜34, the variables on $\mathsf{P}_{p}$ and $\mathsf{P}_{q}$ are all distinct. Consider the following possible cases:
Case 1:
If $r_{k+1}\cap r_{p}=r_{k+1}\cap r_{q}$ and $\chi(e_{l})=\chi(e_{r})$, then $\mathsf{P}_{p},\mathsf{P}_{q}$ form a Berge-cycle.
Case 2:
If $r_{k+1}\cap r_{p}=r_{k+1}\cap r_{q}$ and $\chi(e_{l})\neq\chi(e_{r})$, then $\mathsf{P}_{p},r_{a},\mathsf{P}_{q}$ form a Berge-cycle.
Case 3:
If $r_{k+1}\cap r_{p}\neq r_{k+1}\cap r_{q}$ and $\chi(e_{l})=\chi(e_{r})$, then $\mathsf{P}_{p},\mathsf{P}_{q},r_{k+1}$ form a Berge-cycle.
Case 4:
If $r_{k+1}\cap r_{p}\neq r_{k+1}\cap r_{q}$ and $\chi(e_{l})\neq\chi(e_{r})$, then $\mathsf{P}_{p},r_{a},\mathsf{P}_{q},r_{k+1}$ form a Berge-cycle.
All possible cases above contradict Berge-acyclicity, therefore, algorithm˜1 must correctly assign $r_{p}$ as the parent of $r_{k+1}$. ∎
6 Converting a Binary Join Plan to a Join Tree
A recent approach [20, 2] that converts a binary join plan into a join tree has gained popularity as it allows system builders to leverage existing query optimizers designed for binary join plans. In this section, we focus on an algorithm by Hu et al. [20] that converts left-deep linear join plans into join trees. The algorithm can convert all acyclic queries in standard benchmarks, including those in table˜1 of appendix˜F. We prove that the algorithm converts any connected left-deep linear join plan into a join tree if and only if the query is $\gamma$-acyclic. Our result yields a new characterization of $\gamma$-acyclic queries. We formally define binary join plans and describe the algorithm by Hu et al. [20] with algorithm˜6 in appendix˜F.
Definition 36.
A left-deep linear plan is a sequence of relations $r_{1},r_{2},\ldots,r_{n}$. It is connected if for each $r_{i\geq 2}$, $\exists r_{j<i}:r_{i}\cap r_{j}\neq\emptyset$.
Query optimizers strive to produce connected plans, to avoid expensive Cartesian products. Many optimizers produce exclusively left-deep linear plans. Plans that are not left-deep are called bushy, and such plans may still be decomposed into left-deep fragments [43].
Given a left-deep linear plan, Hu et al. [20] generate a join tree with algorithm˜6 in appendix˜F. The algorithm chooses the first relation $r_{1}$ as the root and iterates through the rest of the plan. For each relation $r_{i\in[2,n]}$, it finds the first relation $r_{j}$ that contains all attributes shared by $r_{i}$ with all previous relations $r_{i}\cap\bigcup r_{k<i}$, and assigns $r_{j}$ as the parent of $r_{i}$. The algorithm constructs a join tree if it finds a parent for each $r_{i}$.
Hu et al. [20] proved that the algorithm succeeds whenever the input plan is the reverse of a GYO-reduction order. They also observed that every left-deep linear plan produced for queries in standard benchmarks is indeed the reverse of a GYO-reduction order. This is not a coincidence, as we show that every connected left-deep linear join plan must be the reverse of a GYO-reduction order if and only if the query is $\gamma$-acyclic.
Theorem 37.
A query is $\gamma$-acyclic if and only if every connected left-deep linear join plan for the query is the reverse of a GYO-reduction order.
The “only if” direction is more difficult. We prove its contrapositive by constructing a $\gamma$-cycle from a query plan that is not the reverse of a GYO-reduction order. We start with a simple observation that directly follows from the definition of GYO-reduction order.
Proposition 38.
If a left-deep linear plan is not the reverse of a GYO-reduction order, then there exists a relation $r_{i}$ that has no parent among $\{r_{j<i}\}$, formally $\neg\exists r_{j<i}:(r_{i}\cap\bigcup r_{k<i})\subseteq r_{j}$. We call $r_{i}$ an orphan, and denote it as $\mathring{r_{i}}$.
We introduce an ordering on relations to compare them in a plan.
Definition 39.
Given a relation $r$, we write $r_{1}\geq_{r}r_{2}$ if $r_{1}\cap r\supseteq r_{2}\cap r$.
We first observe that $\geq_{r}$ is a preorder (reflexive and transitive), but not a partial order (antisymmetry may fail). Moreover, if there is a greatest element $r_{p}$ with respect to $\geq_{r}$, i.e., $\forall r^{\prime}:r_{p}\geq_{r}r^{\prime}$, then $r_{p}$ is a parent of $r$. Those observations lead to the following.
Lemma 40.
Let $\hat{r}$ be a maximal relation with respect to $\geq_{\mathring{r}}$ for an orphan $\mathring{r}$, meaning $\neg\exists r:r>_{\mathring{r}}\hat{r}$, then there is a relation $\bar{r}$ incomparable with $\hat{r}$, i.e., $\hat{r}\not\geq_{\mathring{r}}\bar{r}$ and $\bar{r}\not\geq_{\mathring{r}}\hat{r}$.
Proof.
If every relation is comparable with $\hat{r}$, the maximal relation $\hat{r}$ is the greatest element with respect to $\geq_{\mathring{r}}$, thus a parent of $\mathring{r}$. It contradicts the fact that $\mathring{r}$ is an orphan. ∎
Given a pair of incomparable relations $\hat{r}$ and $\bar{r}$ with respect to $\mathring{r}$, we can find at least one variable in each relation that does not appear in the other relation but appears in $\mathring{r}$, namely $\exists\bar{x}\in\bar{r}\cap\mathring{r}:\bar{x}\notin\hat{r}$ and $\exists\hat{x}\in\hat{r}\cap\mathring{r}:\hat{x}\notin\bar{r}$. We say that $\hat{x}$ is a dangling variable of $\hat{r}$ and $\bar{x}$ is a dangling variable of $\bar{r}$. We are now ready to prove theorem˜37. Our strategy is to derive a $\gamma$-cycle from any orphan relation, leading to a contradiction.
Proof of theorem˜37.
To prove the “only if” direction by contradiction, we assume that a connected left-deep linear plan for a $\gamma$-acyclic query is not the reverse of a GYO-reduction order. By proposition˜38, there is an orphan $\mathring{r}$ in the plan. Assuming that $\mathring{r}$ is the $t$-th relation in the plan, we now consider the prefix of the plan up to $\mathring{r}$, namely $\mathsf{pre}({t})=r_{1},\ldots,r_{t-1},\mathring{r}$. Let $\hat{r}\in\mathsf{pre}({t-1})$ be a maximal relation with respect to $\geq_{\mathring{r}}$, lemma˜40 guarantees an $\bar{r}\in\mathsf{pre}({t-1})$ incomparable with $\hat{r}$. There are at least two dangling variables $\bar{x}\in\bar{r}\cap\mathring{r}\setminus\hat{r}$ and $\hat{x}\in\hat{r}\cap\mathring{r}\setminus\bar{r}$.
Because the shorter prefix $\mathsf{pre}({t-1})=r_{1},\ldots,r_{t-1}$ is connected, there is a path in $\mathsf{pre}({t-1})$ between $\hat{r}$ and $\bar{r}$. The incomparability guarantees that $\bar{r}\cap\mathring{r}\not\subseteq\hat{r}\cap\mathring{r}$. We can choose a shortest path in $\mathsf{pre}({t-1})$ that connects a variable in $\hat{r}\cap\mathring{r}$ to a variable in $\mathring{r}\setminus\hat{r}$. We now consider two exhaustive cases where each yields a $\gamma$-cycle:
Case 1:
If the shortest path consists of a single relation $\tilde{r}$, $\tilde{r}\not>_{\mathring{r}}\hat{r}$ (because $\hat{r}$ is maximal) and $\tilde{r}\not\leq_{\mathring{r}}\hat{r}$ (because $\tilde{r}$ contains some variable in $\mathring{r}\setminus\hat{r}$). Therefore, $\tilde{r}$ is incomparable with $\hat{r}$. Since $\tilde{r}\cap\hat{r}\neq\emptyset$, $\hat{r},\hat{x},\mathring{r},\tilde{x},\tilde{r},\breve{x}$ form a $\gamma$-cycle where $\hat{x}\in\hat{r}\cap\mathring{r}$, $\tilde{x}\in\tilde{r}\cap\mathring{r}$, and $\breve{x}\in\tilde{r}\cap\hat{r}$.
Case 2:
If the shortest path consists of at least two relations, let $r_{0}=\mathring{r}$, and the path be $r_{1},\ldots,r_{k-1}$. We now show that $r_{0},r_{1},\ldots,r_{k-1}$ form a pure cycle where each $r_{i}$ is only adjacent to $r_{(i-1)\bmod k}$ and $r_{(i+1)\bmod k}$. First, if a $r_{i\in[2,k-2]}$ on the path is adjacent to a $r_{j\in[1,k-1]}$ other than its neighbors $r_{i-1}$ or $r_{i+1}$, the path can be shortened, contradicting the shortest path assumption. Second, if the relation $r_{i\in[2,k-2]}$ is adjacent to $r_{0}$, then $r_{i}$ must contain variables in either $r_{0}\cap\hat{r}$ or $r_{0}\setminus\hat{r}$. We can shrink the path to $r_{1},\ldots,r_{i}$ or $r_{i},r_{k-1}$, which also contradicts the shortest path assumption. Therefore, each $r_{i\in[0,k-1]}$ on the cycle is only adjacent to $r_{(i-1)\bmod k}$ and $r_{(i+1)\bmod k}$. The cycle $r_{0},r_{1},\ldots,r_{k-1}$ is a pure cycle of length $k\geq 3$. A pure cycle is always a $\gamma$-cycle.
This concludes the proof for the “only if” direction.
Now we prove the “if” direction by contrapositive, namely that if a query contains a $\gamma$-cycle, then there exists a connected left-deep linear plan that is not the reverse of a GYO-reduction order. Let the $\gamma$-cycle of length $k\geq 3$ be $r_{0},\ldots,r_{i},\ldots,r_{k-1}$ where $r_{i\in[1,k-2]}$. By definition, there is an $x_{i-1}$ appearing exclusively in $r_{i-1}$ and $r_{i}$ and an $x_{i}$ appearing exclusively in $r_{i}$ and $r_{i+1}$. We modify the original plan by moving $r_{i}$ after $r_{k-1}$ such that $r_{0},\ldots,r_{i-1},r_{i+1},\ldots,r_{k-1},r_{i}$ becomes part of the new query plan.
The prefix $\mathsf{pre}({k-1})=r_{0},\ldots,r_{i-1},r_{i+1},\ldots,r_{k-1}$ is connected. Therefore, the new query plan is connected. However, $r_{i}$ has no parent, because no relation in $\mathsf{pre}({k-1})$ contains both $x_{i-1}$ and $x_{i}$. This implies the plan is not the reverse of a GYO-reduction order. ∎
7 Conclusion and Future Work
We proposed three approaches for constructing join trees. Our enumeration algorithm in section˜4 generates join trees by edits with amortized constant delay, forming a basis for cost-based optimization of acyclic joins. In section˜5, we showed that the Maximum Cardinality Search algorithm constructs the unique shallowest join tree for any Berge-acyclic query, supporting very large queries and enabling parallel execution. Finally, in section˜6, we gave theoretical justification for converting binary join plans to join trees, allowing reuse of existing optimization infrastructure.
Future work includes compact representations of join trees for dynamic programming, as in binary plan optimizers, and the challenging cost estimation for Yannakakis-style algorithms: the random-walk approach [26] models joint probabilities for binary joins, whereas an efficient and accurate solution for semijoins remains to be found.
Our results also raise theoretical questions. Can join tree enumeration achieve worst-case constant delay? We proved Berge-acyclicity sufficient for the existence and uniqueness of the canonical join tree, but it is not necessary, and $\gamma$-acyclicity is insufficient as shown in fig.˜6 of appendix˜F. What is the precise characterization of hypergraphs that admit a unique canonical join tree for any root, or for some root?
Appendix A Extended Preliminaries
Definition 41 (Hypergraph Acyclicity).
A hypergraph $H$ is:
  • $\alpha$-acyclic if it admits a join tree;
  • $\beta$-acyclic if every subgraph of $H$ is $\alpha$-acyclic;
  • $\gamma$-acyclic if it does not contain any $\gamma$ cycle. A $\gamma$ cycle is a sequence of length $k\geq 3$ of distinct hyperedges and distinct vertices $(r_{0},x_{0},\dots,r_{k-1},x_{k-1})$ such that every $x_{i\in[0,k-2]}$ belongs to $r_{i}\cap r_{i+1}$ and no other $r_{j}$ while $x_{k-1}$ belongs to $r_{k-1}\cap r_{0}$ and possibly other hyperedges;
  • Berge-acyclic if it does not contain any Berge cycle. A Berge cycle is a sequence of length $k\geq 2$ of distinct vertices and distinct hyperedges $(r_{0},x_{0},\dots,r_{k-1},x_{k-1})$ such that $\forall\ i\in[k]:x_{i}\in r_{i}\cap r_{(i+1)\bmod k}$.
The four notions of acyclicity form a strict hierarchy that Berge-acyclic $\Rightarrow$ $\gamma$-acyclic $\Rightarrow$ $\beta$-acyclic $\Rightarrow$ $\alpha$-acyclic [13].
Given a hypergraph $H$ with $n$ hyperedges, algorithm˜1 labels each hyperedge with an integer $i\in[1,n]$ in ascending order and assigns a parent $p(r_{i})$ to each hyperedge $r_{i}$. Upon labeling an edge the algorithm marks all previously unmarked vertices contained in that edge. Initially, all hyperedges are unlabeled and all vertices are unmarked. The algorithm begins by assigning nullnull as the parent of the root $r$ and initializing label $i$ to 0. In each following while loop iteration, the algorithm increments $i$ by one and labels the current hyperedge as $r_{i}$. For each unmarked vertex $x$ in $r_{i}$ (a vertex is considered marked when it is removed from $X$), it assigns $r_{i}$ tentatively as the parent of each unlabeled hyperedge $r^{\prime}$ containing $x$. The algorithm marks all unmarked vertices in $r_{i}$ and removes $r_{i}$ from future consideration. Finally, it selects a hyperedge with the most marked vertices to label next, breaking ties arbitrarily. The algorithm terminates when all hyperedges are labeled. Tarjan and Yannakakis [40] describes an efficient implementation of algorithm˜1 running in $\mathcal{O}(|H|)$.
Example 42.
Considering the hypergraph $H_{6}$ shown in fig.˜2(a), we run algorithm˜1 from the hyperedge $P$. It is labeled as $r_{1}$ and all its vertices are marked. $P$ is also tentatively assigned as the parent of five hyperedges such that $p(S)=p(T)=p(U)=p(W)=p(Y)=r_{1}$. We proceed to label the next hyperedge with the most marked vertices. At this point, each unlabeled hyperedge has one marked vertex, so we can choose any of them, say $S$, and label it as $r_{2}$.
Because the labeling of $S$ marks no additional vertices in $T,U,W,Y$, their parents remain unchanged. These remaining hyperedges are again tied with the same number of marked vertices, so we can choose any of them, say $T$, and label it as $r_{3}$. Since the labeling of $T$ marks a vertex $c$ in $U,W$, their parents are updated to $p(U)=p(W)=r_{3}$.
We continue the process until all hyperedges are labeled and all vertices are marked. By connecting each labeled hyperedge with its parent, we construct the join tree $T^{M}$ shown in fig.˜2(d).
Notice that algorithm˜1 can generate different join trees for the same hypergraph by breaking ties differently. However, it does not guarantee to generate all possible trees. For example, it never generates the Join Tree $T^{G}$ shown in fig.˜2(c).
Appendix B Data Structures for Constant Time Operations on Graphs
Given line graph $L$ and MCS tree $T$, algorithm˜3 iterates over non-tree edges of $L$. Moreover, the algorithm uses data structures $0pt(\cdot)$, $\mathsf{LA}(\cdot)$ and $\mathsf{LCA}(\cdot)$, in determining the endpoints of a non-tree edge’s LCA edges. In this section we describe briefly how iterating over non-tree edges can be achieved in time linear in the number of vertices and edges of $L$ and how determining the endpoints of a non-tree edge’s LCA edges can be achieved in constant time.
Given a graph $G=(V,E)$, we assume the vertices are represented by integers from 1 to $|V|$. Given a tree $T$ of $G$ rooted at node $r$, recall that data structure $0pt_{T}(\cdot)$, stores for each node $r_{i}$ the length of the shortest path between $r$ and $r_{i}$ in $T$. Data structure $\mathsf{LCA}_{T}(\cdot)$, stores for each pair of nodes $r_{i},r_{j}$ the lowest common ancestor of $r_{i},r_{j}$ in $T$. Finally, data structure $\mathsf{LA}_{T}(\cdot)$ stores for each node-integer pair $r_{i},j$, the ancestor of node $r_{i}$ at depth $j$. Each of these data structures can be established from a tree $T$ in time linear in the size of the tree and support constant time look-ups.
Observation 43.
Given rooted tree $T$, data structures $0pt(\cdot)$, $\mathsf{LCA}(\cdot)$ and $\mathsf{LA}(\cdot)$ can be established in time $\mathcal{O}(|T|)$ and support $\mathcal{O}(1)$ time look-ups [9, 33, 3, 4].
Given the line graph $L$ of an $\alpha$-acyclic hypergraph $H$ and an MCS tree $T$ of $H$, recall that the LCA edges $\lambda(e)$ of a non-tree edge $e=(r_{i},r_{j})$ of $T$ is defined as the set of edges incident to $\mathsf{LCA}(r_{i},r_{j})$ that lie on the unique path between $r_{i}$ and $r_{j}$ in $T$. The following observation shows that an edges LCA edges can be found in constant time given data structures $0pt(\cdot)$, $\mathsf{LCA}(\cdot)$ and $\mathsf{LA}(\cdot)$.
Observation 44.
Given the line graph $L$ of an $\alpha$-acyclic hypergraph $H$, an MCS tree $T$ and data structures $0pt(\cdot)$, $\mathsf{LCA}(\cdot)$ and $\mathsf{LA}(\cdot)$, the LCA edges of any non-tree edge can be found in constant time.
Proof.
Given a non-tree edge $e=(r_{i},r_{j})$ we find $\mathsf{LCA}(r_{i},r_{j})=l$ in constant time by ˜43. The depth $d=0pt(l)+1$ is likewise found in constant time. The edge $e_{1}=(l,r_{1})$, where node $r_{1}=\mathsf{LA}(r_{i},d)$ is found by a constant time look-up, is adjacent to $l$ and lies on the unique path between $r_{i}$ and $r_{j}$. The same holds for edge $e_{2}=(l,r_{2})$ where node $r_{1}=\mathsf{LA}(r_{j},d)$. ∎
In what remains of this section we describe how algorithm˜3 iterates over and performs sliding transformations on non-tree edges.
We use an array $L$ to represent the adjacency list of a line graph where entry $L[i]$ consists of a linked list of tuples $(j,\omega)$ where $j$ is a neighbor of $i$ and $\omega$ is the weight of the edge $(i,j)$. A rooted tree $T_{r}$ is likewise implemented as an array, where the index $i$ represents the vertex $i$, and the $i$-th entry $T_{r}[i]$ stores $(p,\omega)$ where $p$ is the parent of $i$ and $\omega$ is the weight of the edge $(i,p)$. The depth table of the tree is also implemented as an array to enable constant time lookup for the depth of a vertex $0pt(T_{r},i)$.
appendix˜C shows how to construct the weighted line graph and the hypergraph given the original query. Then, using the hypergraph we can construct $T_{r}$ and $0pt$ during the MCS algorithm in linear time [40].
As we iterate over each edge $(j,\omega)$ of $L[i]$, we check if the current edge $(i,j)$ is a tree edge by performing constant-time query of $T_{r}[i]$ and $T_{r}[j]$ to see if one of them is the parent of the other. From ˜44 we know that we can find the LCA edges $\lambda(i,j)=\{e_{1},e_{2}\}$ in constant time. There’s no general solution to lookup an entry in a linked list in constant time. Instead of iterating over the linked-list representation of $L$ to find the edge weights of $e_{1}$ and $e_{2}$, we can lookup their weights from the array representation of $T_{r}$ in constant time. Let $e_{1}=(r_{1},p(r_{1}))$ and $e_{2}=(r_{2},p(r_{2}))$ with weights $\omega_{1}$ and $\omega_{2}$. Because both $e_{1}$ and $e_{2}$ are tree edges, we can find their weights by querying $T_{r}[r_{1}]=(p(r_{1}),\omega_{1})$ and $T_{r}[r_{2}]=(p(r_{2}),\omega_{2})$. Each of these operations takes constant time. When we need to delete an edge $e^{*}=(i,j)$ from $L$, we first check if $e^{*}$ is a tree edge by checking $T_{r}[i]$ and $T_{r}[j]$. Deleting $e^{*}$ from the linked list takes constant time. Sliding an edge involves updating the endpoints of the edge which takes constant time.
Appendix C From Queries to (Hyper-)Graphs
algorithm˜3 assumes the weighted line graph and an initial MCS tree are given as input. In practice, we need to construct these graphs from the original SQL query.
We model a query involving $m$ relations as a set of join predicates each in the form of $R_{a}.x_{i}=R_{b}.x_{j}$ such that $a\neq b\wedge a,b\in[m]$, namely no self-joins or duplicates, and assume each relation is uniquely identified by an integer in $[m]$. For each relation $R_{a}$ with arity $k_{a}$, each attribute $x_{i\in[k_{a}]}$ is uniquely identified by an integer in $k_{a}]$. Therefore each predicate is uniquely identified by a tuple $(a,i,b,j)$. Each query $Q$ is represented by a set of such tuples. We will assume that query $Q$ is transatively closed over join predicates, i.e. if $R_{a}.x_{i}=R_{b}.x_{j}$ and $R_{b}.x_{j}=R_{c}.x_{k}$ are in $Q$, then $R_{a}.x_{i}=R_{c}.x_{k}$ is also in $Q$. The tuple $(a,i,b,j)$ is present in $Q$ if and only if $R_{a}.x_{i}=R_{b}.x_{j}$ is a join predicate over relations $R_{a},R_{b}$. We define the size of a query $|Q|$ as the number of join predicates in $Q$, equivalent to the number of $(a,i,b,j)$ tuples. Our goal is to construct the query hypergraph and the weighted line graph from this set of tuples.
C.1 Construction of Hypergraph
We construct the hypergraph $H$ by finding connected components of the predicate graph, which is a graph where there is a vertex for each attribute $R_{a}.x_{i}$, and there is an edge between $R_{a}.x_{i}$ and $R_{b}.x_{j}$ for every predicate $R_{a}.x_{i}=R_{b}.x_{j}$. Each connected component of the predicate graph corresponds to an “equivalence class” of attributes. We then assign a unique variable to each connected component, and map each attribute to the variable of its component. This can be done in linear time using standard algorithms for computing connected components [9, 39].
algorithm˜4 generates hypergraph $H$ from query $Q$. $H$ is represented as an array of size $m$ where each entry is a list of integers. The algorithm starts by initializing $H$ to an array of size $m$ and $P$ to an array of size $p$, where $p$ is the number of distinct relation-attribute pairs $R_{a}.x_{i}$ in $Q$ as shown by Lines 4-4. If numbers $m$ and $p$ are not known a priori they can be found in a constant number of passes over elements in $Q$. Array $H$ will serve as a hyperedge list of the query hypergraph and array $P$ will serve as an adjacency list representation of the query predicate graph. The subsequent for-loop fills out adjacency list $P$ at Lines 4-4. We assume here that each $(a,i)$ pair is mapped to a unique integer in $[p]$ by a perfect hash function $f$ at Line 4, otherwise such a mapping can be established in a constant number of passes over $Q$. Next, a standard algorithm [9, 39] for the enumeration of connected components of a graph is called in Line 4. This produces array $\mathcal{C}$ where each entry is a list of vertices that forms a connected component in $P$. The subsequent nested for-loop iterates over each relation in each component to fill out hyperedge list $H$ at Lines 4-4.
Input: Query $Q$
Output: Query hypergraph $H$
1
2for $i\in[1\ldots m]$ do
3 $H[i]\leftarrow\texttt{linkedList}()$
4
5for $i\in[1\ldots p]$ do
6 $P[i]\leftarrow\texttt{linkedList}()$
7
8for $(a,i,b,j)\in Q$ do
9 $p=f(a,i)$
10 $P[p].\text{append}(b,j)$
[…]
12
13for $i\in[|\mathcal{C}|]$ do
14 for $R_{a}\in\mathcal{C}[i]$ do
15    $H[a].\text{append}(i)$
16
17
return $H$
Algorithm 4 Generating query hypergraph $H$ from query $Q$.
Observation 45.
algorithm˜4 generates the hypergraph of $Q$ in time $\mathcal{O}(|Q|)$.
Proof.
Arrays $H$ and $P$ can be initialized in time $\mathcal{O}(|Q|)$ since $m\leq|Q|$ and $p\leq|Q|$. The subsequent for-loop iterates over all elements in $|Q|$ making a single insertion to $P$ thus performing at most $\mathcal{O}(|Q|)$ steps. Predicate graph $P$ therefore has size at most $\mathcal{O}(|Q|)$. Generating a list of connected components of predicate graph $P$ is done in time linear in $|P|=\mathcal{O}(|Q|)$ by standard algorithms [9, 39]. The subsequent nested for-loop iterates over each relation in each component performing a single insertion on each iteration. The sum of the sizes of each component is exactly number of vertices of $P$ and therefore $\mathcal{O}(|Q|)$ elementary operations are performed. ∎
C.2 Construction of Weighted Line Graph
To construct the weighted line graph, we will ignore the attribute indices $(i,j)$ and focus on the relation indices $(a,b)$ for each join predicate in $Q$. The set of predicates can then be seen as representing the edge list of a multigraph. algorithm˜5 generates the weighted line graph of $Q$ by converting this multigraph into a weighted graph $L$ where the weight of each edge $(u,v)$ is the number of parallel edges between $u$ and $v$ in the multigraph.
The algorithm first initializes arrays $L$ and $M$ both of size $m$, where $M$ will be an adjacency list representation of a multigraph and $L$ will be adjacency list representation of the final weighted line graph in Line 5-5. Auxiliary array weight is initialized in Line 5. Entries of weight are initialized to be an empty linked list. In the subsequent for-loop, "edges" in $Q$ are inserted into multigraph $M$ in Line 5. Array weight[j] stores the weight of the edge $(i,j)$ for the current $i$. Entries of weight[j] are initialized to 0. The subsequent for-loop iterates over adjacency list $M$ in Line 5. The first nested for-loop at Line 5 iterates over each neighbor $j$ of $i$ stored in $M[i]$. For each neighbor $j$, array weight[j] is incremented by 1. Array weight[j] now contains the weights of all edges adjacent to $i$ at its non-zero entries. The second nested for-loop from Line 5 to Line 5 iterates over each neighbor $j$ one more time and inserts neighbor and weight pairs into array $L$ if weight[j] is not 0, otherwise skipping to the next neighbor. After an insertion is made, weight[j] is set to 0. A weight of 0 at weight[j] indicates the edge at $(i,j)$ and its weight has already been outputted. This prevents the insertion of duplicate edges and resets array weight[j] to be reused for the next relation in $M$. Since each array operation takes constant time and we visit each edge a constant number of times, the overall complexity of this procedure is linear in the size of the input.
Input: Query $Q$
Output: Weighted line graph $L$
1
2for $i\in[1\ldots m]$ do
3 $L[i]\leftarrow\texttt{linkedList}()$
4 $M[i]\leftarrow\texttt{linkedList}()$
5 $\texttt{weight}[i]\leftarrow 0$
6
7for $(a,i,b,j)\in Q$ do
8 $M[a].\text{append}(b)$
9
10for $R_{i}\in M$ do
11 for $R_{j}\in M[i]$ do
12    $\texttt{weight}[j]\leftarrow\texttt{weight}[j]+1$
13 for $R_{j}\in M[i]$ do
14    if $\texttt{weight}[j]\neq 0$ then
15       $L[i].\text{append}(R_{j}.\texttt{weight}[j])$
16       $\texttt{weight}[j]\leftarrow 0$
17
18
19
return $L$
Algorithm 5 Generating weighted line graph $L$ from query $Q$.
Observation 46.
algorithm˜5 generates the weighted line graph of $Q$ in time $\mathcal{O}(|Q|)$.
Proof.
Array initializations for $L$ and $M$ take $\mathcal{O}(m)$ time where $m\leq|Q|$. The first for–loop scans all predicates in $Q$ and inserts the corresponding edges in the multigraph $M$, performing $\mathcal{O}(|Q|)$ insertions, so $|M|=\mathcal{O}(|Q|)$. The auxiliary array weight has size $m$ and is initialized in $\mathcal{O}(m)$ time.
In the subsequent traversal of $M$, each adjacency list is visited once. For every neighbor $j$ of a vertex $i$, the first inner loop performs a constant–time update of weight[j]; the second inner loop performs at most one constant–time insertion of $(j,\texttt{weight[j]})$ into $L$ and resets weight[j]. Thus each stored neighbor (i.e., each edge occurrence in $M$) incurs only $\mathcal{O}(1)$ work overall. Since the total number of stored neighbors is $\Theta(|M|)=\mathcal{O}(|Q|)$, the time for these loops is $\mathcal{O}(|Q|)$.
Summing up, all steps are linear in $|Q|$, hence algorithm˜5 runs in time $\mathcal{O}(|Q|)$. ∎
Appendix D Missing Proofs in section˜4
lemma˜15 If $H^{\prime}\twoheadrightarrow H$ then $\mathcal{T}(H)\subseteq\mathcal{T}(H^{\prime})$.
Proof.
Let $T$ be a join tree of $H$, and $x^{\prime}$ be any vertex in $H^{\prime}$. Consider the neighborhood $H^{\prime}|_{x^{\prime}}$ of $x^{\prime}$ in $H^{\prime}$ comprising the set of hyperedges $\left\{r^{\prime}\in H^{\prime}\mid x^{\prime}\in r^{\prime}\right\}$. Since $H^{\prime}\twoheadrightarrow H$, we have that $x^{\prime}\in r^{\prime}$ if and only if $\mathsf{f}(x^{\prime})\in\mathsf{f}(r^{\prime})=r^{\prime}$. The neighborhood $H|_{\mathsf{f}(x^{\prime})}$ of $\mathsf{f}(x^{\prime})$ in $H$ comprises of the set of hyperedges $\left\{r\in H\mid\mathsf{f}(x^{\prime})\in r\right\}$. Therefore, $r^{\prime}\in H^{\prime}|_{x^{\prime}}\iff\mathsf{f}(r^{\prime})=r^{\prime}\in H|_{\mathsf{f}(x^{\prime})}$. Because $T$ is a join tree of $H$, $H|_{\mathsf{f}(x^{\prime})}$ is connected in $T$, therefore $H^{\prime}|_{x^{\prime}}$ is also connected in $T$ and $T$ is also a join tree of $H^{\prime}$. ∎
lemma˜16 Let $T$ be an MCS tree. For any edge $e\in T$ that has a parent, then $e\not\subseteq p(e)$.
Proof.
In algorithm˜1, a child $r_{c}$ is only connected to a parent $r$ if they share some previously unmarked vertex $x$. Because every vertex is marked only once, and the parent $r_{p}$ of $r$ was labeled before $r$, $x$ cannot be marked by $r$’s parent $r_{p}$. Let the $e$ be edge between $r$ and $r_{c}$, then the edge between $r$ and $r_{p}$ is $p(e)$, and $x\in e$ but $x\notin p(e)$, therefore $e\not\subseteq p(e)$. ∎
lemma˜17 Let $T$ be an MCS tree. For two edges $e,e^{\prime}\in T$ that have parents,
$(e\setminus p(e))\ \cap\ (e^{\prime}\setminus p(e^{\prime}))\ \neq\ \emptyset\quad\implies\quad p(e)=p(e^{\prime}).$
Proof.
We prove by contrapositive, assuming $p(e)\neq p(e^{\prime})$.
Case 1:
$p(e^{\prime})=e$, then $(e\setminus p(e))\cap(e^{\prime}\setminus e)\subseteq(e\cap e^{\prime})\setminus e=\emptyset$.
Case 2:
$p(e^{\prime})\neq e$. Without loss of generality, we suppose $e$ is no further from the root than $e^{\prime}$. Let $p(e^{\prime})=e^{\prime\prime}$, then $e^{\prime\prime}$ is on the path between $e$ and $e^{\prime}$ in $T$. If $(e\setminus p(e))\cap(e^{\prime}\setminus e^{\prime\prime})\neq\emptyset$, then $\exists\ x\in(e^{\prime}\cap e)\setminus e^{\prime\prime}$, which violates the running intersection property of join trees.
lemma˜19 Let $H^{\prime}$ be a duplicated hypergraph of $H$. Then,
  1. 1.
    $H^{\prime}\twoheadrightarrow H$ and $H^{\prime}\twoheadleftarrow H$
  2. 2.
    $\mathcal{T}(H^{\prime})=\mathcal{T}(H)$
  3. 3.
    $L(H^{\prime})=L(H)$
Proof.
1. We only need to show that $H^{\prime}\twoheadrightarrow H$ ($H^{\prime}\twoheadleftarrow H$) holds for $H^{\prime}$ derived from $H$ by a single vertex duplication. $H^{\prime}\twoheadrightarrow H$ ($H^{\prime}\twoheadleftarrow H$) will then hold for any duplicate $H^{\prime}$ derived by a series of vertex duplications since strong homomorphisms are closed under composition.
Suppose $H^{\prime}$ is derived from $H$ by duplicating $x\in X(H)$ to $x^{\prime}$.
Consider a pair of functions $(\mathsf{f}_{X}:X(H^{\prime})\mapsto X(H),R(H)_{\texttt{id}})$ where $\mathsf{f}_{X}$ maps $x^{\prime}$ to $x$. Otherwise, it is an identity map $X(H)_{\texttt{id}}$. This function pair constitutes a strong hypergraph homomorphism $H^{\prime}\twoheadrightarrow H$. Likewise, the pair of functions $(X(H)_{\texttt{id}},R(H)_{\texttt{id}})$ constitutes a homomorphism $H\twoheadrightarrow H^{\prime}$.
2. It follows from lemma˜15 and the facts above.
3. By the strong homomorphisms, the hyperedge sets have the same cardinality $|R(H)|=|R(H^{\prime})|$. So do the vertex sets of the line graphs, $R(L(H))=R(H)$ and $R(L(H^{\prime}))=R(H^{\prime})$. By construction, the vertex duplication adds $x^{\prime}$ to each hyperedge containing $x$, therefore does not create any additional edges in the line graph, $E(L(H^{\prime}))=E(L(H))$. ∎
Appendix E Missing Proofs in section˜5
proposition˜26 An $\alpha$-acyclic hypergraph is Berge-acyclic if and only if it is linear.
Proof.
Fagin [13] showed that every Berge-acyclic hypergraph is linear. We prove the other direction by contradiction, by assuming that $H$ is $\alpha$-acyclic and linear. We also assume that there is a Berge cycle $(r_{0},x_{0},\dots,r_{k-1},x_{k-1})$ of length $k\geq 2$. Without loss of generality, we suppose $r_{1}$ is the first hyperedge on the cycle to be removed by GYO reduction. Then $r_{1}$’s parent $r_{p}$ must satisfy $r_{p}\cap r_{1}\supseteq(r_{1}\cap\bigcup_{i\in[k]\wedge i\neq 1}r_{i})=\{x_{0},x_{1}\}$, contradicting $H$’s linearity. ∎
lemma˜31 The line graph $L$ of a Berge-acyclic hypergraph $H$ is a block graph.
Proof.
A Berge-acyclic hypergraph is $\gamma$-acyclic [13]. Zhu proves that the line graph of a $\gamma$-acyclic hypergraph is chordal [48]. We now prove that the line graph is diamond-free by showing that every diamond must be part of a $K_{4}$. Assuming a diamond $(r_{1},r_{2},r_{3},r_{4})$ in $L$ without the edge $(r_{1},r_{3})$, it can be treated as two triangles $(r_{1},r_{2},r_{4})$ and $(r_{2},r_{4},r_{3})$ sharing an edge $(r_{2},r_{4})$. We now consider the possible edge labelings in each triangle:
Case 1:
If all edges are labeled with distinct variables, each triangle forms a Berge-cycle, contradicting the Berge-acyclicity.
Case 2:
If two edges are labeled with the same variable, all three vertices share that variable. The third edge must be labeled with the same variable.
Since the two triangles share the edge $(r_{2},r_{4})$, all their edges are labeled with the same variable $x$, namely $x\in r_{1}\cap r_{2}\cap r_{3}\cap r_{4}$. There has to be an edge $(r_{1},r_{3})\in L$, contradicting the assumption that $(r_{1},r_{2},r_{3},r_{4})$ is a diamond. ∎
lemma˜34 Let $T_{r}$ be the canonical join tree rooted at $r$ of a Berge-acyclic hypergraph. Along its root-to-leaf path $r,r_{1},\ldots,r_{k}$, each pair of adjacent vertices shares a distinct variable.
Proof.
Let $r_{0}=r$, proposition˜26 requires that $|r_{i-1}\cap r_{i}|=1$ for all $i\in[k]$. By theorem˜33, the path $r_{0},\ldots,r_{k}$ is the shortest path from $r_{0}$ to $r_{k}$. The variable shared by two adjacent vertices $r_{i-1},r_{i}$ is equivalent to the edge label $\chi(r_{i-1},r_{i})$. The edge labels along the path are all distinct. Otherwise, two consecutive edges with the same label allows for shortening the path by removing either, contradicting the shortest path property. Two non-consecutive edges with the same label violate the running intersection property. ∎
Appendix F Additional Figures, Tables and Algorithms
Name # Queries # $\alpha$-Acyclic # Composite-Key Joins # Berge-Acyclic
TPC-H[41] 22 21 2 19
JOB[24] 113 113 0 113
STATS-CEB[17] 2603 2603 0 2603
CE[8] 3004 1839 0 1839
Spider-NLP[46] 4712 4709 6 4703
Table 1: Acyclic queries in the benchmarks (all $\alpha$-acyclic queries are also $\gamma$-acyclic.)
Input: Left-deep plan $r_{1},r_{2},\dots,r_{n}$
Output: Tree rooted at $r_{1}$
1
[…]
3 for $i\leftarrow 2$ to $n$ do
4   key $\leftarrow r_{i}\cap\bigcup r_{k<i}$
5
6 for $j\leftarrow 1$ to i1i-1 do
7    if $\text{key}\subseteq r_{j}$ then
8       $p(r_{i})\leftarrow r_{j}$
9       break
10
11      end if
12
13   end for
14
15 end for
Algorithm 6 Converting a binary plan to a join tree
Figure 5: Possible cases of how $r_{k+1}$ is connected to $G_{k}$
(a) $H_{3}$
(b) $T_{S}(H_{3})$
(c) $H_{5}$
(d) $T_{W}(H_{5})$
(e) $T^{\prime}_{W}(H_{5})$
Figure 6: $H_{3}$ is a non-Berge hypergraph that admits a unique canonical join tree $T_{S}(H_{3})$ with any relation chosen as root, such as $T_{S}(H_{3})$. $H_{5}$ is a $\gamma$ hypergraph that does not admit unique canonical join trees at any relation chosen as root. For example, $T_{W}(H_{5})$ and $T^{\prime}_{W}(H_{5})$ are MCS trees generated by algorithm˜1. Neither of them is a canonical join tree rooted at $W$.
Input: Hypergraph $H$, root $r$
Output: weighted join tree $T_{r}$ and unweighted line graph $L$
1
[…]
4while $R\neq\emptyset$ do
5 $i\leftarrow i+1;r_{i}\leftarrow r$
6 for $x\in r_{i}\cap X$ do
7    for distinct $r^{\prime},r^{\prime\prime}\in R:x\in r^{\prime}\cap r^{\prime\prime}$ do
8       $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}E(L)\leftarrow E(L)\cup\{(r^{\prime},r^{\prime\prime})\}$
9       $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\omega(r^{\prime},r_{i})\leftarrow|r^{\prime}\cap r_{i}|$
10       $p(r^{\prime})\leftarrow r_{i}$
11
12
13 $X\leftarrow X\setminus r_{i};R\leftarrow R\setminus\{r_{i}\}$
14 $r\leftarrow\operatorname*{arg\,max}_{r\in R}|r\setminus X|$
15
return $L,p(\cdot),\omega(\cdot)$
Algorithm 7 MCS for the $\gamma$-acyclic
Input: weighted MCS tree $T_{r}(H)$ weight function $\omega$ and unweighted union join graph $L(H)$
1 $G^{\equiv}\leftarrow L(H)$
2for $e^{*}=(r_{i},r_{j})\in R(L(H))\setminus R(T_{r}(H))$ do
 // where $0pt(T_{r}(H),r_{i})\leq 0pt(T_{r}(H),r_{j})$
3
4 $l\leftarrow\mathsf{LCA}(r_{i},r_{j});d\leftarrow 0pt(l)+1$
5 $e_{1}=(l,r_{1}\leftarrow\mathsf{LA}(r_{i},d));w_{1}=\omega(e_{1})$
6 $e_{2}=(l,r_{2}\leftarrow\mathsf{LA}(r_{j},d));w_{2}=\omega(e_{2})$
7 if $e_{1}=e_{2}$ then
8      slide $r_{j}$ to $r_{1}$
9
10 else if $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}w_{1}=w_{2}$ then
11      slide $r_{i}$ to $r_{1}$ and $r_{j}$ to $r_{2}$
12
13 else if $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}w_{1}<w_{2}$ then
14      slide $r_{i}$ to $l$ and $r_{j}$ to $r_{1}$
15
16 else if $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}w_{2}<w_{1}$ then
17      slide $r_{i}$ to $l$ and $r_{j}$ to $r_{2}$
18
19return $G^{\equiv}$
Algorithm 8 buildEGbuildEG for the $\gamma$-acyclic

### v2

@@
 In particular, while avoiding Cartesian products requires each subplan to form a spanning tree of the corresponding subquery’s join graph, [^1] The join graph of a query has a vertex for each relation and an edge for each pair of relations that join with each other.
-We later define this as the line graph of the query hypergraph in definition 2.
+We later define this as the line graph of the query hypergraph in Section 3.
 our algorithms find maximum spanning trees of the weighted join graph.
@@
 Example 3.
-fig.˜2(a) shows an example hypergraph of size 14.
-The line graph $L_{6}=L(H_{6})$ is shown in fig.˜2(b).
+Figure˜2(a) shows an example hypergraph of size 14.
+The line graph $L_{6}=L(H_{6})$ is shown in Figure˜2(b).
 For instance, hyperedges $Y$ and $U$ share two vertices $a$ and $d$, so they are connected by an edge in the line graph.
@@
 The GYO reduction algorithm finds such an order iteratively, and attaches each hyperedge to its parent to form a join tree.
-It generates the join tree $T^{G}$ as shown in fig.˜2(c).
+It generates the join tree $T^{G}$ as shown in Figure˜2(c).
 Readers may refer to the work of Yu and Ozsoyoglu [45] for details.
 Four common notions of hypergraph acyclicity are defined in decreasing order of strictness, following the seminal work by Fagin [13], namely Berge-acyclic $\Rightarrow$ $\gamma$-acyclic $\Rightarrow$ $\beta$-acyclic $\Rightarrow$ $\alpha$-acyclic.
-The detailed definitions are given by definition˜41 in appendix˜A.
+The detailed definitions are given by Appendix˜A in Appendix˜A.
 Another algorithm for checking acyclicity of hypergraphs and constructing join trees is the Maximum Cardinality Search (MCS) algorithm by Tarjan and Yannakakis [40].
 Algorithm˜1 shows a simplified pseudocode of MCS.
-We discuss the properties of MCS and how it constructs the join trees $T^{M}$ in fig.˜2(d) with example˜42 in appendix˜A.
+We discuss the properties of MCS and how it constructs the join trees $T^{M}$ in Figure˜2(d) with Appendix˜A in Appendix˜A.
 Computation model.
@@
 (d) $T^{\prime}_{W}$
-Figure 2: $T_{W}$ is an MST of the line graph in fig.˜2(b) where the black solid lines stand for the tree edges, the dashed line for an MST edge $(S,U)$, the dotted line for a non-MST edge $(T,Y)$ and gray solid lines for the remaining non-tree edges.
+Figure 2: $T_{W}$ is an MST of the line graph in Figure˜2(b) where the black solid lines stand for the tree edges, the dashed line for an MST edge $(S,U)$, the dotted line for a non-MST edge $(T,Y)$ and gray solid lines for the remaining non-tree edges.
 $G^{\equiv}$ is the resulting equivalent graph where the thick lines highlight the parallel edges.
@@
 Given weighted graph $G$ with MST $T$ rooted, an equivalent graph $G^{\equiv}$ is obtained from $G$ by exhaustively sliding every edge $e=(v,w)\in E(G)$ along another $e^{*}=(u,v)\in T$ as long as $0pt(T,u)<0pt(T,v)$.
-Sliding transformation preserves the identity of each edge $e\in E(G)$ as explained in example˜10. Every spanning tree of $G^{\equiv}$ is an MST of $G$ and vice versa.
+Sliding transformation preserves the identity of each edge $e\in E(G)$ as explained in Section˜4.1. Every spanning tree of $G^{\equiv}$ is an MST of $G$ and vice versa.
 Theorem 9 (by Eppstein [12]).
@@
 Example 10.
-Given the 6-clique line graph $L_{6}$ in fig.˜2(b), we find a rooted MST $T_{W}$ as shown in fig.˜2(a).
+Given the 6-clique line graph $L_{6}$ in Figure˜2(b), we find a rooted MST $T_{W}$ as shown in Figure˜2(a).
 Among all the tree edges shown as black solid lines, we can only apply the sliding transformation to the edge $(T,P)$ whose weight $\omega(T,P)=|\{a\}|=1$ is lighter than its parent tree edge $(W,T)$ with $\omega(W,T)=|\{a,c\}|=2$.
-We slide along the tree edge $(W,T)$ to the root so that the edge $(T,P)$ becomes $(W,P)$ as shown by the solid curve in fig.˜2(b).
+We slide along the tree edge $(W,T)$ to the root so that the edge $(T,P)$ becomes $(W,P)$ as shown by the solid curve in Figure˜2(b).
 Non-tree edges can slide similarly.
-We consider two examples, $(S,U)$ illustrated with a dashed line and $(Y,T)$ with a dotted line in fig.˜2(a).
+We consider two examples, $(S,U)$ illustrated with a dashed line and $(Y,T)$ with a dotted line in Figure˜2(a).
 All other non-tree edges are shown in light gray.
-We can slide along tree edge $(U,W)$ to the root so that the edge $(S,U)$ becomes $(S,W)$ as shown by the dashed curve in fig.˜2(b).
-Both ends of $(Y,T)$ can slide along the tree edges to the root so that the edge becomes a self-loop as shown by the dotted loop in fig.˜2(b).
+We can slide along tree edge $(U,W)$ to the root so that the edge $(S,U)$ becomes $(S,W)$ as shown by the dashed curve in Figure˜2(b).
+Both ends of $(Y,T)$ can slide along the tree edges to the root so that the edge becomes a self-loop as shown by the dotted loop in Figure˜2(b).
 This edge will not appear in any spanning tree of $G^{\equiv}$, and therefore not a part of any MST of $L_{6}$.
 We refer to such an edge as a non-MST edge, as opposed to an MST edge.
-By applying sliding transformations to a fixpoint, we obtain an equivalent graph $G^{\equiv}$, where there are two sets of parallel edges highlighted by thick lines in fig.˜2(c).
+By applying sliding transformations to a fixpoint, we obtain an equivalent graph $G^{\equiv}$, where there are two sets of parallel edges highlighted by thick lines in Figure˜2(c).
 For example, the tree edge $(T,P)$ in $T_{W}$ and non-tree edges $(P,U),(P,T),(P,Y)$ are parallel in $G^{\equiv}$ between $P$ and $W$.
-We can easily verify that each spanning tree of $G^{\equiv}$ corresponds to an MST of $L_{6}$, such as $T^{\prime}_{W}$ in fig.˜2(d).
+We can easily verify that each spanning tree of $G^{\equiv}$ corresponds to an MST of $L_{6}$, such as $T^{\prime}_{W}$ in Figure˜2(d).
 The choice of the initial spanning tree can affect the structure of the equivalent graph, but the order of sliding transformations performed has no impact [12].
@@
 Example 12.
-$H^{*}_{6}$ in fig.˜4(a) is a hypergraph with similar structure to $H_{6}$ in fig.˜2(a).
+$H^{*}_{6}$ in Figure˜4(a) is a hypergraph with similar structure to $H_{6}$ in Figure˜2(a).
 $H^{*}_{6}$ differs from $H_{6}$ by one vertex $d^{\prime}$ and admits the same set of join trees.
 Therefore, finding an equivalent graph $G^{\equiv}(H^{*}_{6})$ is sufficient for enumerating the join trees of $H_{6}$.
-$H^{*}_{6}$ also admits a monotonic weight join tree $\hat{T}_{P}$ in fig.˜4(b).
+$H^{*}_{6}$ also admits a monotonic weight join tree $\hat{T}_{P}$ in Figure˜4(b).
 This means each tree edge in $\hat{T}_{P}$ is already in place, because it cannot slide along the lighter parent.
@@
 To prove the existence, we construct such a hypergraph $H^{*}$ by vertex duplication.
-lemma˜19 guarantees that $\mathcal{T}(H^{*})=\mathcal{T}(H)$.
+Section˜4.2 guarantees that $\mathcal{T}(H^{*})=\mathcal{T}(H)$.
 We only need to show the existence of $\hat{T}\in\mathcal{T}(H^{*})$.
@@
 The duplication is performed only if $\Delta>0$.
-lemma˜16 guarantees that $\chi(e)\setminus\chi(p(e))\neq\emptyset$.
+Section˜4.2 guarantees that $\chi(e)\setminus\chi(p(e))\neq\emptyset$.
 Therefore, we can always duplicate a vertex $x\in\chi(e)\setminus\chi(p(e))$ for $\Delta$ times such that $\omega(e)>\omega(p(e))$.
@@
 Therefore, the duplication of $x$ does not affect the weight of any edge closer to the root than $e$ in $T$.
-By lemma˜17, duplication in a sibling edge $e^{\prime}\in s(e)$ would not affect the weight of $e$ either, and $\omega(e)>\omega(p(e))$ will continue to hold.
+By Section˜4.2, duplication in a sibling edge $e^{\prime}\in s(e)$ would not affect the weight of $e$ either, and $\omega(e)>\omega(p(e))$ will continue to hold.
 The tree has monotonic weights when Algorithm˜2 terminates.
@@
 If $r_{i}$, $r_{j}$ are ancestor and child, then $|\lambda(e)|=1$, otherwise, $|\lambda(e)|=2$.
-In the process of building $G^{\equiv}(H^{*})$ from a monotonic weight join tree $\hat{T}$ of $H^{*}$, definition˜7 and definition˜11 guarantee that:
+In the process of building $G^{\equiv}(H^{*})$ from a monotonic weight join tree $\hat{T}$ of $H^{*}$, Section˜4.1 and Section˜4.2 guarantee that:
   • no tree edge in $\hat{T}$ needs to be slid;
@@
 (c) Equivalent Graph $G^{\equiv}(H^{*}_{6})$
-Figure 4: algorithm˜3 on $H^{*}_{6}$ of fig.˜4(a)
+Figure 4: Algorithm˜3 on $H^{*}_{6}$ of Figure˜4(a)
 Example 21.
-We consider $\hat{T}_{P}$ in fig.˜4(b) of an equivalent hypergraph $H^{*}_{6}$ in fig.˜4(a).
-Three of $\hat{T}_{P}$’s non-tree edges are shown in fig.˜5(a) where the non-MST edges $(Y,W),(Y,T)$ are dotted and the MST edge $(Y,S)$ is dashed.
+We consider $\hat{T}_{P}$ in Figure˜4(b) of an equivalent hypergraph $H^{*}_{6}$ in Figure˜4(a).
+Three of $\hat{T}_{P}$’s non-tree edges are shown in Figure˜5(a) where the non-MST edges $(Y,W),(Y,T)$ are dotted and the MST edge $(Y,S)$ is dashed.
 We can identify $\lambda(Y,W)=\{(T,W),(T,U)\}$ in constant time.
@@
 Because they have the same weight with $\omega(Y,S)=\omega(P,T)=\omega(P,S)=|\{a\}|=1$, we slide $(Y,S)$ directly to $(T,S)$.
-The intermediate graph with those three non-tree edges processed is shown in fig.˜5(b).
-The resulting equivalent graph $G^{\equiv}(H^{*}_{6})$ is shown in fig.˜5(c).
+The intermediate graph with those three non-tree edges processed is shown in Figure˜5(b).
+The resulting equivalent graph $G^{\equiv}(H^{*}_{6})$ is shown in Figure˜5(c).
 The next result follows immediately from the construction of the equivalent hypergraph.
@@
 ∎
-lemma˜22 allows us to produce $G^{\equiv}(H^{*})$ by keeping all tree edges of the MCS tree $T(H)$ in place and sliding the non-tree edges directly to their LCA edges.
-For example, running algorithm˜1 on the MCS tree $T(H_{6})$ in fig.˜2(d) produces an identical $G^{\equiv}(H^{*}_{6})$ in fig.˜5(c).
+Section˜4.2 allows us to produce $G^{\equiv}(H^{*})$ by keeping all tree edges of the MCS tree $T(H)$ in place and sliding the non-tree edges directly to their LCA edges.
+For example, running Algorithm˜1 on the MCS tree $T(H_{6})$ in Figure˜2(d) produces an identical $G^{\equiv}(H^{*}_{6})$ in Figure˜5(c).
 Theorem 23.
@@
 Proof.
-One can readily check that for any non-tree edge $e$ of $T$, the conditions of each if statement in algorithm˜3 hold under $\omega$ if and only if they hold under $\omega^{*}$ by lemma˜22. ∎
+One can readily check that for any non-tree edge $e$ of $T$, the conditions of each if statement in Algorithm˜3 hold under $\omega$ if and only if they hold under $\omega^{*}$ by Section˜4.2. ∎
 Theorem 24.
@@
 An $\alpha$-acyclic hypergraph is Berge-acyclic if and only if it is linear.
-By proposition˜26, every edge in the line graph of a Berge-acyclic hypergraph is labeled with a single variable, thus having a weight of 1.
+By Section˜5, every edge in the line graph of a Berge-acyclic hypergraph is labeled with a single variable, thus having a weight of 1.
 Every spanning tree is a maximum spanning tree, therefore a join tree.
@@
 The line graph $L$ of a Berge-acyclic hypergraph $H$ is a block graph.
-Together with the fact that every block graph is geodetic (it has a unique shortest path between any two vertices) [31], lemma˜31 implies the following corollary.
+Together with the fact that every block graph is geodetic (it has a unique shortest path between any two vertices) [31], Section˜5.1 implies the following corollary.
 Corollary 32.
@@
 Proof.
-A spanning tree of $L$ is its join tree by corollary˜27. To show $T_{r}$ is a join tree, it is sufficient to show $T_{r}$ is spanning, connected and acyclic.
+A spanning tree of $L$ is its join tree by Section˜5. To show $T_{r}$ is a join tree, it is sufficient to show $T_{r}$ is spanning, connected and acyclic.
 $T_{r}$ is spanning and connected because it contains a path from $r$ to each $r^{\prime}\in R(L)$.
@@
 Assuming that $T_{r}|_{R_{d>0}}$ is acyclic, we consider a vertex $r^{\prime}\in R_{d+1}\setminus R_{d}$.
-corollary˜32 guarantees a unique shortest path between each pair of vertices $r,r^{\prime}\in R(L)$.
+Section˜5.1 guarantees a unique shortest path between each pair of vertices $r,r^{\prime}\in R(L)$.
 Each $r^{\prime}$ is connected to a unique neighbor $r^{\prime\prime}\in R_{d}\setminus R_{d-1}$ that is at distance $d$ from $r$.
@@
 The join tree $T_{r}$ is canonical, because the path from $r$ to each $r^{\prime}\in R(L)$ is the shortest and therefore minimizing the depth $0pt(T_{r},r^{\prime})$.
-The canonical tree is unique by corollary˜32.
+The canonical tree is unique by Section˜5.1.
 ∎
@@
 Denote their lowest common ancestor in $T_{r}$ as $r_{a}=\mathsf{LCA}(r_{p},r_{q})$.
-Let $\lambda(r_{p},r_{q})=\{e_{l},e_{r}\}$ as shown by the solid lines in fig.˜5 of appendix˜F, $\mathsf{P}_{p}$ be the path from $r_{a}$ to $r_{p}$ excluding $e_{l}$ and $\mathsf{P}_{q}$ be the path from $r_{a}$ to $r_{q}$ excluding $e_{r}$ as shown by the dashed lines in fig.˜5.
-By lemma˜34, the variables on $\mathsf{P}_{p}$ and $\mathsf{P}_{q}$ are all distinct.
+Let $\lambda(r_{p},r_{q})=\{e_{l},e_{r}\}$ as shown by the solid lines in Figure˜5 of Appendix˜F, $\mathsf{P}_{p}$ be the path from $r_{a}$ to $r_{p}$ excluding $e_{l}$ and $\mathsf{P}_{q}$ be the path from $r_{a}$ to $r_{q}$ excluding $e_{r}$ as shown by the dashed lines in Figure˜5.
+By Section˜5.2, the variables on $\mathsf{P}_{p}$ and $\mathsf{P}_{q}$ are all distinct.
 Consider the following possible cases:
@@
 To prove the “only if” direction by contradiction, we assume that a connected left-deep linear plan for a $\gamma$-acyclic query is not the reverse of a GYO-reduction order.
-By proposition˜38, there is an orphan $\mathring{r}$ in the plan.
+By Section˜6, there is an orphan $\mathring{r}$ in the plan.
 Assuming that $\mathring{r}$ is the $t$-th relation in the plan, we now consider the prefix of the plan up to $\mathring{r}$, namely $\mathsf{pre}({t})=r_{1},\ldots,r_{t-1},\mathring{r}$.
-Let $\hat{r}\in\mathsf{pre}({t-1})$ be a maximal relation with respect to $\geq_{\mathring{r}}$, lemma˜40 guarantees an $\bar{r}\in\mathsf{pre}({t-1})$ incomparable with $\hat{r}$.
+Let $\hat{r}\in\mathsf{pre}({t-1})$ be a maximal relation with respect to $\geq_{\mathring{r}}$, Section˜6 guarantees an $\bar{r}\in\mathsf{pre}({t-1})$ incomparable with $\hat{r}$.
 There are at least two dangling variables $\bar{x}\in\bar{r}\cap\mathring{r}\setminus\hat{r}$ and $\hat{x}\in\hat{r}\cap\mathring{r}\setminus\bar{r}$.
@@
 Can join tree enumeration achieve worst-case constant delay?
-We proved Berge-acyclicity sufficient for the existence and uniqueness of the canonical join tree, but it is not necessary, and $\gamma$-acyclicity is insufficient as shown in fig.˜6 of appendix˜F.
+We proved Berge-acyclicity sufficient for the existence and uniqueness of the canonical join tree, but it is not necessary, and $\gamma$-acyclicity is insufficient as shown in Figure˜6 of Appendix˜F.
 What is the precise characterization of hypergraphs that admit a unique canonical join tree for any root, or for some root?
@@
 Example 42.
-Considering the hypergraph $H_{6}$ shown in fig.˜2(a), we run algorithm˜1 from the hyperedge $P$.
+Considering the hypergraph $H_{6}$ shown in Figure˜2(a), we run Algorithm˜1 from the hyperedge $P$.
 It is labeled as $r_{1}$ and all its vertices are marked.
@@
 Since the labeling of $T$ marks a vertex $c$ in $U,W$, their parents are updated to $p(U)=p(W)=r_{3}$.
-We continue the process until all hyperedges are labeled and all vertices are marked. By connecting each labeled hyperedge with its parent, we construct the join tree $T^{M}$ shown in fig.˜2(d).
+We continue the process until all hyperedges are labeled and all vertices are marked. By connecting each labeled hyperedge with its parent, we construct the join tree $T^{M}$ shown in Figure˜2(d).
 Notice that Algorithm˜1 can generate different join trees for the same hypergraph by breaking ties differently.
 However, it does not guarantee to generate all possible trees.
-For example, it never generates the Join Tree $T^{G}$ shown in fig.˜2(c).
+For example, it never generates the Join Tree $T^{G}$ shown in Figure˜2(c).
 Appendix B Data Structures for Constant Time Operations on Graphs
@@
 Proof.
-Given a non-tree edge $e=(r_{i},r_{j})$ we find $\mathsf{LCA}(r_{i},r_{j})=l$ in constant time by ˜43.
+Given a non-tree edge $e=(r_{i},r_{j})$ we find $\mathsf{LCA}(r_{i},r_{j})=l$ in constant time by Appendix˜B.
 The depth $d=0pt(l)+1$ is likewise found in constant time.
@@
 As we iterate over each edge $(j,\omega)$ of $L[i]$, we check if the current edge $(i,j)$ is a tree edge by performing constant-time query of $T_{r}[i]$ and $T_{r}[j]$ to see if one of them is the parent of the other.
-From ˜44 we know that we can find the LCA edges $\lambda(i,j)=\{e_{1},e_{2}\}$ in constant time.
+From Appendix˜B we know that we can find the LCA edges $\lambda(i,j)=\{e_{1},e_{2}\}$ in constant time.
 There’s no general solution to lookup an entry in a linked list in constant time.
@@
 Appendix D Missing Proofs in Section˜4
-lemma˜15 If $H^{\prime}\twoheadrightarrow H$ then $\mathcal{T}(H)\subseteq\mathcal{T}(H^{\prime})$.
+Section˜4.2 If $H^{\prime}\twoheadrightarrow H$ then $\mathcal{T}(H)\subseteq\mathcal{T}(H^{\prime})$.
 Proof.
@@
 ∎
-lemma˜16 Let $T$ be an MCS tree. For any edge $e\in T$ that has a parent, then $e\not\subseteq p(e)$.
+Section˜4.2 Let $T$ be an MCS tree. For any edge $e\in T$ that has a parent, then $e\not\subseteq p(e)$.
 Proof.
@@
 ∎
-lemma˜17 Let $T$ be an MCS tree. For two edges $e,e^{\prime}\in T$ that have parents,
+Section˜4.2 Let $T$ be an MCS tree. For two edges $e,e^{\prime}\in T$ that have parents,
 $(e\setminus p(e))\ \cap\ (e^{\prime}\setminus p(e^{\prime}))\ \neq\ \emptyset\quad\implies\quad p(e)=p(e^{\prime}).$
@@
 If $(e\setminus p(e))\cap(e^{\prime}\setminus e^{\prime\prime})\neq\emptyset$, then $\exists\ x\in(e^{\prime}\cap e)\setminus e^{\prime\prime}$, which violates the running intersection property of join trees.
-lemma˜19 Let $H^{\prime}$ be a duplicated hypergraph of $H$. Then,
+Section˜4.2 Let $H^{\prime}$ be a duplicated hypergraph of $H$. Then,
   1. 1.
@@
 Likewise, the pair of functions $(X(H)_{\texttt{id}},R(H)_{\texttt{id}})$ constitutes a homomorphism $H\twoheadrightarrow H^{\prime}$.
-2. It follows from lemma˜15 and the facts above.
+2. It follows from Section˜4.2 and the facts above.
 3.
@@
 Appendix E Missing Proofs in Section˜5
-proposition˜26 An $\alpha$-acyclic hypergraph is Berge-acyclic if and only if it is linear.
+Section˜5 An $\alpha$-acyclic hypergraph is Berge-acyclic if and only if it is linear.
 Proof.
@@
 ∎
-lemma˜31 The line graph $L$ of a Berge-acyclic hypergraph $H$ is a block graph.
+Section˜5.1 The line graph $L$ of a Berge-acyclic hypergraph $H$ is a block graph.
 Proof.
@@
 ∎
-lemma˜34 Let $T_{r}$ be the canonical join tree rooted at $r$ of a Berge-acyclic hypergraph.
+Section˜5.2 Let $T_{r}$ be the canonical join tree rooted at $r$ of a Berge-acyclic hypergraph.
 Along its root-to-leaf path $r,r_{1},\ldots,r_{k}$, each pair of adjacent vertices shares a distinct variable.
 Proof.
-Let $r_{0}=r$, proposition˜26 requires that $|r_{i-1}\cap r_{i}|=1$ for all $i\in[k]$.
+Let $r_{0}=r$, Section˜5 requires that $|r_{i-1}\cap r_{i}|=1$ for all $i\in[k]$.
 By Theorem˜33, the path $r_{0},\ldots,r_{k}$ is the shortest path from $r_{0}$ to $r_{k}$.

### v3

@@
 We propose three approaches to constructing join trees for acyclic queries.
-First, we give an algorithm to enumerate all join trees of an $\alpha$-acyclic query by edits with amortized constant delay, which forms the basis of a cost-based optimizer for acyclic joins.
-Second, we show that the Maximum Cardinality Search algorithm by Tarjan and Yannakakis constructs a unique shallowest join tree, rooted at any relation, for a Berge-acyclic query; this tree enables parallel execution of large join queries.
-Finally, we prove that any connected left-deep linear plan for a $\gamma$-acyclic query can be converted into a join tree by a simple algorithm, allowing reuse of optimization infrastructure developed for binary joins.
+First, we give an algorithm to enumerate all join trees of an $\alpha$-acyclic query by edits in linear time with amortized constant delay, which forms the basis of a cost-based optimizer for acyclic joins.
+Second, we show the Maximum Cardinality Search algorithm by Tarjan and Yannakakis constructs the unique shallowest join tree for any Berge-acyclic query, thus enabling parallel execution of large join queries.
+Finally, we prove that a simple algorithm by Hu et al.
+converts any connected left-deep linear plan of a $\gamma$-acyclic query into a join tree, allowing reuse of optimizers developed for binary joins.
 1 Introduction
 The query optimizer sits at the heart of a database system.
-It takes a query as input and generates a plan for efficient execution, allowing users to program declaratively without worrying about evaluation.
+It takes a query as input and generates a plan for efficient execution, allowing the user to program declaratively without worrying about performance.
 Among the many relational algebra operators, join has received significant attention in optimization research.
@@
 A query optimizer typically has two parts: a plan generator and a cost model to assess each plan. This paper focuses on plan generation and presents three approaches:
-  • An algorithm that enumerates all join trees of an acyclic query by edits, with amortized constant time delay.
-  • A proof that Maximum Cardinality Search [40] yields a unique shallowest join tree for any Berge-acyclic query, enabling parallel execution for very large queries.
-  • A characterization of when the algorithm by Hu et al. [20] can convert a binary join plan into a valid join tree, allowing reuse of existing optimizers.
-The rest of the paper is organized as follows: Section˜2 discusses related work; Section˜3 introduces relevant concepts and notations; Section˜4 presents the join tree enumeration algorithm; Section˜5 introduces canonical join trees and their construction; Section˜6 discusses the conversion from binary plans to join trees; finally, Section˜7 points to avenues for future work and concludes the paper.
-All missing proofs can be found in the appendix, which also contains additional figures, tables and algorithms.
+  1. 1.
+    Given the line graph $L$ (section˜3) of an $\alpha$-acyclic query $H$, algorithm˜2 enumerates all join trees by edits [^1] To avoid redundant work, enumeration by edits [24] outputs the difference between consecutive elements.
+with amortized constant delay, i.e., in $\mathcal{O}(|L|+k)$ where $k$ is the number of trees generated (Theorem˜27).
+If the query is $\gamma$-acyclic, algorithm˜4 further reduces the total time complexity to $\mathcal{O}(|H|+k)$ (Theorem˜28).
+  2. 2.
+    Given a Berge-acyclic query, we prove that the classic Maximum Cardinality Search algorithm by Tarjan and Yannakakis [41] yields a unique shallowest join tree (Theorem˜35), enabling parallel execution of very large queries.
+  3. 3.
+    Given a connected left-deep linear join plan for any $\gamma$-acyclic query, we prove that a simple algorithm by Hu et al.
+[22] always converts the plan into a valid join tree (section˜6), allowing reuse of existing optimizers.
+The rest of the paper is organized as follows: section˜2 discusses related work; section˜3 introduces relevant concepts and notations; section˜4 presents the join tree enumeration algorithms; section˜5 introduces the unique shallowest join tree, namely the canonical join tree and its construction; section˜6 discusses the conversion from binary plans to join trees; section˜7 concludes the paper by pointing to avenues for future work.
+Additional technical details, proofs and a notation table (table˜2) can be found in the appendix.
 2 Related Work
-Join order optimization is well studied, with algorithms based on dynamic programming (DP) from the bottom up [28, 34, 30], cost-based pruning from the top down [11, 15], greedy heuristics [5, 14, 38], and randomized search [36].
+Join order optimization is well studied, with algorithms based on dynamic programming (DP) from the bottom up [30, 36, 32], cost-based rewriting from the top down [13, 17], greedy heuristics [6, 16, 40], and randomized search [38].
 Since the plan space is exponential, most methods prune it: some restrict to left-deep plans [23, 25], while others avoid Cartesian products [30, 31].
@@
 This is desirable because subplans are grouped into equivalence classes, and the optimal plan can be constructed in a bottom-up manner.
-Our enumeration algorithm avoids redundant work by enumerating join trees by edits, i.e., it outputs the difference between consecutive join trees instead of the trees themselves.
 Future work may explore constructing compact representations of join trees, suitable for dynamic programming.
@@
 3 Preliminaries
-We focus on full conjunctive queries [1] in this paper and identify each query with its hypergraph, where there is a vertex for each variable and a hyperedge for each relation.
+We focus on full conjunctive queries [1] in this paper and identify each query with its hypergraph, where each vertex represents a variable and each hyperedge represents a relation.
+Definition 1 (Hypergraph).
+A hypergraph $H=(X,R,\chi)$ consists of a set of vertices $X$, a set of hyperedges $R$ and an incidence function $\chi:R\to\mathcal{P}(X)$.
+We only consider hypergraphs without isolated vertices, empty hyperedges or duplicated hyperedges over the same set of vertices.
+We assume each hyperedge $r$ contains a bounded number ($\mathcal{O}(1)$) of vertices exclusive to $r$.
+When there is no ambiguity, we will use $r$ interchangeably with $\chi(r)$, $x\in r$ with $x\in\chi(r)$, and $H(X,R)$ with $H(X,R,\chi)$.
+We will also apply common set operations directly to hyperedges, e.g., $r_{1}\cap r_{2}$ for $\chi(r_{1})\cap\chi(r_{2})$.
+In fact, the reader can often ignore $\chi$ and identify a hyperedge with its set of vertices; the purpose of $\chi$ is to relate hyperedges over different vertex sets across hypergraphs in the correctness proofs of our algorithms.
+We write $X(H)$ and $R(H)$ to refer to the vertices and edges of $H$ respectively.
+The size of a hypergraph is the total size of all hyperedges $|H|=\sum_{r\in R}|r|$.
+We let $H|_{x}$ denote the neighborhood of $x\in X$, consisting of all hyperedges containing $x$.
+Definition 2 (Multigraph).
+A multigraph $G=(R,E,\rho)$ consists of a set of vertices $R$, a set of edges $E$, and the incidence function $\rho:E\to\mathcal{P}(R)$ such that $1\leq|\rho(e)|\leq 2$ for any $e\in E$.
+An edge $e$ is a self-loop if $|\rho(e)|=1$.
+Edges $e_{1},e_{2}\in E$ are parallel if $\rho(e_{1})=\rho(e_{2})$.
+We denote the set of vertices in $G$ with $R$, because we will soon define the line graph $L$ of a hypergraph where each vertex in $L$ represents a hyperedge.
+We write $R(G)$ and $E(G)$ to refer to the vertices and edges of $G$ respectively.
+Definition 3 (Simple Graph).
+A simple graph is a multigraph with no parallel edges or self-loops, i.e., the incidence function is injective and always returns two distinct vertices.
+Because each edge $e$ in a simple graph can be identified with its two endpoints $\rho(e)=\{r_{1},r_{2}\}$, we will omit $\rho$ and write $\{r_{1},r_{2}\}$ for $e$ when there is no ambiguity.
+Definition 4 (Cycle, Clique and Diamond).
+In a simple graph $G=(R,E)$, a sequence of $n\geq 3$ distinct vertices $(v_{0},\ldots,v_{n-1})$ where $v_{i}\in R$ is
+  • a cycle if there is an edge between $v_{i}$ and $v_{(i+1)\bmod n}$ for all $0\leq i\leq n-1$;
+  • an n-clique $K_{n}$ if the induced graph $G|_{\{v_{0},\dots,v_{n}\}}$ has an edge between each pair of vertices;
+  • a diamond if $n=4$ and $G|_{\{v_{0},\dots,v_{3}\}}$ has one fewer edge than a $K_{4}$.
+A weighted graph is a (multi-)graph where each edge is assigned a weight:
+Definition 5 (Weighted Graph).
+A weighted graph $(G,\omega)$ consists of a multigraph $G=(R,E,\rho)$ and weight function $\omega:E\to\mathbb{N}$ assigning a natural number to each edge in $G$.
+The line graph of a hypergraph is the “intersection graph” of its hyperedges:
+Definition 6 (Line Graph).
+For a hypergraph $H$, the line graph $L(H)=(G,\chi)$ consists of a simple graph $G=(R,E)$ and an edge labeling function $\chi:E\to\mathcal{P}(X)$.
+The vertices of $G$ are the hyperedges $R$ of $H$, and there is an edge $e=\{r_{1},r_{2}\}\in E$ when $r_{1}\cap r_{2}\neq\emptyset$.
+The edge labeling function maps each edge $e$ to the intersection of the hyperedges represented by the endpoints of $e$: $\chi(\{r_{1},r_{2}\})=r_{1}\cap r_{2}$.
+In addition, we define a weight function mapping each edge to the size of its label:
+Definition 7 (Line Graph Edge Weight).
+For each edge $e\in E(L(H))$, the weight function $\omega:E(L(H))\to\mathbb{N}$ is defined as $\omega(e)=|\chi(e)|$.
+When there is no ambiguity from the context, we simply write $L(H)$ as $L$, and we may treat the line graph $(G,\chi)$ as a weighted graph $(G,\omega)$.
+The size of a line graph is the sum of all edge weights $|L|=\sum_{e\in E(L)}\omega(e)$.
+For any subgraph $G$ of $L(H)$, we let $G|_{x}$ denote the subgraph of $G$ induced by $H|_{x}$.
+Example 8.
+fig.˜2(a) shows an example hypergraph of size 14.
+The line graph $L_{6}=L(H_{6})$ is shown in fig.˜2(b).
+For instance, hyperedges $Y$ and $U$ share two vertices $a$ and $d$, so they are connected by an edge in the line graph.
+The edge weight is given by $\omega(\{Y,U\})=|\{a,d\}|=2$.
+Each unannotated edge in $L_{6}$ connects two hyperedges sharing only the vertex $a$.
+Otherwise, the common vertices shared by a pair of hyperedges are annotated next to the corresponding edge.
+The size of the line graph is $|L_{6}|=19$.
+Definition 9 (Join Tree).
+A join tree $T$ of hypergraph $H$ is a spanning tree of $L(H)$ such that $T|_{x}$ is a connected subtree for each $x\in X(H)$.
+If a certain vertex is specified as the root, $T$ becomes a rooted join tree.
+We use $R(T)$ to denote the set of nodes and $E(T)$ to denote the set of edges of tree $T$.
+The requirement that $T|_{x\in X(H)}$ be a connected subtree is also known as the running intersection property [15].
+We write $\mathcal{T}(L(H))$ or $\mathcal{T}(H)$ to denote the set of unrooted join trees of $H$.
+We use $T_{r}$ to denote a join tree rooted at $r\in R(T)$.
+When there is no ambiguity, we simply write $T$.
+The depth of a node $r_{i}$ in the rooted tree $T_{r}$, denoted $0pt(T_{r},r_{i})$, is defined as its distance from the root.
+Four common notions of hypergraph acyclicity are defined in increasing order of strictness [15], namely $\alpha$-acyclic $\supset$ $\beta$-acyclic $\supset$ $\gamma$-acyclic $\supset$ Berge-acyclic.
+Definition 10 (Hypergraph Acyclicity).
+A hypergraph $H$ is:
+  • $\alpha$-acyclic if it admits a join tree as in section˜3;
+  • $\beta$-acyclic if every subgraph of $H$ is $\alpha$-acyclic;
+  • $\gamma$-acyclic if it does not contain any $\gamma$ cycle.
+A $\gamma$ cycle is a sequence of length $k\geq 3$ of distinct hyperedges and distinct vertices $(r_{0},x_{0},\dots,r_{k-1},x_{k-1})$ such that every $x_{i\in[0,k-2]}$ belongs to $r_{i}\cap r_{i+1}$ and no other $r_{j}$ while $x_{k-1}$ belongs to $r_{k-1}\cap r_{0}$ and possibly other hyperedges;
+  • Berge-acyclic if it does not contain any Berge cycle.
+A Berge cycle is a sequence of length $k\geq 2$ of distinct vertices and distinct hyperedges $(r_{0},x_{0},\dots,r_{k-1},x_{k-1})$ such that $\forall\ i\in[k]:x_{i}\in r_{i}\cap r_{(i+1)\bmod k}$.
+The following classic result relates join trees of $H$ to maximum spanning trees of $L(H)$:
+Theorem 11 (Maier [29]).
+Given an $\alpha$-acyclic hypergraph $H=(X,R)$, a tree with nodes in $R$ is a join tree of $H$ if and only if it is a maximum spanning tree of $L(H)$.
+Join trees for acyclic hypergraphs can be constructed by a procedure called GYO reduction.
 Definition 12.
-A hypergraph is a tuple $(X,R,\chi)$ where $X$ is a set of vertices, $R$ a set of hyperedges, and $\chi:R\to 2^{X}$ is a function that assigns to each hyperedge a set of vertices.
-Without loss of generality, we limit the scope of our discussion to queries which admit connected hypergraphs.
-For queries with disconnected hypergraphs, we treat each component separately.
-We do not consider isolated vertices, empty hyperedges or duplicated hyperedges containing the same set of vertices.
-Therefore, each hyperedge can be identified with its vertex set.
-We will use $x\in r$ interchangeably with $x\in\chi(r)$ and apply the common set operations to hyperedges.
-The size of a hypergraph is the sum of all hyperedges $|H|=\sum_{r\in R}|r|$.
-We let $H|_{x}$ denote the neighborhood of $x$ in $H$, consisting of all hyperedges containing vertex $x$.
-Input: Hypergraph $H$, root $r$
+A GYO reduction order is a sequence of hyperedges $r_{1},r_{2},\ldots,r_{k}$ such that for each $r_{i<k}$, there is some $r_{p>i}$, called the parent of $r_{i}$, such that $\forall r_{j>i}:r_{i}\cap r_{j}\subseteq r_{p}$.
+The GYO reduction algorithm [46] finds such an order iteratively, and attaches each hyperedge to its parent to form a join tree.
+It generates the join tree $T^{G}$ as shown in fig.˜2(c), where, for example, $T$ is the parent of $W$, denoted as $p(W)=T$ and the root $P$ has no parent.
+Input: $\alpha$-acyclic hypergraph $H$ and root $r$
 Output: Join tree $T_{r}$ rooted at $r$
@@
 […]
-[…]
-4while $R\neq\emptyset$ do
-5 $i\leftarrow i+1$
-6 $r_{i}\leftarrow r$
-7 for $x\in r_{i}\cap X$ do
+3while $R\neq\emptyset$ do
+4
+5 for $x\in r\cap X$ do
 6    for $r^{\prime}\in R:x\in r^{\prime}$ do
-9       $p(r^{\prime})\leftarrow r_{i}$
-10
-11
-12 $X\leftarrow X\setminus r_{i}$
-13 $R\leftarrow R\setminus\{r_{i}\}$
+7       $p(r^{\prime})\leftarrow r$
+8
+9
+10 $X\leftarrow X\setminus r$
+11 $R\leftarrow R\setminus\{r\}$
+12 $E(T_{r})\leftarrow E(T_{r})\cup\{\{r,p(r)\}\}$
 13 $r\leftarrow\operatorname*{arg\,max}_{r\in R}|r\setminus X|$
+return $T_{r}$
 Algorithm 1 MCS from $r\in R(H)$
@@
 Figure 1: A hypergraph $H_{6}$, its line graph $L_{6}$, and two join trees $T^{G}$ and $T^{M}$.
-Definition 2.
-The line graph $L(H)$ of a hypergraph $H=(X,R,\chi)$ is an undirected simple graph $(R,E)$ whose vertices $R$ are the hyperedges of $H$, and there is an edge $(r_{1},r_{2})\in E$ whenever $r_{1}\cap r_{2}\neq\emptyset$.
-We extend $\chi$ such that $\chi(e)=\chi(r_{1},r_{2}):=r_{1}\cap r_{2}$ for each edge $e=(r_{1},r_{2})\in E$, and define the edge-weight function $\omega(r_{1},r_{2}):=|r_{1}\cap r_{2}|$.
-When there is no ambiguity from the context, we simply write $L(H)$ as $L$.
-The size of a line graph is the sum of all edge weights $|L|=\sum_{e\in E(L)}\omega(e)$.
-It reflects the number of join clauses in a SQL query with one clause per variable shared between two relations.
-We let $L|_{x}$ denote the subgraph of $L(H)$ induced by $H|_{x}$.
+Another way to construct join trees is via the Maximum Cardinality Search (MCS) algorithm [41], and we present a simplified version in algorithm˜1.
+The join tree produced by the algorithm is called an MCS tree.
+Given an input hypergraph $H$ and a hyperedge $r\in R(H)$ chosen as the root, algorithm˜1 constructs a rooted MCS tree by iteratively expanding toward hyperedges connected with those already processed.
+The algorithm maintains a set $R$ of unprocessed hyperedges and a set $X$ of unvisited vertices.
+Initially, the MCS tree contains no edges, and the first hyperedge to be processed is the root $r$.
+In each iteration, the algorithm processes the current hyperedge $r$ by examining all of its unvisited vertices in $r\cap X$.
+Whenever such a vertex $x$ is encountered, every remaining hyperedge $r^{\prime}\in R$ that contains $x$ records $r$ as its tentative parent $p(r^{\prime})\leftarrow r$.
+Once all vertices of $r$ are marked as visited, they are removed from $X$.
+The hyperedge $r$ is marked as processed and removed from $R$.
+After processing each $r$, the edge between $r$ and its recorded parent $p(r)$ is added to the MCS tree as $\{r,p(r)\}$.
+The algorithm selects as the next hyperedge an element of $R$ that maximizes the number of already visited vertices, namely a hyperedge $r$ that maximizes $|r\setminus X|$, breaking ties arbitrarily.
+This process continues until all hyperedges have been processed, when the recorded edges $\{r,p(r)\}$ form an MCS tree rooted at $r$.
 Example 13.
-Figure˜2(a) shows an example hypergraph of size 14.
-The line graph $L_{6}=L(H_{6})$ is shown in Figure˜2(b).
-For instance, hyperedges $Y$ and $U$ share two vertices $a$ and $d$, so they are connected by an edge in the line graph.
-The edge weight is given by $\omega(Y,U)=|\{a,d\}|=2$.
-Each unannotated edge in $L_{6}$ connects two hyperedges sharing only the vertex $a$.
-Otherwise, the common vertices shared by a pair of hyperedges are annotated next to the corresponding edge.
-The size of the line graph is $|L_{6}|=19$.
-Definition 4.
-A join tree $T$ of hypergraph $H$ is a spanning tree of $L(H)$ such that $T|_{x}$ is a connected subtree for each $x\in X(H)$.
-If a certain vertex is specified as the root, $T$ becomes a rooted join tree where the edges are oriented away from the root.
-A join tree $T(L)$ is a subgraph of $L$, so we use $R(T)$ to denote the set of tree nodes and $E(T)$ to denote the set of tree edges.
-Because $T$ is spanning and acyclic, we always have $R(T)\equiv R(L)$ and $E(T)\subseteq E(L)$.
-The requirement that $T|_{x\in X(H)}$ is connected is also known as the running intersection property [13].
-We write $\mathcal{T}(L(H))$ or $\mathcal{T}(H)$ to denote the set of unrooted join trees.
-The union join graph $\mathcal{U}(L(H))$ or $\mathcal{U}(H)$ is a spanning subgraph of $L(H)$ where the edge set is the union of all join trees $E(\mathcal{U}(H))=\bigcup_{T\in\mathcal{T}(H)}E(T)$.
-We use $T_{r}$ to denote a join tree rooted at $r\in R(T)$.
-When there is no ambiguity, we simply write $T$.
-The depth of a node $r_{i}$ in the rooted tree $T_{r}$, denoted $0pt(T_{r},r_{i})$, is defined as its distance from the root.
-Join trees can be constructed by a procedure called GYO reduction.
-Definition 5.
-A GYO reduction order is a sequence of hyperedges $r_{1},r_{2},\ldots,r_{k}$ such that for each $r_{i<k}$, there is some $r_{p>i}$, called the parent of $r_{i}$, such that $\forall r_{j>i}:r_{i}\cap r_{j}\subseteq r_{p}$.
-The GYO reduction algorithm finds such an order iteratively, and attaches each hyperedge to its parent to form a join tree.
-It generates the join tree $T^{G}$ as shown in Figure˜2(c).
-Readers may refer to the work of Yu and Ozsoyoglu [45] for details.
-Four common notions of hypergraph acyclicity are defined in decreasing order of strictness, following the seminal work by Fagin [13], namely Berge-acyclic $\Rightarrow$ $\gamma$-acyclic $\Rightarrow$ $\beta$-acyclic $\Rightarrow$ $\alpha$-acyclic.
-The detailed definitions are given by Appendix˜A in Appendix˜A.
-Another algorithm for checking acyclicity of hypergraphs and constructing join trees is the Maximum Cardinality Search (MCS) algorithm by Tarjan and Yannakakis [40].
-Algorithm˜1 shows a simplified pseudocode of MCS.
-We discuss the properties of MCS and how it constructs the join trees $T^{M}$ in Figure˜2(d) with Appendix˜A in Appendix˜A.
+On $H_{6}$ in fig.˜2(a) with $r=P$, algorithm˜1 examines all vertices of $P$.
+$P$ also becomes the tentative parent of five hyperedges, $p(S)=p(T)=p(U)=p(W)=p(Y)\leftarrow P$.
+It proceeds to check the next hyperedge with the most visited vertices as in Line 1.
+At this point, each of the five has one visited vertex, algorithm˜1 breaks the tie arbitrarily, say proceeding with $S$, and examining all its vertices.
+Its parent is finalized as $p(S)=P$.
+Each of the remaining four again has one visited vertex, so algorithm˜1 proceeds to check $T$.
+Its parent is finalized as $p(T)=P$.
+While checking $T$, algorithm˜1 examines the vertex $c$ that is also in $U,W$, their parents are updated to $p(U)=p(W)\leftarrow T$.
+It continues until all hyperedges are processed.
+By connecting each hyperedge with its parent, algorithm˜1 constructs the MCS tree $T^{M}$ in fig.˜2(d).
+algorithm˜1 can generate different join trees by varying the choice of root and tie-breaking strategy.
+However, it does not guarantee to generate all possible trees.
+For example, it never generates the join tree $T^{G}$ shown in fig.˜2(c).
+We extend the notions of parent, children and siblings to edges of a rooted tree.
+Let $e\in E(T)$ be an edge of tree $T$ rooted at $r$ (and $e$ not incident to $r$).
+Its parent $p(e)\in E(T)$ is the unique tree edge incident to $e$ and lying closer to the root $r$.
+We let $c(e)$ denote the set of children of $e$ in $T$.
+The siblings of $e$ is the set of edges incident to the parent of $e$, namely $s(e)=\left\{e^{\prime}\in E(T)\setminus\{e\}\mid p(e^{\prime})=p(e)\right\}$.
+Edges incident to the root do not have a parent and are all siblings.
+section˜3 shows that an MCS tree is “somewhat monotonic” in that every edge must contain some variable not in its parent. If two edges share any variable not in their parents, they must be siblings.
+Lemma 14.
+Let $T$ be an MCS tree. For an edge $e\in T$ that has a parent, then
+  1. 1.
+    $e\not\subseteq p(e)$
+  2. 2.
+    For another edge $e^{\prime}\in T:(e\setminus p(e))\ \cap\ (e^{\prime}\setminus p(e^{\prime}))\ \neq\ \emptyset\quad\implies\quad p(e)=p(e^{\prime}).$
 Computation model.
@@
 4 Enumerating Join Trees
-Our strategy for enumerating join trees is based on the following classic result relating join trees of a hypergraph to maximum spanning trees (MSTs) of its line graph:
-Proposition 6 (Maier [27]).
-A subgraph of $L(H)$ is a join tree of $H$ if and only if it is a maximum spanning tree of $L(H)$ according to the weight function $\omega$.
-A naïve approach is to apply an off-the-shelf algorithm for enumerating the maximum spanning trees of $L(H)$.
-The best known algorithm for this purpose is due to Eppstein [12], and it works by deriving from the input graph $G$ a so-called equivalent graph $G^{\equiv}$.
+Our strategy for enumerating the join trees of a hypergraph $H$ starts from Theorem˜11 which allows us to reduce the problem to the enumeration of maximum spanning trees of the line graph $L(H)$.
+The best known algorithm for MST enumeration [^3] Eppstein focuses on minimum spanning trees, but the same algorithm applies to maximum spanning trees with flipped comparisons.
+is due to Eppstein [14] by deriving from the input graph $G$ a so-called equivalent graph $G^{\equiv}$.
 Every spanning tree of $G^{\equiv}$ corresponds to an MST of $G$ and vice versa.
-Eppstein proves a lower bound of $\Omega(m+n\log n)$ on constructing the equivalent graph from an arbitrary weighted graph with $m$ edges and $n$ vertices.
-Then to enumerate all $k$ MSTs of $G$, Eppstein applies existing algorithms to enumerate the spanning trees of $G^{\equiv}$.
-Since there are optimal spanning tree enumeration algorithms that run in $\mathcal{O}(m+n+k)$ [22, 35], the overall time complexity to enumerate MSTs is $\mathcal{O}(m+n\log n+k)$.
-The main result of this section is an algorithm for enumerating join trees leveraging the structure of acyclic hypergraphs and their line graphs. In particular:
-  • for any $\alpha$-acyclic hypergraph $H$, we can construct an equivalent graph in $\mathcal{O}(|L|)=\mathcal{O}(m+n)$, where $m$ and $n$ are the numbers of edges and vertices in $L(H)$, thus enumerating the join trees in $\mathcal{O}(|L|+k)$.
-  • for any $\gamma$-acyclic hypergraph $H$, we can construct an equivalent graph from the union join graph $\mathcal{U}(H)$ which can be built in $\mathcal{O}(|H|+|\mathcal{U}(H)|)$, lowering the overall time complexity of enumeration to $\mathcal{O}(|H|+k)$.
-In the rest of this section, we formally introduce the notion of the equivalent graph $G^{\equiv}$.
-Thereafter, we outline the main algorithm for enumerating join trees, prove its correctness and analyze its complexity.
-Finally, we consider the case of $\gamma$-acyclic queries.
+Eppstein gives an algorithm (and proves a matching lower bound) in time $\mathcal{O}(m+n\log n)$ to construct the equivalent graph from an arbitrary weighted graph with $m$ edges and $n$ vertices.
+Then to enumerate all $k$ MSTs of $G$, he applies existing algorithms to enumerate the spanning trees of $G^{\equiv}$.
+Since there are optimal spanning tree enumeration algorithms that run in $\mathcal{O}(m+n+k)$ time [24, 37], the overall time complexity to enumerate MSTs is $\mathcal{O}(m+n\log n+k)$.
+The main result of this section is an algorithm for enumerating join trees leveraging the structure of acyclic hypergraphs and their line graphs.
+In particular:
+  • Given the line graph $L$ of an $\alpha$-acyclic hypergraph, we can construct an equivalent graph of $L$ in $\mathcal{O}(|L|)$, thus enumerating the join trees in $\mathcal{O}(|L|+k)$ time (Theorem˜27).
+  • Given any $\gamma$-acyclic hypergraph $H$, we can construct an equivalent graph of $L(H)$ in $\mathcal{O}(|H|)$, lowering the overall time complexity of enumeration to $\mathcal{O}(|H|+k)$ (Theorem˜28).
+Note that $|L(H)|$ can be quadratic in $|H|$ while $|H|\in\mathcal{O}(|L(H)|)$, so the second item above yields a better bound.
+In the rest of this section, we first define what is an equivalent graph.
+Then, we present the algorithm for enumerating join trees of $\alpha$-acyclic hypergraphs.
+Finally, we adapt the algorithm to enumerate join trees of $\gamma$-acyclic hypergraphs more efficiently.
 4.1 Equivalent Graph
-The construction of the equivalent graph $G^{\equiv}$ for a given edge-weighted graph $G$, is based on the following local graph modification, sliding transformation.
+The key idea of Eppstein’s algorithm [14] is to construct an equivalent graph $G^{\equiv}$ whose spanning trees one-to-one correspond to the MSTs of the input graph $G$, thereby reducing MST enumeration to spanning tree enumeration.
+Definition 15 (Equivalent Graph [14]).
+Given a weighted graph $G=((R,E,\rho),\omega)$, a multigraph $G^{\equiv}=((R,E,\rho^{\equiv}),\omega)$ is an equivalent graph of $G$ if the spanning trees of $G^{\equiv}$ one-to-one correspond to the maximum spanning trees of $G$: a set of edges $E^{T}\subseteq E$ induces a spanning tree of $G^{\equiv}$ if and only if $E^{T}$ induces an MST of $G$.
+In what follows, we may use the notation $G^{\equiv}(L(H))$ or $G^{\equiv}(H)$ to refer to an equivalent graph of the line graph of hypergraph $H$.
+Note that $G^{\equiv}$ shares the same vertices and edges of $G$, and they only differ in the incidence function mapping each edge to its endpoints.
+Such an equivalent graph can be constructed from $G$ by applying a series of sliding transformations: for two incident edges $e,e^{\prime}$ with $\rho(e)=\{r_{1},r_{2}\}$ and $\rho(e^{\prime})=\{r_{2},r_{3}\}$, we can slide $e^{\prime}$ along $e$ by updating $\rho(e^{\prime})=\{r_{1},r_{3}\}$, if $\omega(e^{\prime})<\omega(e)$.
+Note that the edge $e^{\prime}$ retains its identity after sliding, which explains the need for the incidence function $\rho$.
+Furthermore, we will use a given rooted MST $T$ as a guide, and only slide edges towards the root.
+More formally:
 Definition 16 (Sliding Transformation [14]).
-Let $G$ be an edge-weighted graph with weight function $\omega$.
-Given edges $e^{*}=(u,v)$ and $e=(v,w)$ where $\omega(e)<\omega(e^{*})$, sliding $e$ along $e^{*}$ moves $e$’s end node $v$ to $u$ such that $e=(u,w)$.
+Let $T$ be a MST of $G$ rooted at $r$.
+Let $e$ be an edge such that $\rho(e)=\{r_{1},r_{2}\}$ with $r_{1}$ closer to the root than $r_{2}$ in $T$.
+If another edge $e^{\prime}$ shares $r_{2}$ with $e$, i.e., $\rho(e^{\prime})=\{r_{2},r_{3}\}$, and $\omega(e^{\prime})<\omega(e)$, then sliding $e^{\prime}$ along $e$ results in a graph $G^{\prime}$ that is identical to $G$, except that the incidence function maps $e^{\prime}$ to $\{r_{1},r_{3}\}$.
+The key result by Eppstein shows that applying sliding transformations on $G$ along a rooted MST $T$ to a fixpoint results in an equivalent graph of $G$.
+To state this formally, we first define an ordering on graphs based on sliding transformations.
+It is easy to verify the following is a partial order:
+Definition 17 (Sliding Partial Order).
+We write $G\leq_{T}G^{\prime}$ if $G^{\prime}$ can be obtained from $G$ by applying a sequence of sliding transformations along a rooted MST $T$ of $G$.
+Applying sliding transformations to a fixpoint therefore yields a maximal element under the sliding partial order.
+Theorem 18 (Sliding produces $G^{\equiv}$ at fixpoint [14]).
+Given a weighted graph $G$ and a rooted MST $T$, and let $G^{\equiv}$ be a maximal element under $\leq_{T}$. Then $G^{\equiv}$ is an equivalent graph of $G$.
+The choice of the initial rooted MST can affect the structure of the equivalent graph, but the order of sliding transformations performed has no impact [14].
 (a) $T_{W}$
@@
 (d) $T^{\prime}_{W}$
-Figure 2: $T_{W}$ is an MST of the line graph in Figure˜2(b) where the black solid lines stand for the tree edges, the dashed line for an MST edge $(S,U)$, the dotted line for a non-MST edge $(T,Y)$ and gray solid lines for the remaining non-tree edges.
+Figure 2: $T_{W}$ is an MST of the line graph in fig.˜2(b) where the black solid lines stand for the tree edges, the dashed line for an MST edge $(S,U)$, the dotted line for a non-MST edge $(T,Y)$ and gray solid lines for the remaining non-tree edges.
 $G^{\equiv}$ is the resulting equivalent graph where the thick lines highlight the parallel edges.
-We can construct an equivalent graph by repeatedly performing sliding transformations. Sliding may create parallel edges, cycles, and self-loops, so the result is a multigraph.
-Definition 8 (Equivalent Graph [12]).
-Given weighted graph $G$ with MST $T$ rooted, an equivalent graph $G^{\equiv}$ is obtained from $G$ by exhaustively sliding every edge $e=(v,w)\in E(G)$ along another $e^{*}=(u,v)\in T$ as long as $0pt(T,u)<0pt(T,v)$.
-Sliding transformation preserves the identity of each edge $e\in E(G)$ as explained in Section˜4.1. Every spanning tree of $G^{\equiv}$ is an MST of $G$ and vice versa.
-Theorem 9 (by Eppstein [12]).
-The maximum spanning trees of a weighted graph $G$ are in one-to-one correspondence with the spanning trees of its equivalent graph $G^{\equiv}$.
-[^2] Minimum spanning trees are considered in the original literature [12].
-In the following, we use $p(r)$ to denote the parent of $r$ in a given rooted tree, and $c(r)$ to denote its children.
-We extend the notation of parent to edges as follows: for any edge $e\in E(T)$, its parent $p(e)\in E(T)$ is another tree edge incident to $e$ and closer to the root $r$.
-The edges incident to the root do not have a parent and are all siblings.
-Otherwise, the siblings of $e$ are incident to the same parent, denoted as $s(e)=\left\{e^{\prime}\in E(T)\setminus\{e\}\mid p(e^{\prime})=p(e)\right\}$.
 Example 19.
-Given the 6-clique line graph $L_{6}$ in Figure˜2(b), we find a rooted MST $T_{W}$ as shown in Figure˜2(a).
-Among all the tree edges shown as black solid lines, we can only apply the sliding transformation to the edge $(T,P)$ whose weight $\omega(T,P)=|\{a\}|=1$ is lighter than its parent tree edge $(W,T)$ with $\omega(W,T)=|\{a,c\}|=2$.
-We slide along the tree edge $(W,T)$ to the root so that the edge $(T,P)$ becomes $(W,P)$ as shown by the solid curve in Figure˜2(b).
+Given the 6-clique line graph $L_{6}$ in fig.˜2(b), we find a rooted MST $T_{W}$ as shown in fig.˜2(a).
+Among all the tree edges shown as black solid lines, we can only apply the sliding transformation to the edge $\{T,P\}$ whose weight $\omega(\{T,P\})=|\{a\}|=1$ is lighter than its parent tree edge $\{W,T\}$ with $\omega(\{W,T\})=|\{a,c\}|=2$.
+We slide along the tree edge $\{W,T\}$ to the root so that the edge $\{T,P\}$ becomes $\{W,P\}$ as shown by the solid curve in fig.˜2(b).
 Non-tree edges can slide similarly.
-We consider two examples, $(S,U)$ illustrated with a dashed line and $(Y,T)$ with a dotted line in Figure˜2(a).
+We consider two examples, $\{S,U\}$ illustrated with a dashed line and $\{Y,T\}$ with a dotted line in fig.˜2(a).
 All other non-tree edges are shown in light gray.
-We can slide along tree edge $(U,W)$ to the root so that the edge $(S,U)$ becomes $(S,W)$ as shown by the dashed curve in Figure˜2(b).
-Both ends of $(Y,T)$ can slide along the tree edges to the root so that the edge becomes a self-loop as shown by the dotted loop in Figure˜2(b).
-This edge will not appear in any spanning tree of $G^{\equiv}$, and therefore not a part of any MST of $L_{6}$.
+We can slide $\{S,U\}$ along tree edge $\{U,W\}$ to become $\{S,W\}$ as shown by the dashed curve in fig.˜2(b).
+Both ends of $\{Y,T\}$ can slide along the tree edges to the root so that the edge becomes a self-loop as shown by the dotted loop in fig.˜2(b).
+This edge will not appear in any spanning tree of $L_{6}^{\equiv}$, and therefore not a part of any MST of $L_{6}$.
 We refer to such an edge as a non-MST edge, as opposed to an MST edge.
-By applying sliding transformations to a fixpoint, we obtain an equivalent graph $G^{\equiv}$, where there are two sets of parallel edges highlighted by thick lines in Figure˜2(c).
-For example, the tree edge $(T,P)$ in $T_{W}$ and non-tree edges $(P,U),(P,T),(P,Y)$ are parallel in $G^{\equiv}$ between $P$ and $W$.
-We can easily verify that each spanning tree of $G^{\equiv}$ corresponds to an MST of $L_{6}$, such as $T^{\prime}_{W}$ in Figure˜2(d).
-The choice of the initial spanning tree can affect the structure of the equivalent graph, but the order of sliding transformations performed has no impact [12].
+Denoting the graph after sliding as $L_{6}^{\prime}$, ordering $L_{6}\leq_{T_{W}}L_{6}^{\prime}$ holds.
+By applying sliding transformations to a fixpoint, we obtain an equivalent graph $L_{6}^{\equiv}$, where there are two sets of parallel edges highlighted by thick lines in fig.˜2(c) (self-loops are omitted).
+For example, the tree edge $\{T,P\}$ in $T_{W}$ and non-tree edges $\{P,U\},\{P,T\},\{P,Y\}$ become parallel in $L_{6}^{\equiv}$ between $P$ and $W$.
+Therefore $L_{6}\leq_{T_{W}}L_{6}^{\prime}\leq_{T_{W}}L_{6}^{\equiv}$ holds under the sliding partial order.
+We can easily verify that each spanning tree of $L_{6}^{\equiv}$ corresponds to an MST of $L_{6}$, such as $T^{\prime}_{W}$ in fig.˜2(d).
 4.2 Enumerating Join Trees of $\alpha$-Acyclic Hypergraphs
-The bottleneck of Eppstein’s algorithm for constructing equivalent graphs of general graphs comes from a subroutine that identifies where each edge will eventually slide to.
+The bottleneck of Eppstein’s algorithm for constructing equivalent graphs of arbitrary weighted graphs lies in a subroutine that identifies where each edge will eventually slide to.
 Because each edge can only slide along a heavier edge, it will eventually be “blocked” by a lighter or equally weighted edge along its path to the root.
 The subroutine essentially performs binary search to find the blocking edge, leading to the $\log n$ factor in the overall complexity.
-The key to our improvement is to show that for every acyclic hypergraph $H$, we can construct an “equivalent hypergraph” $H^{*}$ that shares the same structure with $H$ and has the same set of join trees, but $H^{*}$ admits a special join tree with monotonically increasing weight from root to leaf, which enables constant-time identification of the blocking edge.
+The key to our improvement is to show that for every acyclic hypergraph $H$, we can construct an equivalent hypergraph $H^{*}$ whose join trees one-to-one correspond to those of $H$, but one special join tree of $H^{*}$ has monotonically increasing weight from root to leaf, which enables constant-time identification of the blocking edge.
 Definition 20.
-A monotonic weight join tree $\hat{T}\in\mathcal{T}(H)$ is a rooted join tree of $H$ such that for any $e\in\hat{T}$ that has a parent $p(e)$, $\omega(e)>\omega(p(e))$.
+[…]
+We can always construct such a $H^{*}$ given any $\alpha$-acyclic hypergraph $H$:
+Theorem 21.
+Given an $\alpha$-acyclic hypergraph $H=(X,R,\chi)$, there exists an equivalent hypergraph $H^{*}=(X^{*},R,\chi^{*})$ over the same hyperedge set $R$ that
+  • admits a monotonic weight join tree $\hat{T}\in\mathcal{T}(H^{*})$, and
+  • $\mathcal{T}(H^{*})=\mathcal{T}(H)$.
+(a) $H^{*}_{6}$
+(b) $\hat{T}_{p}$
+(c) $\hat{T}_{p}$ with 3 non-tree edges
+Figure 3: An equivalent hypergraph $H^{*}_{6}$ and its monotonic weight join tree $\hat{T}_{P}$ rooted at $P$.
 Example 22.
-$H^{*}_{6}$ in Figure˜4(a) is a hypergraph with similar structure to $H_{6}$ in Figure˜2(a).
+$H^{*}_{6}$ in fig.˜3(a) is a hypergraph with similar structure to $H_{6}$ in fig.˜2(a).
 $H^{*}_{6}$ differs from $H_{6}$ by one vertex $d^{\prime}$ and admits the same set of join trees.
-Therefore, finding an equivalent graph $G^{\equiv}(H^{*}_{6})$ is sufficient for enumerating the join trees of $H_{6}$.
-$H^{*}_{6}$ also admits a monotonic weight join tree $\hat{T}_{P}$ in Figure˜4(b).
-This means each tree edge in $\hat{T}_{P}$ is already in place, because it cannot slide along the lighter parent.
-For each non-tree edge, we can identify its destination in constant time using algorithms for finding lowest common ancestors [3] and level ancestors [4], as detailed in Algorithm˜3.
-The example shows we can enumerate join trees of $H$ by constructing $H^{*}$ and its equivalent graph $G^{\equiv}(H^{*})$, then enumerate spanning trees of $G^{\equiv}(H^{*})$.
-However, constructing $H^{*}$ has a cost that may exceed our target time complexity.
-We show that, perhaps surprisingly, running Algorithm˜3 directly on $H$ produces the same equivalent graph $G^{\equiv}(H^{*})$, so we can skip the costly construction of $H^{*}$.
-Before presenting the algorithm, we first formalize the relationship between $H$ and $H^{*}$ using the concept of hypergraph homomorphisms.
-Note the definition of homomorphisms is not standardized in the literature, and we choose ours carefully to simplify the proofs.
-In particular, our definition is equivalent to the homomorphism of incidence graphs [32].
+Therefore, finding an equivalent graph $G^{\equiv}(L(H^{*}_{6}))$ is sufficient for enumerating the join trees of $H_{6}$.
+$H^{*}_{6}$ also admits a monotonic weight join tree $\hat{T}_{P}$ in fig.˜3(b).
+A monotonic weight join tree $\hat{T}$ can accelerate the construction of the equivalent graph in several ways.
+First, all tree edges are already “in place”, because they cannot slide along their lighter parent edges.
+We therefore need only consider non-tree edges.
+For each non-tree edge $e=\{r_{i},r_{j}\}$, there are two possible cases.
+First, if one endpoint, say $r_{i}$, is an ancestor of the other, say $r_{j}$, then $e$ can only slide to become parallel to the edge right below $r_{i}$ on the path connecting $r_{i}$ and $r_{j}$ in $\hat{T}$.
+This is because all edges on that path must be no lighter than $e$ due to the running intersection property of join trees, but since $\hat{T}$ has monotonically increasing weight, at most one edge (the one right below $r_{i}$) can have equal weight to $e$.
+If this edge has weight equal to $e$, it becomes the blocking edge, and $e$ slides to become parallel to it; otherwise, $e$ slides to a self-loop at $r_{i}$.
+In the second case, if neither endpoint is an ancestor of the other, there are at most two blocking edges, namely those below the lowest common ancestor of $r_{i}$ and $r_{j}$ on the path between them.
+We therefore only need to compare the weight of $e$ with these two potentially blocking edges to determine the destination of $e$.
+Example 23.
+Consider again $\hat{T}_{p}$ in fig.˜3(c).
+First note that every tree edge is heavier than its parent edge, and each non-tree edge is no heavier than any edge on the path connecting its endpoints in the tree.
+There is one edge, $\{T,Y\}$, where one endpoint ($T$) is an ancestor of the other ($Y$).
+The only potentially blocking edge is $\{T,U\}$ right below $T$.
+But because $\{T,U\}$ is heavier, $\{T,Y\}$ slides to become a self-loop at $T$.
+For the edge $\{W,Y\}$, all edges on the tree path connecting the endpoints are heavier, so $\{W,Y\}$ also slides to a self-loop at their LCA $T$.
+Finally, $\{Y,S\}$ is blocked by the two edges $\{P,T\}$ and $\{P,S\}$ below their LCA $P$, as all three edges have a weight of 1.
+Therefore $\{Y,S\}$ slides to $\{T,S\}$.
+Our algorithm for constructing the equivalent graph leverages the insight above to slide each edge in constant time.
+In the following, we first define a few helpful notations and describe a preprocessing step to build helper data structures, before presenting the algorithm.
+Given a rooted MCS tree $T_{r}$ of a line graph $L$, a non-tree edge is $e=\{r_{i},r_{j}\}\in E(L)\setminus E(T_{r})$.
+There is a path in $T_{r}$ between $r_{i}$ and $r_{j}$ via their lowest common ancestor $\mathsf{LCA}(r_{i},r_{j})$.
+We define the LCA edges $\lambda(e)=\lambda(r_{i},r_{j})$ as a set of at most two tree edges on the path and incident to $\mathsf{LCA}(r_{i},r_{j})$.
+If $r_{i}$, $r_{j}$ are ancestor and child, then $|\lambda(e)|=1$, otherwise, $|\lambda(e)|=2$.
+During pre-processing, we first conduct a breadth-first search on $T_{r}$ to obtain the depth table $0pt$ of each tree node in $\mathcal{O}(|T_{r}|)$.
+Then we build two data structures in $\mathcal{O}(|T_{r}|)$ to facilitate the constant-time query of the lowest common ancestor (LCA) [4] given two nodes, and the level ancestor (LA) [3] of a node $r$, which is the ancestor of $r$ at a given depth.
+algorithm˜2 assumes these data structures are prebuilt and available, which allows finding $\lambda(e)$ of any non-tree edge $e$ in constant time as follows:
+$\lambda(\{r_{i},r_{j}\})=\begin{cases}\{\{r_{i},\mathsf{LA}(r_{j},0pt(r_{i})+1)\}\}&\text{if }r_{i}=\mathsf{LCA}(r_{i},r_{j})\\ \{\{r_{j},\mathsf{LA}(r_{i},0pt(r_{j})+1)\}\}&\text{if }r_{j}=\mathsf{LCA}(r_{i},r_{j})\\ \{\{l,\mathsf{LA}(r_{i},d)\},\{l,\mathsf{LA}(r_{j},d)\}\}&\text{otherwise, }l:=\mathsf{LCA}(r_{i},r_{j}),d:=0pt(l)+1\end{cases}$
+Input: Line graph $L=((R,E,\rho),\omega)$ of $\alpha$-acyclic $H$ and an MCS tree $T_{r}$
+Output: Equivalent Graph $G^{\equiv}$
+1 $G^{\equiv}(R,E,\rho^{\equiv})\leftarrow(R,E,\rho)$
+2for $e\in E(L)\setminus E(T_{r})$ do
+3 $\omega^{*}\leftarrow\omega(e)$
+4 $r_{i},r_{j}\leftarrow\rho(e)$
+5 if $|\lambda(e)|=1$ then
+6    $\{\{l,l^{\prime}\}\}\leftarrow\lambda(e)$
+7    if $\omega^{*}<\omega(l,l^{\prime})$ then
+8         set $\rho^{\equiv}(e)$ to $\{l\}$
+9    else
+10         set $\rho^{\equiv}(e)$ to $\{l,l^{\prime}\}$
+11
+12 else
+13    $\{\{l,l_{1}\},\{l,l_{2}\}\}\leftarrow\lambda(e)$
+14    if $\omega^{*}<\omega(l,l_{1})\wedge\omega^{*}<\omega(l,l_{2})$ then
+15         set $\rho^{\equiv}(e)$ to $\{l\}$
+16    else if $\omega^{*}=\omega(l,l_{1})$ then
+17         set $\rho^{\equiv}(e)$ to $\{l,l_{1}\}$
+18    else if $\omega^{*}=\omega(l,l_{2})$ then
+19         set $\rho^{\equiv}(e)$ to $\{l,l_{2}\}$
+20    else
+21         set $\rho^{\equiv}(e)$ to $\{l_{1},l_{2}\}$
+22
+23
+return $G^{\equiv}$
+Algorithm 2 buildEGbuildEG
+(a) MCS Tree $T_{P}(H_{6})$
+(b) 3 Non-tree Edges Processed
+(c) Equivalent Graph $G^{\equiv}(H_{6})$
+Figure 4: algorithm˜2 on $H_{6}$ of fig.˜2(a)
+We are now ready to present algorithm˜2 which constructs $G^{\equiv}$ by sliding each non-tree edge $e=\{r_{i},r_{j}\}$ with weight $\omega^{*}$.
+There are two cases.
+First, if $\lambda(e)$ returns a single edge, then one endpoint is an ancestor of the other.
+By the definition above $\lambda(e)$ always returns the ancestor node $l$ first, while the other node is denoted as $l^{\prime}$.
+The weight of $\{l,l^{\prime}\}$ is then compared with $\omega^{*}$: if $\omega^{*}$ is lighter, $e$ slides to a self-loop at $l$; otherwise the weights must be equal, and $e$ slides to be parallel to $\{l,l^{\prime}\}$.
+Second, if $\lambda(e)$ returns two edges, then neither endpoint is an ancestor of the other.
+In this case we compare $\omega^{*}$ with the two edges in $\lambda(e)$.
+If $\omega^{*}$ is lighter than both, $e$ slides to a self-loop at $l$; if $\omega^{*}$ equals one of them, $e$ slides to be parallel to that edge; otherwise, $\omega^{*}$ equals both, and $e$ slides to be incident to each of the highest depth endpoints of both edges.
+The four cases are exhaustive, as the non-tree edge never outweighs any of its LCA edges by the running intersection property.
+Theorem 24.
+Given a hypergraph $H^{*}=(X^{*},R,\chi^{*})$ with a monotonic weight join tree $\hat{T}$, algorithm˜2 produces an Equivalent Graph $G^{\equiv}(L(H^{*}))$.
+The notation $H^{*}$ highlights later applications to equivalent hypergraphs, although Theorem˜24 applies to any hypergraph admitting a monotonic weight join tree.
+Although algorithm˜2 constructs an equivalent graph of $L(H^{*})$ in constant time per edge, such an $H^{*}$ may be asymptotically larger than $H$.
+Perhaps surprisingly, running algorithm˜2 directly on $H$ produces the same equivalent graph!
+This is because $H^{*}$ is constructed in a way that preserves the structure of $H$, keeping the relative weights of an edge and its LCA edges invariant.
+Intuitively, we construct $H^{*}$ by making “local copies” of the vertices in $H$: this perturbs the weights to be monotonic along $\hat{T}$, while ensuring the weights are adjusted consistently.
+In particular, when making a new copy $x^{\prime}$ of a vertex $x$, we add $x^{\prime}$ to all hyperedges that contain $x$, namely $H|_{x}$.
+This way, $\chi(e)\subset\chi(e^{\prime})$ in $L(H)$ if and only if $\chi(e)\subset\chi(e^{\prime})$ in $L(H^{*})$.
+As a result, all comparisons in algorithm˜2 return the same result when running on $H$ or $H^{*}$.
+We illustrate this with an example, while deferring the detailed construction of $H^{*}$ to algorithm˜6 in appendix˜C.
+Example 25.
+In fig.˜3(a), we made a copy $d^{\prime}$ of $d$ and added it to all hyperedges containing $d$ in $H_{6}$.
+This maintains the relative weights between the tree edge $\{U,Y\}$ and the non-tree edges highlighted in fig.˜3(c): all three non-tree edges remain lighter than $\{U,Y\}$.
+Were we to make a copy of $a$ to add to $H_{6}|_{a}$ (which contains every hyperedge in $H_{6}$), the weights of the non-tree edges relative to their respective LCA edges would also remain unchanged.
+Finally, because $H$ and $H^{*}$ share the same join trees, there must be $G^{\equiv}(L(H))$ that is the same as $G^{\equiv}(L(H^{*}))$.
+Together with Theorem˜24, this means running algorithm˜2 directly on $L(H)$ produces an equivalent graph $G^{\equiv}(L(H))$ of $L(H)$:
+Theorem 26.
+Given the line graph $L$ and an MCS tree $T_{r}$ of an $\alpha$-acyclic hypergraph $H$, algorithm˜2 returns an equivalent graph $G^{\equiv}(L)$ in time $\mathcal{O}(|L|)$.
+Once we compute the equivalent graph $G^{\equiv}(L)$, we can apply existing algorithms [24, 37] to enumerate all its spanning trees by edits with amortized constant delay.
+And because the MCS tree can be constructed in linear time from $H$ which in turn can be recovered from $L(H)$ in linear time, the overall time complexity of enumerating all join trees of $H$ is linear in the size of $L$ plus the number of join trees:
+Theorem 27.
+Given the line graph $L$ of an $\alpha$-acyclic hypergraph $H$, the join trees of $H$ can be enumerated in time $\mathcal{O}(|L|+|\mathcal{T}(H)|)$.
+4.3 $\gamma$-Acyclic Queries
+The run time of algorithm˜2 depends on the size of the line graph which can be quadratically larger than the input hypergraph.
+If $H$ is $\gamma$-acyclic, we can bring the total time complexity of enumeration down to $\mathcal{O}(|H|+|\mathcal{T}(H)|)$.
+Using an algorithm by Leitert [27], we can construct $L(H)$ from $H$ in time $\mathcal{O}(|H|+|L(H)|)$ when $H$ is $\gamma$-acyclic; furthermore, because every edge in $L(H)$ of a $\gamma$-acyclic $H$ is an MST edge [27], we have $|H|+|\mathcal{T}(H)|\in\Omega(|L(H)|)$, and so $\mathcal{O}(|H|+|L(H)|+|\mathcal{T}(H)|)=\mathcal{O}(|H|+|\mathcal{T}(H)|)$, hiding the $|L(H)|$ term in the overall complexity.
+However, the $L(H)$ constructed by Leitert is unweighted, so we need to modify algorithm˜2 slightly to slide each non-tree edge based on the weights of its LCA edges alone.
+We first modify MCS (algorithm˜1) to track the weight of each tree edge as in algorithm˜3.
+Then, we update the cases of algorithm˜2 to those in algorithm˜4.
+First, if $\lambda(e)$ returns a single edge, we directly slide $e$ to be parallel to it, because we know $e$ is an MST edge and therefore will not become a self-loop.
+Otherwise, if $\lambda(e)$ returns two edges, we compare the weights of those two edges, instead of comparing them with $\omega(e)$.
+This is again because we know $e$ is an MST edge, so it must have weight equal to at least one of its LCA edges.
+If one of the LCA edges is lighter, $e$ must have equal weight to that edge, and we slide $e$ to be parallel to it.
+Otherwise, if both LCA edges have the same weight, $e$ must also have the same weight, and we slide $e$ to be under both edges to form a triangle.
+In all cases we only use the weights of the tree edges and do not need to know $\omega(e)$.
+Input: $\gamma$-acyclic Hypergraph $H$ and root $r$
+Output: $T_{r}$ and $\omega$
+1
+[…]
+3while $R\neq\emptyset$ do
+4
+5 for $x\in r\cap X$ do
+6    for $r^{\prime}\in R:x\in r^{\prime}$ do
+7       $p(r^{\prime})\leftarrow r$
+8
+9
+10 $X\leftarrow X\setminus r$
+11 $R\leftarrow R\setminus\{r\}$
+12 $E(T_{r})\leftarrow E(T_{r})\cup\{\{r,p(r)\}\}$
+13 $\omega(\{r,p(r)\})\leftarrow|r\cap p(r)|$
+14 $r\leftarrow\operatorname*{arg\,max}_{r\in R}|r\setminus X|$
+return $T_{r}$, $\omega$
+Algorithm 3 MCS ($\gamma$-acyclic)
+Input: Weighted $T_{r}$, unweighted $L(H)$
+Output: Equivalent Graph $G^{\equiv}$
+1
+2 for $e\in E(L)\setminus E(T_{r})$ do
+3
+4 $\dots$
+5 if $|\lambda(e)|=1$ then
+6    $\dots$ set $\rho^{\equiv}(e)$ to $\{l,l^{\prime}\}$
+7 else
+8    $\{\{l,l_{1}\},\{l,l_{2}\}\}\leftarrow\lambda(e)$
+9    else if $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\omega(l,l_{1})<\omega(l,l_{2})$ then
+10         set $\rho^{\equiv}(e)$ to $\{l,l_{1}\}$
+11
+12    else if $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\omega(l,l_{1})>\omega(l,l_{2})$ then
+13         set $\rho^{\equiv}(e)$ to $\{l,l_{2}\}$
+14
+15    else set $\rho^{\equiv}(e)$ to $\{l_{1},l_{2}\}$
+16
+17
+18return $G^{\equiv}$
+Algorithm 4 buildEGbuildEG ($\gamma$-acyclic)
+Theorem 28.
+The join trees of a $\gamma$-acyclic $H$ can be enumerated in time $\mathcal{O}(|H|+|\mathcal{T}(H)|)$.
+5 The Canonical Join Tree of a Berge-Acyclic Query
+An acyclic query can have exponentially many join trees with respect to its size.
+For example, the line graph of a clique query with $n$ relations is an $n$-clique $K_{n}$ with $n^{n-2}$ join trees by Cayley’s formula [8].
+Enumerating all join trees can be prohibitive for large queries.
+The query optimizer does not need to consider all possible join trees to achieve good performance.
+For example, the implementation of Yannakakis’ algorithm by Zhao et al.
+[48] achieves similar performance on any join tree rooted at the largest relation.
+An alternative to enumeration is therefore to simply construct one join tree for a given root.
+This can be done in linear time by the Maximum Cardinality Search (MCS) algorithm from a chosen relation as shown in algorithm˜1.
+In this section, we prove that for a Berge-acyclic query the MCS algorithm produces a shallowest tree, where the depth of each tree node is minimized.
+We prove that this shallowest tree is unique, and therefore call it the canonical join tree.
+Definition 29.
+A join tree $T_{r}$ rooted at $r$ is canonical if $0pt(T_{r},r_{i})\leq 0pt(T^{\prime}_{r},r_{i})$ for any other join tree $T^{\prime}_{r}$ rooted at $r$ and any $r_{i}\in R(T_{r})=R(T^{\prime}_{r})$.
+A shallow join tree has practical benefits.
+For example, the depth of the join tree determines the number of sequential steps required in a parallel join algorithm.
+A shallow join tree also tends to be wide and have more leaves, allowing better utilization of indices.
+Although Berge-acyclicity was thought to be too restrictive when it was first introduced to database theory [15], we found it to be general enough to cover almost all acyclic queries encountered in the wild.
+As shown in table˜1, among 10 45410\,454 queries from five popular benchmarks, $9285$ are $\alpha$-acyclic, and only 8 of these are not Berge-acyclic.
+In retrospect, this should not be surprising, as most joins in relational databases are over primary/foreign keys.
+Emerging workloads in graph databases usually involve simple graphs and seldom require composite key joins.
+A query without composite key joins admits a linear hypergraph, where each pair of hyperedges shares at most one vertex.
+The following result establishes an equivalence between $\alpha$-acyclicity with linearity and Berge-acyclicity.
+Name # Queries # $\alpha$-Acyclic # Composite-Key Joins # Berge-Acyclic
+TPC-H[42] 22 21 2 19
+JOB[26] 113 113 0 113
+STATS-CEB[19] 2603 2603 0 2603
+CE[9] 3004 1839 0 1839
+Spider-NLP[47] 4712 4709 6 4703
+Table 1: Acyclic queries in the benchmarks (all $\alpha$-acyclic queries are also $\gamma$-acyclic.)
+Proposition 30.
+An $\alpha$-acyclic hypergraph is Berge-acyclic if and only if it is linear.
+By section˜5, every edge in the line graph of a Berge-acyclic hypergraph has a weight of 1. Every spanning tree is a maximum spanning tree, therefore a join tree.
+Corollary 31.
+For a Berge-acyclic hypergraph $H$, any spanning tree of $L(H)$ is a join tree.
+In the rest of this section, we prove the existence and uniqueness of the canonical join tree rooted at any relation of a Berge-acyclic hypergraph, and show that it can be constructed by MCS as in algorithm˜1.
+The key insight leading to the existence and uniqueness of the canonical join tree is that the line graph $L$ of a Berge-acyclic hypergraph is geodetic [33], meaning that there is a unique shortest path between any pair of vertices.
+The canonical join tree can then be constructed by taking the union of all shortest paths from the root to each other vertex, as this will guarantee minimal depth for each vertex.
+The key step to establish the geodetic property is to show that $L$ is a special class of chordal graphs called block graphs.
+Chordal graphs are intimately related to acyclic hypergraphs [11], and a block graph is a special chordal graph defined as follows:
+Definition 32 (Chordal and Block).
+A simple graph $G=(R,E)$ is
+  • chordal [12] if every cycle of length at least 4 has a chord, i.e., an edge that is not part of the cycle but connects two vertices of the cycle;
+  • a block graph [20] if it is chordal and diamond-free, i.e., no subgraph induced by any $R^{\prime}\subseteq R$ is a diamond (section˜3).
+Lemma 33.
+The line graph $L$ of a Berge-acyclic hypergraph $H$ is a block graph.
+Together with the fact that every block graph is geodetic (with a unique shortest path between any two vertices) [33], section˜5 implies the following corollary.
+Corollary 34.
+Let $L$ be the line graph of a Berge-acyclic hypergraph $H$. There is a unique shortest path between any two vertices in $L$.
+We are now ready to prove the existence and uniqueness of the canonical join tree.
+Theorem 35.
+A Berge-acyclic hypergraph $H$ has a unique canonical tree.
+Proof.
+Let $L$ be the line graph of $H$, and $\mathsf{P}(r,r^{\prime})$ be the shortest path in $L$ between $r,r^{\prime}\in R(L)$, we prove that $T_{r}=\bigcup_{r^{\prime}\in R(L)}\mathsf{P}(r,r^{\prime})$ is the unique canonical join tree for $H$ rooted at $r$.
+By section˜5 any spanning tree of $L$ is a join tree.
+$T_{r}$ is connected and spans all vertices in $R(L)$, because it contains the shortest path from $r$ to every $r^{\prime}\in R(L)$.
+It remains to show that $T_{r}$ is acyclic, which we prove by induction on the distance $\mathsf{dist}(r,r^{\prime})$ between $r$ and $r^{\prime}$.
+Let $R_{d}=\left\{r^{\prime}\in R(L)\mid\mathsf{dist}(r,r^{\prime})\leq d\right\}$.
+$R_{0}=\{r\}$ contains only the root.
+The subgraph $T_{r}|_{R_{0}}$ is trivially acyclic.
+Assuming that $T_{r}|_{R_{d>0}}$ is acyclic, we consider a vertex $r^{\prime}\in R_{d+1}\setminus R_{d}$.
+section˜5 guarantees a unique shortest path between each pair of vertices $r,r^{\prime}\in R(L)$.
+Each $r^{\prime}$ is connected to a unique neighbor $r^{\prime\prime}\in R_{d}\setminus R_{d-1}$ that is at distance $d$ from $r$.
+Otherwise, there are at least two distinct shortest paths from $r$ to $r^{\prime}$.
+Therefore $T_{r}|_{R_{d+1}}$ is acyclic, and $T_{r}$ is a spanning tree thus a join tree of $H$.
+The join tree $T_{r}$ is canonical, because the path from $r$ to each $r^{\prime}\in R(L)$ is the shortest and therefore minimizing the depth $0pt(T_{r},r^{\prime})$.
+The canonical tree is unique by section˜5.
+∎
+The canonical join tree can be constructed by the MCS algorithm as shown in algorithm˜1:
+Theorem 36.
+Given a Berge-acyclic hypergraph $H$, running algorithm˜1 from $r\in R(H)$ constructs the canonical tree $T_{r}(H)$.
+6 Converting a Binary Join Plan to a Join Tree
+Input: Left-deep plan $(r_{1},\dots,r_{n})$
+Output: Rooted join tree $T_{r_{1}}$
+1
+[…]
+3for $i\leftarrow 2$ to $n$ do
+4
+5  key $\leftarrow r_{i}\cap\bigcup_{k<i}r_{k}$
+6 for $j\leftarrow 1$ to i1i-1 do
+7    if $\text{key}\subseteq r_{j}$ then
+8
+9       $E(T_{r_{1}})\leftarrow E(T_{r_{1}})\cup\{\{r_{i},r_{j}\}\}$
+10       break
+11
+12
+13 return $T_{r_{1}}$
+Algorithm 5 Binary-to-JT Conversion
+Refer $\RightarrowRefer to caption$
+Figure 5: algorithm˜5[22] on JOB-3a[26]
+Recent approaches [22, 2] that convert a binary join plan into a join tree have gained popularity as they allow system builders to leverage existing query optimizers designed for binary join plans.
+In this section, we focus on an algorithm by Hu et al.
+[22] to convert left-deep linear join plans into join trees as shown in fig.˜5.
+We prove that the algorithm converts any connected left-deep linear join plan into a join tree if and only if the query is $\gamma$-acyclic.
+This can be seen as a new characterization of $\gamma$-acyclic queries.
+We formally define binary join plans and describe the algorithm by Hu et al.
+in algorithm˜5.
+Definition 37.
+A left-deep linear plan is a sequence of hyperedges $(r_{1},r_{2},\ldots,r_{n})$. It is connected if for each $r_{i\geq 2}$, $\exists r_{j<i}:r_{i}\cap r_{j}\neq\emptyset$.
+Query optimizers strive to produce connected plans, to avoid expensive Cartesian products.
+Many optimizers produce exclusively left-deep linear plans.
+Plans that are not left-deep are called bushy, and such plans may still be decomposed into left-deep fragments [44].
+Given a left-deep linear plan, Hu et al.
+[22] generate a join tree with algorithm˜5.
+The algorithm chooses the first relation $r_{1}$ as the root and iterates through the rest of the plan.
+For each relation $r_{i\in[2,n]}$, it finds the first relation $r_{j}$ that contains all attributes shared by $r_{i}$ with all previous relations, $\text{key}:=r_{i}\cap\bigcup_{k<i}r_{k}$, and assigns $r_{j}$ as the parent of $r_{i}$.
+The algorithm constructs a join tree if it finds a parent for each $r_{i}$.
+Hu et al.
+[22] proved that the algorithm succeeds whenever the input plan is the reverse of a GYO-reduction order.
+They also observed that every left-deep linear plan produced for queries in standard benchmarks is indeed the reverse of a GYO-reduction order.
+This is not a coincidence, as we show that every connected left-deep linear join plan must be the reverse of a GYO-reduction order if and only if the query is $\gamma$-acyclic.
+Theorem 38.
+A query is $\gamma$-acyclic if and only if every connected left-deep linear join plan for the query is the reverse of a GYO-reduction order.
+Immediately following Theorem˜38, we can conclude the following.
+Corollary 39.
+For any $\gamma$-acyclic query, algorithm˜5 converts a given connected left-deep linear join plan to a join tree.
+7 Conclusion and Future Work
+We proposed three approaches for constructing join trees.
+Our enumeration algorithm in section˜4 generates join trees by edits with amortized constant delay; in section˜5, we showed that the Maximum Cardinality Search algorithm constructs the unique shallowest join tree for any Berge-acyclic query; in section˜6, we characterize the class of binary join plans that can be converted to join trees.
+Practitioners can choose from the three approaches to integrate instance-optimal algorithms into their system: a cost-based optimizer can use our enumeration algorithm to generate candidate query plans; a system aiming to support very large queries can implement the MCS algorithm to generate shallow join trees to improve parallelism; and a more conservative extension can derive join trees using existing optimization infrastructure, while our final result guarantees the validity of the output.
+Future work includes compact representations of join trees for dynamic programming, as in binary plan optimizers, and the challenging cost estimation for Yannakakis-style algorithms: the random-walk approach [28] models joint probabilities for binary joins, whereas an efficient and accurate solution for semijoins remains to be found.
+(a) $H_{3}$
+(b) $T_{S}(H_{3})$
+(c) $H_{5}$
+(d) $T_{W}(H_{5})$
+(e) $T^{\prime}_{W}(H_{5})$
+Figure 6: $H_{3}$ is not Berge-acyclic but admits a unique canonical join tree $T_{S}(H_{3})$ with any relation chosen as root, such as $T_{S}(H_{3})$.
+$H_{5}$ is $\gamma$-acyclic and does not admit unique canonical join trees at any relation chosen as root.
+For example, $T_{W}(H_{5})$ and $T^{\prime}_{W}(H_{5})$ are MCS trees generated by algorithm˜1.
+Neither of them is a canonical join tree rooted at $W$.
+Our work also raises further theoretical questions.
+Can join tree enumeration achieve worst-case constant delay?
+We proved Berge-acyclicity sufficient for the existence and uniqueness of the canonical join tree, but it is not necessary, and $\gamma$-acyclicity is insufficient as shown in fig.˜6.
+What is the precise characterization of hypergraphs that admit a unique canonical join tree for any root, or for some root?
+How can our algorithms be extended to enumerate tree decompositions à la Carmeli et al.
+[7]?
+One possible direction is to further develop connections between MSTs of the line graph and tree decompositions of the hypergraph.
+For example, we are considering using the determinants of the line graph Laplacian as heuristics to guide the search for good tree decompositions.
+Appendix A Data Structures for Constant Time Operations on Graphs
+Given line graph $L$ and MCS tree $T$, algorithm˜2 iterates over non-tree edges of $L$.
+Moreover, the algorithm uses data structures $0pt(\cdot)$, $\mathsf{LA}(\cdot)$ and $\mathsf{LCA}(\cdot)$, in determining the LCA edges of a non-tree edge.
+In this section, we describe how to enumerate over non-tree edges in time linear in the number of vertices and edges of $L$ and how to determine the LCA edges of a non-tree edge in constant time.
+Given a graph $G=(V,E)$, we assume the vertices are represented by integers from 1 to $|V|$.
+Given a tree $T$ of $G$ rooted at node $r$, data structure $0pt_{T}(\cdot)$ stores for each node $r_{i}$ the length of the shortest path between $r$ and $r_{i}$ in $T$.
+Data structure $\mathsf{LCA}_{T}(\cdot)$ stores for each pair of nodes $r_{i},r_{j}$ the lowest common ancestor of $r_{i},r_{j}$ in $T$.
+Finally, data structure $\mathsf{LA}_{T}(\cdot)$ stores for each node-integer pair $r_{i},j$, the ancestor of node $r_{i}$ at depth $j$.
+Each of these data structures can be established from a tree $T$ in time linear in the size of the tree and support constant time look-ups.
+Observation 40.
+Given rooted tree $T$, data structures $0pt(\cdot)$, $\mathsf{LCA}(\cdot)$ and $\mathsf{LA}(\cdot)$ can be established in time $\mathcal{O}(|T|)$ and support $\mathcal{O}(1)$ time look-ups [10, 35, 4, 3].
+Given the line graph $L$ of an $\alpha$-acyclic hypergraph $H$ and an MCS tree $T$ of $H$, recall that the LCA edges $\lambda(e)$ of a non-tree edge $e=(r_{i},r_{j})$ of $T$ is defined as the set of edges incident to $\mathsf{LCA}(r_{i},r_{j})$ that lie on the unique path between $r_{i}$ and $r_{j}$ in $T$.
+The following observation shows that the LCA edges of a non-tree edge can be found in constant time given data structures $0pt(\cdot)$, $\mathsf{LCA}(\cdot)$ and $\mathsf{LA}(\cdot)$.
+Observation 41.
+Given the line graph $L$ of an $\alpha$-acyclic hypergraph $H$, an MCS tree $T$ and data structures $0pt(\cdot)$, $\mathsf{LCA}(\cdot)$ and $\mathsf{LA}(\cdot)$, the LCA edges of any non-tree edge can be found in constant time.
+Proof.
+Given a non-tree edge $e=(r_{i},r_{j})$ we find $\mathsf{LCA}(r_{i},r_{j})=l$ in constant time by appendix˜A.
+The depth $d=0pt(l)+1$ is likewise found in constant time.
+The edge $e_{1}=(l,r_{1})$, where node $r_{1}=\mathsf{LA}(r_{i},d)$ is found by a constant time look-up, is adjacent to $l$ and lies on the unique path between $r_{i}$ and $r_{j}$.
+The same holds for edge $e_{2}=(l,r_{2})$ where node $r_{2}=\mathsf{LA}(r_{j},d)$.
+∎
+In the remainder of this section, we describe how algorithm˜2 iterates over and performs sliding transformations on non-tree edges.
+We use an array $L$ to represent the adjacency list of a line graph where entry $L[i]$ consists of a linked list of tuples $(j,\omega)$ where $j$ is a neighbor of $i$ and $\omega$ is the weight of the edge $(i,j)$.
+A rooted tree $T_{r}$ is likewise implemented as an array, where the index $i$ represents the vertex $i$, and the $i$-th entry $T_{r}[i]$ stores $(p,\omega)$ where $p$ is the parent of $i$ and $\omega$ is the weight of the edge $(i,p)$.
+The depth table of the tree is also implemented as an array to enable constant time lookup for the depth of a vertex $0pt(T_{r},i)$.
+As we iterate over each edge $(j,\omega)$ of $L[i]$, we check if the current edge $(i,j)$ is a tree edge by performing constant-time query of $T_{r}[i]$ and $T_{r}[j]$ to see if one of them is the parent of the other.
+From appendix˜A we know that we can find the LCA edges $\lambda(i,j)=\{e_{1},e_{2}\}$ in constant time.
+There’s no general solution to lookup an entry in a linked list in constant time.
+Instead of iterating over the linked-list representation of $L$ to find the edge weights of $e_{1}$ and $e_{2}$, we can lookup their weights from the array representation of $T_{r}$ in constant time.
+Let $e_{1}=\{r_{1},p(r_{1})\}$ and $e_{2}=\{r_{2},p(r_{2})\}$ with weights $\omega_{1}$ and $\omega_{2}$.
+Because both $e_{1}$ and $e_{2}$ are tree edges, we can find their weights by querying $T_{r}[r_{1}]=(p(r_{1}),\omega_{1})$ and $T_{r}[r_{2}]=(p(r_{2}),\omega_{2})$.
+Each of these operations takes constant time.
+Appendix B Missing Proofs in section˜3
+We prove section˜3 of section section˜3 below.
+section˜3 Let $T$ be an MCS tree. For an edge $e\in T$ that has a parent, then
+  1. 1.
+    $e\not\subseteq p(e)$
+  2. 2.
+    For another edge $e^{\prime}\in T:(e\setminus p(e))\ \cap\ (e^{\prime}\setminus p(e^{\prime}))\ \neq\ \emptyset\quad\implies\quad p(e)=p(e^{\prime}).$
+Proof.
+In algorithm˜1, a child $r_{c}$ is only connected to a parent $r$ if they share some previously unmarked vertex $x$.
+Because every vertex is marked only once, and the parent $r_{p}$ of $r$ was labeled before $r$, $x$ cannot be marked by $r$’s parent $r_{p}$.
+Let $e$ be edge between $r$ and $r_{c}$, then the edge between $r$ and $r_{p}$ is $p(e)$, and $x\in e$ but $x\notin p(e)$, therefore $e\not\subseteq p(e)$.
+We prove the second statement by contrapositive, assuming $p(e)\neq p(e^{\prime})$.
+We also treat each edge as a set of vertices below, as we can identify each tree edge with its set of incident vertices.
+Case 1:
+$p(e^{\prime})=e$, then $(e\setminus p(e))\cap(e^{\prime}\setminus e)\subseteq(e\cap e^{\prime})\setminus e=\emptyset$.
+Case 2:
+$p(e^{\prime})\neq e$.
+Without loss of generality, we suppose $e$ is no further from the root than $e^{\prime}$.
+Let $p(e^{\prime})=e^{\prime\prime}$, then $e^{\prime\prime}$ is on the path between $e$ and $e^{\prime}$ in $T$.
+If $(e\setminus p(e))\cap(e^{\prime}\setminus e^{\prime\prime})\neq\emptyset$, then $\exists\ x\in(e^{\prime}\cap e)\setminus e^{\prime\prime}$, which violates the running intersection property of join trees.
+Appendix C Missing Proofs in section˜4
+In this section we present the missing proofs of section˜4.
+In particular, in section˜C.1 we show that there exists an equivalent hypergraph $H^{*}$ of a given hypergraph $H$ that admits a monotonic weight join tree and preserves the set of join trees of $H$.
+This is proven in Theorem˜21.
+Thereafter, we prove Theorem˜24 showing that algorithm˜2 produces an equivalent graph when given input $L(H^{*})$ and a monotonic weight join tree of $H^{*}$.
+Next we show that algorithm˜2 behaves identically on the input of line graph $L(H)$ and an MCS tree of $H$ as it does on $L(H^{*})$ and a monotonic weight join tree of $H^{*}$.
+algorithm˜2 therefore produces an equivalent graph of $L(H)$.
+This proven in Theorem˜26.
+Finally, we prove Theorem˜27, showing that the join trees of an $\alpha$-acyclic hypergraph can be enumerated in time $\mathcal{O}(|L|+|\mathcal{T}(H)|)$ given its line graph $L$.
+In section˜C.2 we prove Theorem˜28 showing that the join trees of a $\gamma$-acyclic $H$ can be enumerated in time $\mathcal{O}(|H|+|\mathcal{T}(H)|)$.
+C.1 Missing Proofs in section˜4.2
+We define the notion of an equivalent hypergraph below.
+Definition 42.
+Given an $\alpha$-acyclic hypergraph $H=(X,R,\chi)$, hypergraph $H^{*}=(X^{*},R,\chi^{*})$, over the same hyperedge set $R$, is an equivalent hypergraph of $H$, if $H^{*}$,
+  • admits a monotonic weight join tree $\hat{T}\in\mathcal{T}(H^{*})$, and
+  • $\mathcal{T}(H^{*})=\mathcal{T}(H)$.
+We formalize the relationship between an $\alpha$-acyclic hypergraph $H$ and its equivalent hypergraph $H^{*}$ using the concept of hypergraph homomorphisms.
+Note the definition of homomorphisms is not standardized in the literature, [^4] A textbook on hypergraphs [5] contains two different definitions of hypergraph homomorphism; one is the same as ours, while the other is also used by Scheidt and Schweikardt [34].
+and we choose ours carefully to simplify the proofs.
 Definition 43 (Hypergraph Homomorphism).
-For two hypergraphs $H_{1}$, $H_{2}$, a homomorphism $\mathsf{f}$ is a pair of functions $(\mathsf{f}_{X}:X(H_{1})\rightarrow X(H_{2}),\mathsf{f}_{R}:R(H_{1})\rightarrow R(H_{2}))$ such that:
-$\forall r_{1}\in R(H_{1}),x_{1}\in r_{1}:\mathsf{f}_{X}(x_{1})\in\mathsf{f}_{R}(r_{1})$
-We write $H_{1}\rightarrow H_{2}$ to denote that there is a homomorphism from $H_{1}$ to $H_{2}$.
-When there is no ambiguity from the context, we simply write $\mathsf{f}$ for $\mathsf{f}_{X}$ or $\mathsf{f}_{R}$.
+Let $H_{1}$ and $H_{2}$ be hypergraphs.
+A homomorphism from $H_{1}$ to $H_{2}$ is a pair of functions $(\mathsf{f}_{X}:X(H_{1})\rightarrow X(H_{2}),\mathsf{f}_{R}:R(H_{1})\rightarrow R(H_{2}))$, such that for every hyperedge $r_{1}\in R(H_{1})$ and every vertex $x_{1}\in r_{1}$, the image $\mathsf{f}_{X}(x_{1})$ belongs to the hyperedge $\mathsf{f}_{R}(r_{1})$ in $H_{2}$.
+We write $H_{1}\rightarrow H_{2}$ to denote the existence of such a homomorphism.
+When there is no ambiguity from context, we simply write $\mathsf{f}$ for $\mathsf{f}_{X}$ or $\mathsf{f}_{R}$.
 Intuitively, homomorphisms preserve relations.
@@
 Definition 44 (Strong Homomorphism).
-A homomorphism $\mathsf{f}:H_{1}\rightarrow H_{2}$ is called strong if it also satisfies:
-$\forall r_{1}\in R(H_{1}),x_{1}^{\prime}\not\in r_{1}:\mathsf{f}_{X}(x_{1}^{\prime})\not\in\mathsf{f}_{R}(r_{1})$
-We write $H_{1}\twoheadrightarrow H_{2}$ to denote that there is a strong homomorphism from $H_{1}$ to $H_{2}$.
-In this paper we only consider hypergraphs over the same set of edges, [^3] The hypergraphs share the same hyperedge set $R$ but differ in the hyperedge-to-vertices functions $\chi$, i.e., an edge may have different sets of vertices in different hypergraphs.
+Let $\mathsf{f}:H_{1}\rightarrow H_{2}$ be a hypergraph homomorphism.
+We call $\mathsf{f}$ strong if it also reflects non-incidence, in the following sense: for every hyperedge $r_{1}\in R(H_{1})$ and every vertex $x_{1}\in X(H_{1})$ that does not belong to $r_{1}$, the vertex $\mathsf{f}_{X}(x_{1})$ does not belong to the hyperedge $\mathsf{f}_{R}(r_{1})$ in $H_{2}$.
+We write $H_{1}\twoheadrightarrow H_{2}$ to denote the existence of a strong homomorphism from $H_{1}$ to $H_{2}$.
+Equivalently, under a strong homomorphism, a vertex belongs to a hyperedge in $H_{1}$ if and only if its image belongs to the image of that hyperedge in $H_{2}$.
+We only consider hypergraphs over the same set of edges, [^5] The hypergraphs share the same hyperedge set $R$ but differ in incidence function $\chi$, i.e., the same edge may be incident to different vertices in different hypergraphs.
 and from now on we assume $\mathsf{f}_{R}$ is the identity function.
-We now show that if there is a strong homomorphism $\mathsf{f}:H^{\prime}\twoheadrightarrow H$, then every join tree of $H$ is also a join tree of $H^{\prime}$.
+We show that if there is a strong homomorphism $\mathsf{f}:H^{\prime}\twoheadrightarrow H$, then every join tree of $H$ is also a join tree of $H^{\prime}$.
 Lemma 45.
 If $H^{\prime}\twoheadrightarrow H$ then $\mathcal{T}(H)\subseteq\mathcal{T}(H^{\prime})$.
-We construct $H^{*}$ such that $H^{*}\twoheadrightarrow H$ and $H^{*}$ admits a monotonic weight join tree $\hat{T}$, by first using the MCS algorithm to construct a join tree $T$ of $H$ called the MCS tree.
-Then, we perturb the weights of $T$ so that they increase monotonically from the root to leaves.
-Finally, we construct $H^{*}$ such that the perturbed tree is a join tree of $H^{*}$.
-The next result shows every MCS tree is already “somewhat monotonic”, in that every edge must contain some variable not in its parent.
-With abuse of notation we write $e_{1}\subseteq e_{2}$ for $\chi(e_{1})\subseteq\chi(e_{2})$ and use the common set operations directly on edges to reduce clutter.
-Lemma 16.
-Let $T$ be an MCS tree. For any edge $e\in T$ that has a parent, then $e\not\subseteq p(e)$.
-Furthermore, if two edges share any variable not in their parents, they must be siblings.
-Lemma 17.
-Let $T$ be an MCS tree. For two edges $e,e^{\prime}\in T$ that have parents,
-$(e\setminus p(e))\ \cap\ (e^{\prime}\setminus p(e^{\prime}))\ \neq\ \emptyset\quad\implies\quad p(e)=p(e^{\prime}).$
-We now show how to derive from an $\alpha$-acyclic hypergraph $H$ an equivalent hypergraph $H^{\prime}$ that preserves $\mathcal{T}(H)$ and admits a monotonic weight join tree $\hat{T}$.
+Proof.
+Let $T$ be a join tree of $H$, and $x^{\prime}$ be any vertex in $H^{\prime}$.
+Consider the neighborhood $H^{\prime}|_{x^{\prime}}$ of $x^{\prime}$ in $H^{\prime}$ comprising the set of hyperedges $\left\{r^{\prime}\in R(H^{\prime})\mid x^{\prime}\in r^{\prime}\right\}$.
+Since $H^{\prime}\twoheadrightarrow H$, we have that $x^{\prime}\in r^{\prime}$ if and only if $\mathsf{f}(x^{\prime})\in\mathsf{f}(r^{\prime})=r^{\prime}$.
+The neighborhood $H|_{\mathsf{f}(x^{\prime})}$ of $\mathsf{f}(x^{\prime})$ in $H$ comprises of the set of hyperedges $\left\{r\in R(H)\mid\mathsf{f}(x^{\prime})\in r\right\}$.
+Therefore, $r^{\prime}\in R(H^{\prime}|_{x^{\prime}})\iff\mathsf{f}(r^{\prime})=r^{\prime}\in R(H|_{\mathsf{f}(x^{\prime})})$.
+Because $T$ is a join tree of $H$, $H|_{\mathsf{f}(x^{\prime})}$ is connected in $T$, therefore $H^{\prime}|_{x^{\prime}}$ is also connected in $T$ and $T$ is also a join tree of $H^{\prime}$.
+∎
+Our goal will be to modify an input hypergraph $H$ such that its line graph admits a monotonic weight join tree.
+The lemma above ensures that if this is done in such a way that a strong homomorphism between $H$ and its modification is preserved then then the set of join trees of $H$ is preserved.
 Definition 46 (Vertex Duplication).
@@
     $L(H^{\prime})=L(H)$
-We are now ready to construct $H^{*}$ by vertex duplication.
-Input: MCS tree $T_{r}(H)$
+Proof.
+1.
+We only need to show that $H^{\prime}\twoheadrightarrow H$ ($H^{\prime}\twoheadleftarrow H$) holds for $H^{\prime}$ derived from $H$ by a single vertex duplication.
+$H^{\prime}\twoheadrightarrow H$ ($H^{\prime}\twoheadleftarrow H$) will then hold for any duplicate $H^{\prime}$ derived by a series of vertex duplications since strong homomorphisms are closed under composition.
+Suppose $H^{\prime}$ is derived from $H$ by duplicating $x\in X(H)$ to $x^{\prime}$.
+Consider a pair of functions $(\mathsf{f}_{X}:X(H^{\prime})\mapsto X(H),R(H)_{\texttt{id}})$ where $\mathsf{f}_{X}$ maps $x^{\prime}$ to $x$.
+Otherwise, it is an identity map $X(H)_{\texttt{id}}$.
+This function pair constitutes a strong hypergraph homomorphism $H^{\prime}\twoheadrightarrow H$.
+Likewise, the pair of functions $(X(H)_{\texttt{id}},R(H)_{\texttt{id}})$ constitutes a homomorphism $H\twoheadrightarrow H^{\prime}$.
+2. It follows from section˜C.1 and the facts above.
+3.
+By the strong homomorphisms, the hyperedge sets have the same cardinality $|R(H)|=|R(H^{\prime})|$.
+So do the vertex sets of the line graphs, $R(L(H))=R(H)$ and $R(L(H^{\prime}))=R(H^{\prime})$.
+By construction, the vertex duplication adds $x^{\prime}$ to each hyperedge containing $x$, therefore does not create any additional edges in the line graph, $E(L(H^{\prime}))=E(L(H))$.
+∎
+Input: $\alpha$-acyclic hypergraph $H$ and rooted MCS tree $T_{r}$
+Output: $H^{*}$ in Theorem˜21
 1
-2for $r^{\prime}\in c(r)$ do
-3 $e\leftarrow(r,r^{\prime});\quad q$.enqueue($e$)
-4
-5while $q\neq\emptyset$ do
+[…]
+[…]
+4for $r^{\prime}\in c(r)$ in $T_{r}$ do
+5 $e\leftarrow\{r,r^{\prime}\};\quad q$.enqueue($e$)
+6
+7while $q\neq\emptyset$ do
 8 $e\leftarrow q.$dequeue()
@@
 11    $x\leftarrow\text{a vertex in }\chi(e)\setminus\chi(p(e))$
-10     make $\Delta$ duplicates of $x$
-11
+12     make $\Delta$ duplicates of $x$ in $X^{*}$
+13
 14 for $e^{\prime}\in c(e)$ do
 15    $q$.enqueue($e^{\prime}$)
-14
-15
+16
+17
+18return $H^{*}$
 Algorithm 6 Construction of $H^{*}$
-(a) $H^{*}_{6}$
-(b) $\hat{T}_{p}$
-Figure 3: An equivalent hypergraph $H^{*}_{6}$ and its monotonic weight join tree $\hat{T}_{P}$
-Theorem 20.
-Given an MCS tree $T$ of an $\alpha$-acyclic hypergraph $H$, there exists a hypergraph $H^{*}$ that
-  • admits a monotonic weight join tree $\hat{T}\in\mathcal{T}(H^{*})$ whose edges are in a one-to-one correspondence to those of $T$
-  • satisfies $\mathcal{T}(H^{*})=\mathcal{T}(H)$.
+Vertex duplication allows us to modify a hypergraph while preserving its set of join trees.
+In particular, duplicating vertices can selectively increase the weights of chosen hyperedge intersections without affecting the line graph or the join tree structure (cf.
+section˜C.1).
+We use this operation to enforce monotonicity of edge weights along a given join tree.
+Given an $\alpha$-acyclic hypergraph $H$ together with a rooted MCS tree $T$, we can construct an equivalent hypergraph $H^{*}$ by duplicating vertices along the edges of $T$.
+The construction proceeds in a breadth-first manner from the root of $T$ and duplicates vertices only when necessary to ensure that each tree edge is strictly heavier than its parent.
+Obtaining an equivalent graph in this manner always succeeds.
+Theorem˜21 Given an $\alpha$-acyclic hypergraph $H=(X,R,\chi)$, there exists an equivalent hypergraph $H^{*}=(X^{*},R,\chi^{*})$ over the same hyperedge set $R$ that
+  • admits a monotonic weight join tree $\hat{T}\in\mathcal{T}(H^{*})$, and
+  • $\mathcal{T}(H^{*})=\mathcal{T}(H)$.
 Proof.
-To prove the existence, we construct such a hypergraph $H^{*}$ by vertex duplication.
-Section˜4.2 guarantees that $\mathcal{T}(H^{*})=\mathcal{T}(H)$.
-We only need to show the existence of $\hat{T}\in\mathcal{T}(H^{*})$.
-Algorithm˜2 constructs $H^{*}$ by vertex duplication in a breadth-first manner from the root of $T$ to make it a monotonic weight join tree in $H^{*}$.
-Algorithm˜2 starts with enqueuing each edge incident to the root.
+To prove existence, we construct hypergraph $H^{*}$ by vertex duplication (algorithm˜6).
+section˜C.1 will guarantee that $\mathcal{T}(H^{*})=\mathcal{T}(H)$.
+Therefore, we only need to show the existence of monotonic weight join tree $\hat{T}\in\mathcal{T}(H^{*})$.
+algorithm˜6 takes as input hypergraph $H$ along with an MCS tree $T$ of $H$.
+The algorithm constructs $H^{*}$ by vertex duplication in a breadth-first manner from the root of $T$ to make it a monotonic weight join tree in $H^{*}$.
+algorithm˜6 starts with enqueuing each edge of $T$ incident to the root.
 Once a tree edge $e\in E(T_{r})$ is dequeued, we check for the number of duplications needed to make it heavier than its parent edge $p(e)$, namely $\Delta=\omega(p(e))-\omega(e)+1$.
 The duplication is performed only if $\Delta>0$.
-Section˜4.2 guarantees that $\chi(e)\setminus\chi(p(e))\neq\emptyset$.
+section˜3 guarantees that $\chi(e)\setminus\chi(p(e))\neq\emptyset$.
 Therefore, we can always duplicate a vertex $x\in\chi(e)\setminus\chi(p(e))$ for $\Delta$ times such that $\omega(e)>\omega(p(e))$.
-Let $e=(r_{u},r_{d})$ where $0pt(T,r_{u})<0pt(T,r_{d})$, $x$ only occurs in the subtree rooted at $r_{u}$.
+We prove that this produces a monotonic weight join tree by induction on a breadth-first ordering (with respect to root $r$) of the edges in $T$.
+Base case: The base case is an edge $e$ incident to the root. Since $e$ has no parent $\omega(e)>\omega(p(e))$ holds vacuously.
+Ind. step: Let $(e_{1},\ldots,e_{m})$ be a breadth-first ordering of the edges in $T$ and assume $\omega(e_{i})>\omega(p(e_{i}))$ for all i<k1i<k-1.
+Let $e_{k}=(r_{u},r_{d})$ where $0pt(T,r_{u})<0pt(T,r_{d})$ and let $x\in\chi(e_{k})\setminus\chi(p(e_{k}))$.
+Vertex $x$ only occurs in the subtree rooted at $r_{u}$.
 Otherwise, it violates the running intersection property.
-Therefore, the duplication of $x$ does not affect the weight of any edge closer to the root than $e$ in $T$.
-By Section˜4.2, duplication in a sibling edge $e^{\prime}\in s(e)$ would not affect the weight of $e$ either, and $\omega(e)>\omega(p(e))$ will continue to hold.
-The tree has monotonic weights when Algorithm˜2 terminates.
-∎
-Given a rooted join tree $T$ of a hypergraph $H$, a non-tree edge is $e=(r_{i},r_{j})\in E(L(H))\setminus E(T)$.
-There is a path in $T$ between $r_{i}$ and $r_{j}$ via their lowest common ancestor $\mathsf{LCA}(r_{i},r_{j})$.
-We define the LCA edges $\lambda(e)=\lambda(r_{i},r_{j})$ as a set of at most two tree edges on the path and incident to $\mathsf{LCA}(r_{i},r_{j})$.
-If $r_{i}$, $r_{j}$ are ancestor and child, then $|\lambda(e)|=1$, otherwise, $|\lambda(e)|=2$.
-In the process of building $G^{\equiv}(H^{*})$ from a monotonic weight join tree $\hat{T}$ of $H^{*}$, Section˜4.1 and Section˜4.2 guarantee that:
-  • no tree edge in $\hat{T}$ needs to be slid;
-  • each non-tree edge $e$ can be removed if it is lighter than all $e^{\prime}\in\lambda(e)$;
-  • otherwise, the non-tree edge $e$ can be slid directly to be incident to each $e^{\prime}\in\lambda(e)$.
-Algorithm˜3 constructs $G^{\equiv}(H^{*})$ from the line graph $L(H)$ weighted by $\omega$ and an MCS tree $T_{r}$.
-Full details on the constant-time operations on graphs and data structures to support operations performed in Algorithm˜3 can be found in Appendix˜B.
-As a part of pre-processing, we first conduct a breadth-first search on $\hat{T}_{r}$ to obtain the depth table $0pt$ of each tree node in $\mathcal{O}(|R(\hat{T}_{r})|)$.
-We will also use the algorithms proposed by Michael Bender [3, 4] to build the data structures in $\mathcal{O}(|R(\hat{T}_{r})|)$ to facilitate the constant-time query of lowest common ancestors (LCAs) and level ancestors (LAs).
-The constant-time query of depth, LCA and LA allows for finding the $\lambda(e)$ of any non-tree edge $e$ in constant time as shown from Line 3 to Line 3 in Algorithm˜3.
-Input: MCS tree $T_{r}(H)$ of line graph $L(H)$ with weight function $\omega$
-1 $G^{\equiv}\leftarrow L(H)$
-2for $e^{*}=(r_{i},r_{j})\in R(L(H))\setminus R(T_{r}(H))$ do
- // where $0pt(T_{r}(H),r_{i})\leq 0pt(T_{r}(H),r_{j})$
-3 $w^{*}=\omega(e^{*})$
-4 $l\leftarrow\mathsf{LCA}(r_{i},r_{j});d\leftarrow 0pt(l)+1$
-5 $e_{1}=(l,r_{1}\leftarrow\mathsf{LA}(r_{i},d));w_{1}=\omega(e_{1})$
-6 $e_{2}=(l,r_{2}\leftarrow\mathsf{LA}(r_{j},d));w_{2}=\omega(e_{2})$
-7 if $w^{*}<w_{1}\wedge w^{*}<w_{2}$ then
-8      delete $e^{*}$ from $R(G^{\equiv})$
-9
-10 if $e_{1}=e_{2}\wedge w^{*}=w_{1}$ then
-11      slide $r_{j}$ to $r_{1}$
-12
-13 else if $e_{1}\neq e_{2}\wedge w^{*}=w_{1}=w_{2}$ then
-14      slide $r_{i}$ to $r_{1}$ and $r_{j}$ to $r_{2}$
-15
-16 else if $e_{1}\neq e_{2}\wedge w^{*}=w_{1}<w_{2}$ then
-17      slide $r_{i}$ to $l$ and $r_{j}$ to $r_{1}$
-18
-19 else if $e_{1}\neq e_{2}\wedge w^{*}=w_{2}<w_{1}$ then
-20      slide $r_{i}$ to $l$ and $r_{j}$ to $r_{2}$
-21
-22return $G^{\equiv}$
-Algorithm 3 buildEGbuildEG
-(a) Monotonic Weight Join Tree $\hat{T}_{P}(H^{*}_{6})$
-(b) 3 Non-tree Edges Processed
-(c) Equivalent Graph $G^{\equiv}(H^{*}_{6})$
-Figure 4: Algorithm˜3 on $H^{*}_{6}$ of Figure˜4(a)
-Example 21.
-We consider $\hat{T}_{P}$ in Figure˜4(b) of an equivalent hypergraph $H^{*}_{6}$ in Figure˜4(a).
-Three of $\hat{T}_{P}$’s non-tree edges are shown in Figure˜5(a) where the non-MST edges $(Y,W),(Y,T)$ are dotted and the MST edge $(Y,S)$ is dashed.
-We can identify $\lambda(Y,W)=\{(T,W),(T,U)\}$ in constant time.
-Because $\omega(Y,W)=|\{a\}|=1$ is lighter than $\omega(T,W)=\omega(T,U)=|\{a,c\}|=2$, it is removed.
-$(Y,T)$ is also removed, for it is lighter than its LCA edge $(T,U)$.
-We compare $(Y,S)$ with $\lambda(Y,S)=\{(P,T),(P,S)\}$.
-Because they have the same weight with $\omega(Y,S)=\omega(P,T)=\omega(P,S)=|\{a\}|=1$, we slide $(Y,S)$ directly to $(T,S)$.
-The intermediate graph with those three non-tree edges processed is shown in Figure˜5(b).
-The resulting equivalent graph $G^{\equiv}(H^{*}_{6})$ is shown in Figure˜5(c).
-The next result follows immediately from the construction of the equivalent hypergraph.
+Therefore, the duplication of $x$ does not affect the weight of any edge $e_{i}$, $i\in[k-1]$ outside of subtree rooted at $r_{u}$ and therefore $\omega(e_{i})>\omega(p(e_{i}))$ holds under any number of duplications of $x$.
+For any $e_{i}$ lying within the subtree rooted at $r_{u}$, $e_{i}$ is necessarily a sibling of $e_{k}$ since edges of $T$ are breadth-first ordered.
+Hence $p(e_{i})$ lies outside this subtree and any number of vertex duplications of $x$ cannot decrease the weight of $e_{i}$ with respect to its parent.
+Therefore, $\omega(e_{i})>\omega(p(e_{i}))$ for all $i\in[k]$ after performing $\Delta$ many vertex duplications of $x$.
+Consequently, tree $T$ has monotonic weights when algorithm˜6 terminates. ∎
+The above theorem guarantees that for any $\alpha$-acyclic hypergraph we are able to obtain an equivalent hypergraph $H^{*}$.
+Next we prove Theorem˜24 showing that algorithm˜2 produces the equivalent graph of line graph $L(H^{*})$.
+Theorem˜24 Given a hypergraph $H^{*}=(X^{*},R,\chi^{*})$ with a monotonic weight join tree $\hat{T}$, algorithm˜2 produces an Equivalent Graph $G^{\equiv}(L(H^{*}))$.
+Proof.
+[…]
+Of note is the case where a non-tree edge $e$ weighs less than all edges in $\lambda(e)$.
+The algorithm then slides $e$ to a self loop at $l=\mathsf{LCA}(e)$.
+Such an edge may still be able to slide if edge $\{l,p(l)\}$ weighs more than edge $e$.
+In this sense graph $G$ returned by algorithm˜2 may not be maximal under partial order $\leq_{\hat{T}}$.
+Therefore it remains to show that these self loops do not prevent $G$ from being an equivalent graph.
+To see this, notice that no self loop can take part in a spanning tree.
+Therefore any spanning tree of graph obtained upon sliding self loops of $G$ to a fixpoint ($G^{\equiv}$) is also a spanning tree of $G$.
+Hence, $G$ and $G^{\equiv}$ have the same set of spanning trees and $G$ is an equivalent graph.
+The next result follows from the construction of the equivalent hypergraph.
 Lemma 48.
@@
 ∎
-Section˜4.2 allows us to produce $G^{\equiv}(H^{*})$ by keeping all tree edges of the MCS tree $T(H)$ in place and sliding the non-tree edges directly to their LCA edges.
-For example, running Algorithm˜1 on the MCS tree $T(H_{6})$ in Figure˜2(d) produces an identical $G^{\equiv}(H^{*}_{6})$ in Figure˜5(c).
-Theorem 23.
-Let $H$ be an $\alpha$-acyclic hypergraph with weight function $\omega$, line graph $L$ and MCS tree $T$.
-Let $H^{*}$ be the equivalent hypergraph of $H$ with weight function $\omega^{*}$.
-Applying Algorithm˜3 to inputs $L$, $T$ and $\omega$ produces the same equivalent graph as applying Algorithm˜3 to inputs $L$, $T$ and $\omega^{*}$.
+As section˜4 shows, running algorithm˜2 on the MCS tree $T(H_{6})$ in fig.˜2(d) produces $G^{\equiv}(H_{6})$ identical to $G^{\equiv}(H^{*}_{6})$ in fig.˜5(c).
+algorithm˜2 does this in time linear in the size of the line graph.
+We give the formal proof below.
+Theorem˜26 Given the line graph $L$ and an MCS tree $T_{r}$ of an $\alpha$-acyclic hypergraph $H$, algorithm˜2 returns an equivalent graph $G^{\equiv}(L)$ in time $\mathcal{O}(|L|)$.
 Proof.
-One can readily check that for any non-tree edge $e$ of $T$, the conditions of each if statement in Algorithm˜3 hold under $\omega$ if and only if they hold under $\omega^{*}$ by Section˜4.2. ∎
-Theorem 24.
-Given line graph $L$ of $\alpha$-acyclic hypergraph $H$ with weight function $\omega$ and an MCS tree $T$ of $H$, Algorithm˜3 returns equivalent graph $G^{\equiv}$ of equivalent hypergraph $H^{*}$ in $\mathcal{O}(m+n)$ where $n$ is the number of vertices and $m$ is the number of edges in $L$.
+Let $H^{*}=(X^{*},R,\chi^{*})$ be an equivalent hypergraph derived from $H=(X,R,\chi)$ by vertex duplication.
+One can readily check that for any non-tree edge $e$ of $T$, the conditions of each if statement in algorithm˜2 hold under weight function $\omega$ of $L(H)$ if and only if they hold under weight function $\omega^{*}$ of $L(H^{*})$ by section˜C.1.
+Therefore algorithm˜2 returns an equivalent graph $G^{\equiv}(L)$.
+The construction of each data structure $0pt(\cdot)$, $\mathsf{LA}(\cdot)$ and $\mathsf{LCA}(\cdot)$ can be done in $\mathcal{O}(m+n)$ [4, 3].
+Execution of the rest of the algorithm also takes time $\mathcal{O}(m+n)$, as it requires examining each non-tree edge to slide it directly to its LCA edges in the MCS tree.
+For each non-tree edge, the lookup of at most two LCA edges takes time $\mathcal{O}(1)$, as does the comparison of weights and the direct sliding transformation.
+The overall runtime is thus $\mathcal{O}(|L|)$.
+∎
+Given algorithm˜2, we now have a systematic way of enumerating join trees of an $\alpha$-acyclic hypergraph.
+Starting from the line graph $L(H)$ of an $\alpha$-acyclic hypergraph $H$, we first reconstruct $H$ and apply algorithm˜1 to obtain an MCS tree $T$.
+Using $L(H)$ and $T$ as input, algorithm˜2 constructs an equivalent graph of $L(H)$.
+Finally, a standard spanning tree enumeration algorithm can be applied to this equivalent graph to enumerate all join trees of $H$.
+The overall running time of this process is $\mathcal{O}(|L|+|\mathcal{T}(H)|)$.
+Theorem˜27 Given the line graph $L$ of an $\alpha$-acyclic hypergraph $H$, the join trees of $H$ can be enumerated in time $\mathcal{O}(|L|+|\mathcal{T}(H)|)$.
 Proof.
-The construction of each data structure $0pt(\cdot)$, $\mathsf{LA}(\cdot)$ and $\mathsf{LCA}(\cdot)$ can be done in $\mathcal{O}(m+n)$ [3, 4].
-Execution of the rest of the algorithm also takes time $\mathcal{O}(m+n)$, as it requires examining each non-tree edge to either delete it or slide it directly to its LCA edges in the MCS tree.
-For each non-tree edge, the lookup of at most two LCA edges takes time $\mathcal{O}(1)$, as does the comparison of weights and the direct sliding transformation.
+From $L(H)$ we first retrieve query hypergraph $H$.
+This is done in time linear in the size of the line graph.
+algorithm˜1 constructs an MCS tree $T$ in $\mathcal{O}(|H|)$.
+With the inputs $L(H)$ and $T$, we can run algorithm˜2 to construct $G^{\equiv}(H)$ in $\mathcal{O}(|L|)$ by Theorem˜26.
+Finally, letting $n$ be the number of vertices and $m$ the number of edges in $L(H)$, the spanning tree enumeration algorithm by Kapoor and Ramesh [24] on $G^{\equiv}(H)$ enumerates all join trees $|\mathcal{T}(H)|$ in $\mathcal{O}(m+n+|\mathcal{T}(H)|)=\mathcal{O}(|L(H)|+|\mathcal{T}(H)|)$.
+C.2 Missing Proofs in section˜4.3
+We prove Theorem˜28 below.
+Theorem˜28 The join trees of a $\gamma$-acyclic $H$ can be enumerated in time $\mathcal{O}(|H|+|\mathcal{T}(H)|)$.
+Proof.
+By modifying the algorithm˜1 to algorithm˜3, where changes are highlighted in red, we can construct a weighted MCS tree $T$ and an unweighted line graph $L(H)$ in $\mathcal{O}(|H|+|E(L(H))|)$.
+We simplify algorithm˜2 to algorithm˜4 where the changes are highlighted in red.
+As $\gamma$-acyclicity guarantees that every non-tree edge $(r_{i},r_{j})$ is an MST edge, we check its LCA $l$ and the weights of LCA edges $\omega_{1},\omega_{2}$ to determine how to slide it.
+If the shallower endpoint is the LCA of the deeper $r_{i}=l$, we slide it to be parallel with the only LCA edge as in Line 4.
+Otherwise, if the LCA edges have different weights, we slide $\{r_{i},r_{j}\}$ to be parallel to the heavier one as in Line 4 and Line 4.
+Finally if they have the same weight, we slide $\{r_{i},r_{j}\}$ to form a triangle with them as in Line 4.
+Those four cases are exhaustive.
+The number of join trees is lower bounded by the number of non-tree edges in its line graph, namely $|\mathcal{T}(H)|=\Omega(|E(L(H))|-|E(T)|)$.
+The total time complexity of enumerating all join trees is $\mathcal{O}(|H|+|E(L(H))|+|\mathcal{T}(H)|)=\mathcal{O}(|H|+|\mathcal{T}(H)|)$.
 ∎
-Once we have the equivalent graph $G^{\equiv}$, the existing algorithm allows for enumerating all its spanning trees with an amortized constant delay.
-The total time complexity of enumerating $|\mathcal{T}(H)|$ given the line graph $L(H)$ as input is $\mathcal{O}(|L(H)|+|\mathcal{T}(H)|)$.
-Starting from an $\alpha$-acyclic query $\mathcal{Q}$, Section˜C.1 describes how to convert it into a hypergraph $H$.
-Section˜C.2 describes how to convert it into a weighted line graph $L(H)$.
-Both conversions take linear time $\mathcal{O}(|\mathcal{Q}|)$ with respect to the query size as defined in Appendix˜C.
-Algorithm˜1 constructs an MCS tree $T$ in $\mathcal{O}(|H|)\leq\mathcal{O}(|\mathcal{Q}|)$.
-With the inputs $L(H)$ and $T$, we can run Algorithm˜3 to construct $G^{\equiv}(H^{*})$ in $\mathcal{O}(m+n)\leq\mathcal{O}(|\mathcal{Q}|)$ (where $n$ is the number of vertices and $m$ is the number of edges in $L(H)$).
-Finally, we apply the spanning tree enumeration algorithm by Kapoor and Ramesh [22] on $G^{\equiv}(H^{*})$ to enumerate all join trees $|\mathcal{T}(H)|$ in $\mathcal{O}(m+n+|\mathcal{T}(H)|)\leq\mathcal{O}(|\mathcal{Q}|+|\mathcal{T}(H)|)=\mathcal{O}(|L(H)|+|\mathcal{T}(H)|)$.
-4.3 $\gamma$-Acyclic Queries
-The run time of Algorithm˜3 depends on the size of the line graph which can be quadratically larger than the input hypergraph.
-But if $H$ is $\gamma$-acyclic, we can bring the total time complexity of enumeration down to $\mathcal{O}(|H|+|\mathcal{T}(H)|)$.
-The line graph of a $\gamma$-acyclic hypergraph $H$ is its union join graph [25] $L(H)=\mathcal{U}(H)$, i.e., every edge $e\in E(L(H))$ in the line graph is an MST edge.
-By modifying the MCS algorithm to Algorithm˜7 in Appendix˜F, we can construct a weighted MCS tree $T$ and an unweighted line graph $L(H)$, namely the union join graph, in $\mathcal{O}(|H|+|E(\mathcal{U}(H))|)$.
-We simplify Algorithm˜3 to Algorithm˜8 in Appendix˜F, which compares the LCA edges $\lambda(e)$ of each non-tree edge $e$ and slides it to construct the equivalent graph $G^{\equiv}(H)$ in $\mathcal{O}(|E(\mathcal{U}(H))|)$.
-As $\gamma$-acyclicity guarantees that every non-tree edge $e$ is an MST edge, we can use the weights of the LCA edges to determine how to slide $e$.
-Let $\lambda(e)=\{e_{l},e_{r}\}$.
-If $e_{l}=e_{r}$, we slide $e$ to be parallel to $e_{l}$.
-Otherwise, if $e_{l}$ and $e_{r}$ have different weights, we slide $e$ to be parallel to the heavier one.
-If they have the same weight, we slide $e$ to form a triangle with them.
-The number of join trees is lower bounded by the number of non-tree edges in its line graph, namely $|\mathcal{T}(H)|=\Omega(|E(\mathcal{U}(H))|-|E(T)|)$.
-The total time complexity of enumerating all join trees is $\mathcal{O}(|H|+|E(\mathcal{U}(H))|+|\mathcal{T}(H)|)=\mathcal{O}(|H|+|\mathcal{T}(H)|)$.
-5 The Canonical Join Tree of a Berge-Acyclic Query
-An acyclic query can have exponentially many join trees with respect to its size.
-For example, the line graph of a clique query with $n$ relations is an $n$-clique $K_{n}$ with $n^{n-2}$ join trees by Cayley’s formula [7].
-Enumerating all join trees can be prohibitive for large queries.
-On the other hand, the query optimizer does not need to consider all possible join trees to achieve good performance.
-For example, the implementation of Yannakakis’ algorithm by Zhao et al.
-[47] has similar performance on any join tree rooted at the largest relation.
-An alternative to enumeration is therefore to simply construct an arbitrary join tree for a given root.
-This can be done in linear time by the Maximum Cardinality Search (MCS) algorithm from a chosen relation as shown in Algorithm˜1.
-In this section, we show that the MCS algorithm produces a shallowest tree for a Berge-acyclic query, in the sense that the depth of every node in the tree is minimized.
-We prove that this shallowest tree is unique, and therefore call it the canonical join tree:
-Definition 25.
-A join tree $T_{r}$ rooted at $r$ is canonical if $0pt(T_{r},r_{i})\leq 0pt(T^{\prime}_{r},r_{i})$ for any other join tree $T^{\prime}_{r}$ rooted at $r$ and any $r_{i}\in R(T_{r})=R(T^{\prime}_{r})$.
-A shallow join tree has practical benefits.
-For example, the depth of the join tree determines the number of sequential steps required in a parallel join algorithm.
-A shallow join tree tends to be wide and have more leaves, allowing better utilization of indices.
-Although Berge-acyclicity was thought to be too restrictive when it was first introduced to database theory [13], we found it to be general enough to cover almost all acyclic queries encountered in the wild.
-As shown in Table˜1 of Appendix˜F, among 10 45410\,454 queries from five popular benchmarks, $9285$ are $\alpha$-acyclic, and only 8 of these are not Berge-acyclic.
-In retrospect, this should not be surprising, as most joins in relational databases are over primary/foreign keys.
-Emerging workloads in graph databases usually involve simple graphs and seldom require composite key joins.
-A query without composite key joins admits a linear hypergraph, where each pair of hyperedges shares at most one vertex.
-The following result establishes an equivalence between $\alpha$-acyclicity with linearity and Berge-acyclicity.
-Proposition 26.
-An $\alpha$-acyclic hypergraph is Berge-acyclic if and only if it is linear.
-By Section˜5, every edge in the line graph of a Berge-acyclic hypergraph is labeled with a single variable, thus having a weight of 1.
-Every spanning tree is a maximum spanning tree, therefore a join tree.
-Corollary 27.
-For a Berge-acyclic hypergraph $H$, any spanning tree of $L(H)$ is a join tree.
-In the rest of this section, we prove the existence and uniqueness of the canonical join tree rooted at any relation of a Berge-acyclic hypergraph, and show that it can be constructed by MCS as in Algorithm˜1.
-5.1 Existence & Uniqueness of the Canonical Join Tree
-The key insight leading to the existence and uniqueness of the canonical join tree is that the line graph of a Berge-acyclic hypergraph is geodetic [31], meaning that there is a unique shortest path between any pair of vertices.
-To prove this, we first review some relevant concepts and properties of graphs.
-All graphs below are simple and undirected.
-Definition 28.
-A cycle in a graph is a sequence of distinct vertices $v_{0},\ldots,v_{k-1\geq 2}$ such that there is an edge between $v_{i}$ and $v_{(i+1)\bmod k}$ for all $1\leq i\leq k$.
-A special class of graphs called chordal graphs are intimately related to acyclic hypergraphs:
-Definition 29.
-A graph is chordal if every cycle of length at least 4 has a chord, i.e., an edge that is not part of the cycle but connects two vertices of the cycle.
-The key step in this section is to show the line graph of a Berge-acyclic hypergraph is a block graph, which is a special class of chordal graphs.
-There are many different characterizations of block graphs [18], and we present a recent one:
-Definition 30 (Block Graph [10]).
-A graph $G=(R,E)$ with the vertex set $R$ and the edge set $E$ is a block graph if it is chordal and diamond-free.
-Removing an edge from a 4-clique $K_{4}$ results in a diamond.
-Diamond-free requires that no induced subgraph $G|_{R^{\prime}}$ is a diamond, where $R^{\prime}\subseteq R$ is any subset of vertices.
-We now present the main result relating block graphs to Berge-acyclic hypergraphs:
-Lemma 31.
-The line graph $L$ of a Berge-acyclic hypergraph $H$ is a block graph.
-Together with the fact that every block graph is geodetic (it has a unique shortest path between any two vertices) [31], Section˜5.1 implies the following corollary.
-Corollary 32.
-Let $L$ be the line graph of a Berge-acyclic hypergraph $H$. There is a unique shortest path between any two vertices in $L$.
-We are now ready to prove the existence and uniqueness of the canonical join tree.
-Theorem 33.
-Let $L$ be the line graph of a Berge-acyclic hypergraph $H$, and $\mathsf{P}(r,r^{\prime})$ be the shortest path in $L$ between $r,r^{\prime}\in R(L)$, $T_{r}=\bigcup_{r^{\prime}\in R(L)}\mathsf{P}(r,r^{\prime})$ is the unique canonical join tree for $H$ rooted at $r$.
+Appendix D Missing Proofs in section˜5
+section˜5 An $\alpha$-acyclic hypergraph is Berge-acyclic if and only if it is linear.
 Proof.
-A spanning tree of $L$ is its join tree by Section˜5. To show $T_{r}$ is a join tree, it is sufficient to show $T_{r}$ is spanning, connected and acyclic.
-$T_{r}$ is spanning and connected because it contains a path from $r$ to each $r^{\prime}\in R(L)$.
-We prove $T_{r}$ is acyclic by induction on the distance $\mathsf{dist}(r,r^{\prime})$ between $r$ and $r^{\prime}$.
-Let $R_{d}=\left\{r^{\prime}\in R(L)\mid\mathsf{dist}(r,r^{\prime})\leq d\right\}$.
-$R_{0}=\{r\}$ contains only the root.
-The subgraph $T_{r}|_{R_{0}}$ is trivially acyclic.
-Assuming that $T_{r}|_{R_{d>0}}$ is acyclic, we consider a vertex $r^{\prime}\in R_{d+1}\setminus R_{d}$.
-Section˜5.1 guarantees a unique shortest path between each pair of vertices $r,r^{\prime}\in R(L)$.
-Each $r^{\prime}$ is connected to a unique neighbor $r^{\prime\prime}\in R_{d}\setminus R_{d-1}$ that is at distance $d$ from $r$.
-Otherwise, there are at least two distinct shortest paths from $r$ to $r^{\prime}$.
-Therefore $T_{r}|_{R_{d+1}}$ is acyclic, and $T_{r}$ is a join tree.
-The join tree $T_{r}$ is canonical, because the path from $r$ to each $r^{\prime}\in R(L)$ is the shortest and therefore minimizing the depth $0pt(T_{r},r^{\prime})$.
-The canonical tree is unique by Section˜5.1.
+Fagin [15] showed that every Berge-acyclic hypergraph is linear.
+We prove the other direction by contradiction, assuming that $H$ is $\alpha$-acyclic and linear.
+We also assume that there is a Berge cycle $(r_{0},x_{0},\dots,r_{k-1},x_{k-1})$ of length $k\geq 2$.
+Without loss of generality, we suppose $r_{1}$ is the first hyperedge on the cycle to be removed by GYO reduction.
+Then $r_{1}$’s parent $r_{p}$ must satisfy $r_{p}\cap r_{1}\supseteq(r_{1}\cap\bigcup_{i\in[k]\wedge i\neq 1}r_{i})=\{x_{0},x_{1}\}$, contradicting $H$’s linearity.
 ∎
-5.2 Construction of the Canonical Join Tree
-Given a Berge-acyclic hypergraph $H$ and $r\in R(H)$ as the root, we can construct the canonical tree $T_{r}(H)$ by running Algorithm˜1.
-We first present a useful lemma on the structure of the canonical join tree:
+section˜5 The line graph $L$ of a Berge-acyclic hypergraph $H$ is a block graph.
+Proof.
+A Berge-acyclic hypergraph is $\gamma$-acyclic [15].
+Zhu proves that the line graph of a $\gamma$-acyclic hypergraph is chordal [49].
+We now prove that the line graph is diamond-free by showing that every diamond must be part of a $K_{4}$.
+Assuming a diamond $(r_{1},r_{2},r_{3},r_{4})$ in $L$ without the edge $(r_{1},r_{3})$, it can be treated as two triangles $(r_{1},r_{2},r_{4})$ and $(r_{2},r_{4},r_{3})$ sharing an edge $(r_{2},r_{4})$.
+We now consider the possible edge labels, $r_{i}\cap r_{j}$, in each triangle:
+Case 1:
+If all edges are labeled with distinct variables, each triangle forms a Berge-cycle, contradicting the Berge-acyclicity.
+Case 2:
+If two edges are labeled with the same variable, all three vertices share that variable. The third edge must be labeled with the same variable.
+Since the two triangles share the edge $(r_{2},r_{4})$, all their edges are labeled with the same variable $x$, namely $x\in r_{1}\cap r_{2}\cap r_{3}\cap r_{4}$.
+There has to be an edge $(r_{1},r_{3})\in E(L)$, contradicting the assumption that $(r_{1},r_{2},r_{3},r_{4})$ is a diamond.
+∎
 Lemma 49.
 Let $T_{r}$ be the canonical join tree rooted at $r$ of a Berge-acyclic hypergraph. Along its root-to-leaf path $r,r_{1},\ldots,r_{k}$, each pair of adjacent vertices shares a distinct variable.
-We now show that the MCS algorithm constructs the canonical join tree.
-Theorem 35.
-Given a Berge-acyclic hypergraph $H$, Algorithm˜1 from $r\in R(H)$ constructs the canonical tree $T_{r}(H)$.
 Proof.
-Algorithm˜1 labels each hyperedge in $R(H)$ in ascending order from the root $r_{1}=r$ to $r_{n=|R(H)|}$.
-When Algorithm˜1 has labeled $i\in[n]$ hyperedges, we consider a graph $G_{i}$ whose vertices are the labeled hyperedges $\{r_{1},\ldots,r_{i}\}$ and edges are the parent-child relationships $\{(r_{2},p(r_{2})),\ldots,(r_{i},p(r_{i}))\}$.
+Let $r_{0}=r$, section˜5 requires that $|r_{i-1}\cap r_{i}|=1$ for all $i\in[k]$.
+By Theorem˜35, the path $r_{0},\ldots,r_{k}$ is the shortest path from $r_{0}$ to $r_{k}$.
+The variable shared by two adjacent vertices $r_{i-1},r_{i}$ is the edge label $r_{i-1}\cap r_{i}$.
+The edge labels along the path are all distinct.
+Otherwise, two consecutive edges with the same label allow for shortening the path by removing either, contradicting the shortest path property.
+Two non-consecutive edges with the same label violate the running intersection property.
+∎
+Theorem˜36 Given a Berge-acyclic hypergraph $H$, algorithm˜1 from $r\in R(H)$ constructs the canonical tree $T_{r}(H)$.
+Proof.
+Denoting the edges labeled by algorithm˜1 as $r=r_{1},\dots,r_{n=|R(H)|}$, when algorithm˜1 has labeled the first $i\in[n]$ hyperedges, we consider a graph $G_{i}$ whose vertices are the labeled hyperedges $\{r_{1},\ldots,r_{i}\}$ and edges are the parent-child relationships $\{\{r_{2},p(r_{2})\},\ldots,\{r_{i},p(r_{i})\}\}$.
 To prove the claim by induction, we show a loop invariant that $G_{i}$ is a subtree of $T_{r}$ for all $i\in[n]$.
@@
 Let us consider the possible relationships between $r_{p}$ and $r_{q}$.
-First, $r_{q}$ cannot be a descendant of $r_{p}$, otherwise there would be a shorter path from $r_{k+1}$ to the root $r$ going through $r_{p}$ instead of $r_{q}$, violating Theorem˜33.
+First, $r_{q}$ cannot be a descendant of $r_{p}$, otherwise there would be a shorter path from $r_{k+1}$ to the root $r$ going through $r_{p}$ instead of $r_{q}$, violating Theorem˜35.
 Suppose $r_{q}$ is an ancestor of $r_{p}$.
@@
 Denote their lowest common ancestor in $T_{r}$ as $r_{a}=\mathsf{LCA}(r_{p},r_{q})$.
-Let $\lambda(r_{p},r_{q})=\{e_{l},e_{r}\}$ as shown by the solid lines in Figure˜5 of Appendix˜F, $\mathsf{P}_{p}$ be the path from $r_{a}$ to $r_{p}$ excluding $e_{l}$ and $\mathsf{P}_{q}$ be the path from $r_{a}$ to $r_{q}$ excluding $e_{r}$ as shown by the dashed lines in Figure˜5.
-By Section˜5.2, the variables on $\mathsf{P}_{p}$ and $\mathsf{P}_{q}$ are all distinct.
-Consider the following possible cases:
+Let $\lambda(r_{p},r_{q})=\{e_{l},e_{r}\}$ as shown by the solid lines in fig.˜7, $\mathsf{P}_{p}$ be the path from $r_{a}$ to $r_{p}$ excluding $e_{l}$ and $\mathsf{P}_{q}$ be the path from $r_{a}$ to $r_{q}$ excluding $e_{r}$ as shown by the dashed lines in fig.˜7.
+Figure 7: Possible cases for how the relation $r_{k+1}$ attaches to the partial join tree $G_{k}$ during the construction of the canonical join tree by Maximum Cardinality Search.
+By appendix˜D, the variables on $\mathsf{P}_{p}$ and $\mathsf{P}_{q}$ are all distinct.
+Consider the following possible cases where $\chi^{l}=\bigcap_{r\in\rho(e_{l})}{r}$ and $\chi^{r}=\bigcap_{r\in\rho(e_{r})}{r}$:
 Case 1:
-If $r_{k+1}\cap r_{p}=r_{k+1}\cap r_{q}$ and $\chi(e_{l})=\chi(e_{r})$, then $\mathsf{P}_{p},\mathsf{P}_{q}$ form a Berge-cycle.
+If $r_{k+1}\cap r_{p}=r_{k+1}\cap r_{q}$ and $\chi^{l}=\chi^{r}$, then $\mathsf{P}_{p},\mathsf{P}_{q}$ form a Berge-cycle.
 Case 2:
-If $r_{k+1}\cap r_{p}=r_{k+1}\cap r_{q}$ and $\chi(e_{l})\neq\chi(e_{r})$, then $\mathsf{P}_{p},r_{a},\mathsf{P}_{q}$ form a Berge-cycle.
+If $r_{k+1}\cap r_{p}=r_{k+1}\cap r_{q}$ and $\chi^{l}\neq\chi^{r}$, then $\mathsf{P}_{p},r_{a},\mathsf{P}_{q}$ form a Berge-cycle.
 Case 3:
-If $r_{k+1}\cap r_{p}\neq r_{k+1}\cap r_{q}$ and $\chi(e_{l})=\chi(e_{r})$, then $\mathsf{P}_{p},\mathsf{P}_{q},r_{k+1}$ form a Berge-cycle.
+If $r_{k+1}\cap r_{p}\neq r_{k+1}\cap r_{q}$ and $\chi^{l}=\chi^{r}$, then $\mathsf{P}_{p},\mathsf{P}_{q},r_{k+1}$ form a Berge-cycle.
 Case 4:
-If $r_{k+1}\cap r_{p}\neq r_{k+1}\cap r_{q}$ and $\chi(e_{l})\neq\chi(e_{r})$, then $\mathsf{P}_{p},r_{a},\mathsf{P}_{q},r_{k+1}$ form a Berge-cycle.
+If $r_{k+1}\cap r_{p}\neq r_{k+1}\cap r_{q}$ and $\chi^{l}\neq\chi^{r}$, then $\mathsf{P}_{p},r_{a},\mathsf{P}_{q},r_{k+1}$ form a Berge-cycle.
 All possible cases above contradict Berge-acyclicity, therefore, algorithm˜1 must correctly assign $r_{p}$ as the parent of $r_{k+1}$. ∎
-6 Converting a Binary Join Plan to a Join Tree
-A recent approach [20, 2] that converts a binary join plan into a join tree has gained popularity as it allows system builders to leverage existing query optimizers designed for binary join plans.
-In this section, we focus on an algorithm by Hu et al.
-[20] that converts left-deep linear join plans into join trees.
-The algorithm can convert all acyclic queries in standard benchmarks, including those in Table˜1 of Appendix˜F.
-We prove that the algorithm converts any connected left-deep linear join plan into a join tree if and only if the query is $\gamma$-acyclic.
-Our result yields a new characterization of $\gamma$-acyclic queries.
-We formally define binary join plans and describe the algorithm by Hu et al.
-[20] with Algorithm˜6 in Appendix˜F.
-Definition 36.
-A left-deep linear plan is a sequence of relations $r_{1},r_{2},\ldots,r_{n}$. It is connected if for each $r_{i\geq 2}$, $\exists r_{j<i}:r_{i}\cap r_{j}\neq\emptyset$.
-Query optimizers strive to produce connected plans, to avoid expensive Cartesian products.
-Many optimizers produce exclusively left-deep linear plans.
-Plans that are not left-deep are called bushy, and such plans may still be decomposed into left-deep fragments [43].
-Given a left-deep linear plan, Hu et al.
-[20] generate a join tree with Algorithm˜6 in Appendix˜F.
-The algorithm chooses the first relation $r_{1}$ as the root and iterates through the rest of the plan.
-For each relation $r_{i\in[2,n]}$, it finds the first relation $r_{j}$ that contains all attributes shared by $r_{i}$ with all previous relations $r_{i}\cap\bigcup r_{k<i}$, and assigns $r_{j}$ as the parent of $r_{i}$.
-The algorithm constructs a join tree if it finds a parent for each $r_{i}$.
-Hu et al.
-[20] proved that the algorithm succeeds whenever the input plan is the reverse of a GYO-reduction order.
-They also observed that every left-deep linear plan produced for queries in standard benchmarks is indeed the reverse of a GYO-reduction order.
-This is not a coincidence, as we show that every connected left-deep linear join plan must be the reverse of a GYO-reduction order if and only if the query is $\gamma$-acyclic.
-Theorem 37.
-A query is $\gamma$-acyclic if and only if every connected left-deep linear join plan for the query is the reverse of a GYO-reduction order.
-The “only if” direction is more difficult.
-We prove its contrapositive by constructing a $\gamma$-cycle from a query plan that is not the reverse of a GYO-reduction order.
-We start with a simple observation that directly follows from the definition of GYO-reduction order.
+Appendix E Missing Algorithm and Proofs in section˜6
+Theorem˜38 A query is $\gamma$-acyclic if and only if every connected left-deep linear join plan for the query is the reverse of a GYO-reduction order.
+We first prove “only if” direction of Theorem˜38 by contradiction, assuming a query plan with a $\gamma$-cycle that is not the reverse of a GYO-reduction order.
 Proposition 50.
@@
 We say that $\hat{x}$ is a dangling variable of $\hat{r}$ and $\bar{x}$ is a dangling variable of $\bar{r}$.
-We are now ready to prove Theorem˜37.
+We are now ready to prove Theorem˜38.
 Our strategy is to derive a $\gamma$-cycle from any orphan relation, leading to a contradiction.
-Proof of Theorem˜37.
+Proof of Theorem˜38.
 To prove the “only if” direction by contradiction, we assume that a connected left-deep linear plan for a $\gamma$-acyclic query is not the reverse of a GYO-reduction order.
-By Section˜6, there is an orphan $\mathring{r}$ in the plan.
+By appendix˜E, there is an orphan $\mathring{r}$ in the plan.
 Assuming that $\mathring{r}$ is the $t$-th relation in the plan, we now consider the prefix of the plan up to $\mathring{r}$, namely $\mathsf{pre}({t})=r_{1},\ldots,r_{t-1},\mathring{r}$.
-Let $\hat{r}\in\mathsf{pre}({t-1})$ be a maximal relation with respect to $\geq_{\mathring{r}}$, Section˜6 guarantees an $\bar{r}\in\mathsf{pre}({t-1})$ incomparable with $\hat{r}$.
+Let $\hat{r}\in\mathsf{pre}({t-1})$ be a maximal relation with respect to $\geq_{\mathring{r}}$, appendix˜E guarantees an $\bar{r}\in\mathsf{pre}({t-1})$ incomparable with $\hat{r}$.
 There are at least two dangling variables $\bar{x}\in\bar{r}\cap\mathring{r}\setminus\hat{r}$ and $\hat{x}\in\hat{r}\cap\mathring{r}\setminus\bar{r}$.
@@
 ∎
-7 Conclusion and Future Work
-We proposed three approaches for constructing join trees.
-Our enumeration algorithm in Section˜4 generates join trees by edits with amortized constant delay, forming a basis for cost-based optimization of acyclic joins.
-In Section˜5, we showed that the Maximum Cardinality Search algorithm constructs the unique shallowest join tree for any Berge-acyclic query, supporting very large queries and enabling parallel execution.
-Finally, in Section˜6, we gave theoretical justification for converting binary join plans to join trees, allowing reuse of existing optimization infrastructure.
-Future work includes compact representations of join trees for dynamic programming, as in binary plan optimizers, and the challenging cost estimation for Yannakakis-style algorithms: the random-walk approach [26] models joint probabilities for binary joins, whereas an efficient and accurate solution for semijoins remains to be found.
-Our results also raise theoretical questions.
-Can join tree enumeration achieve worst-case constant delay?
-We proved Berge-acyclicity sufficient for the existence and uniqueness of the canonical join tree, but it is not necessary, and $\gamma$-acyclicity is insufficient as shown in Figure˜6 of Appendix˜F.
-What is the precise characterization of hypergraphs that admit a unique canonical join tree for any root, or for some root?
-Appendix A Extended Preliminaries
-Definition 41 (Hypergraph Acyclicity).
-A hypergraph $H$ is:
-  • $\alpha$-acyclic if it admits a join tree;
-  • $\beta$-acyclic if every subgraph of $H$ is $\alpha$-acyclic;
-  • $\gamma$-acyclic if it does not contain any $\gamma$ cycle.
-A $\gamma$ cycle is a sequence of length $k\geq 3$ of distinct hyperedges and distinct vertices $(r_{0},x_{0},\dots,r_{k-1},x_{k-1})$ such that every $x_{i\in[0,k-2]}$ belongs to $r_{i}\cap r_{i+1}$ and no other $r_{j}$ while $x_{k-1}$ belongs to $r_{k-1}\cap r_{0}$ and possibly other hyperedges;
-  • Berge-acyclic if it does not contain any Berge cycle.
-A Berge cycle is a sequence of length $k\geq 2$ of distinct vertices and distinct hyperedges $(r_{0},x_{0},\dots,r_{k-1},x_{k-1})$ such that $\forall\ i\in[k]:x_{i}\in r_{i}\cap r_{(i+1)\bmod k}$.
-The four notions of acyclicity form a strict hierarchy that Berge-acyclic $\Rightarrow$ $\gamma$-acyclic $\Rightarrow$ $\beta$-acyclic $\Rightarrow$ $\alpha$-acyclic [13].
-Given a hypergraph $H$ with $n$ hyperedges, Algorithm˜1 labels each hyperedge with an integer $i\in[1,n]$ in ascending order and assigns a parent $p(r_{i})$ to each hyperedge $r_{i}$.
-Upon labeling an edge the algorithm marks all previously unmarked vertices contained in that edge.
-Initially, all hyperedges are unlabeled and all vertices are unmarked.
-The algorithm begins by assigning nullnull as the parent of the root $r$ and initializing label $i$ to 0.
-In each following while loop iteration, the algorithm increments $i$ by one and labels the current hyperedge as $r_{i}$.
-For each unmarked vertex $x$ in $r_{i}$ (a vertex is considered marked when it is removed from $X$), it assigns $r_{i}$ tentatively as the parent of each unlabeled hyperedge $r^{\prime}$ containing $x$.
-The algorithm marks all unmarked vertices in $r_{i}$ and removes $r_{i}$ from future consideration.
-Finally, it selects a hyperedge with the most marked vertices to label next, breaking ties arbitrarily.
-The algorithm terminates when all hyperedges are labeled.
-Tarjan and Yannakakis [40] describes an efficient implementation of Algorithm˜1 running in $\mathcal{O}(|H|)$.
-Example 42.
-Considering the hypergraph $H_{6}$ shown in Figure˜2(a), we run Algorithm˜1 from the hyperedge $P$.
-It is labeled as $r_{1}$ and all its vertices are marked.
-$P$ is also tentatively assigned as the parent of five hyperedges such that $p(S)=p(T)=p(U)=p(W)=p(Y)=r_{1}$.
-We proceed to label the next hyperedge with the most marked vertices.
-At this point, each unlabeled hyperedge has one marked vertex, so we can choose any of them, say $S$, and label it as $r_{2}$.
-Because the labeling of $S$ marks no additional vertices in $T,U,W,Y$, their parents remain unchanged.
-These remaining hyperedges are again tied with the same number of marked vertices, so we can choose any of them, say $T$, and label it as $r_{3}$.
-Since the labeling of $T$ marks a vertex $c$ in $U,W$, their parents are updated to $p(U)=p(W)=r_{3}$.
-We continue the process until all hyperedges are labeled and all vertices are marked. By connecting each labeled hyperedge with its parent, we construct the join tree $T^{M}$ shown in Figure˜2(d).
-Notice that Algorithm˜1 can generate different join trees for the same hypergraph by breaking ties differently.
-However, it does not guarantee to generate all possible trees.
-For example, it never generates the Join Tree $T^{G}$ shown in Figure˜2(c).
-Appendix B Data Structures for Constant Time Operations on Graphs
-Given line graph $L$ and MCS tree $T$, Algorithm˜3 iterates over non-tree edges of $L$.
-Moreover, the algorithm uses data structures $0pt(\cdot)$, $\mathsf{LA}(\cdot)$ and $\mathsf{LCA}(\cdot)$, in determining the endpoints of a non-tree edge’s LCA edges.
-In this section we describe briefly how iterating over non-tree edges can be achieved in time linear in the number of vertices and edges of $L$ and how determining the endpoints of a non-tree edge’s LCA edges can be achieved in constant time.
-Given a graph $G=(V,E)$, we assume the vertices are represented by integers from 1 to $|V|$.
-Given a tree $T$ of $G$ rooted at node $r$, recall that data structure $0pt_{T}(\cdot)$, stores for each node $r_{i}$ the length of the shortest path between $r$ and $r_{i}$ in $T$.
-Data structure $\mathsf{LCA}_{T}(\cdot)$, stores for each pair of nodes $r_{i},r_{j}$ the lowest common ancestor of $r_{i},r_{j}$ in $T$.
-Finally, data structure $\mathsf{LA}_{T}(\cdot)$ stores for each node-integer pair $r_{i},j$, the ancestor of node $r_{i}$ at depth $j$.
-Each of these data structures can be established from a tree $T$ in time linear in the size of the tree and support constant time look-ups.
-Observation 43.
-Given rooted tree $T$, data structures $0pt(\cdot)$, $\mathsf{LCA}(\cdot)$ and $\mathsf{LA}(\cdot)$ can be established in time $\mathcal{O}(|T|)$ and support $\mathcal{O}(1)$ time look-ups [9, 33, 3, 4].
-Given the line graph $L$ of an $\alpha$-acyclic hypergraph $H$ and an MCS tree $T$ of $H$, recall that the LCA edges $\lambda(e)$ of a non-tree edge $e=(r_{i},r_{j})$ of $T$ is defined as the set of edges incident to $\mathsf{LCA}(r_{i},r_{j})$ that lie on the unique path between $r_{i}$ and $r_{j}$ in $T$.
-The following observation shows that an edges LCA edges can be found in constant time given data structures $0pt(\cdot)$, $\mathsf{LCA}(\cdot)$ and $\mathsf{LA}(\cdot)$.
-Observation 44.
-Given the line graph $L$ of an $\alpha$-acyclic hypergraph $H$, an MCS tree $T$ and data structures $0pt(\cdot)$, $\mathsf{LCA}(\cdot)$ and $\mathsf{LA}(\cdot)$, the LCA edges of any non-tree edge can be found in constant time.
-Proof.
-Given a non-tree edge $e=(r_{i},r_{j})$ we find $\mathsf{LCA}(r_{i},r_{j})=l$ in constant time by Appendix˜B.
-The depth $d=0pt(l)+1$ is likewise found in constant time.
-The edge $e_{1}=(l,r_{1})$, where node $r_{1}=\mathsf{LA}(r_{i},d)$ is found by a constant time look-up, is adjacent to $l$ and lies on the unique path between $r_{i}$ and $r_{j}$.
-The same holds for edge $e_{2}=(l,r_{2})$ where node $r_{1}=\mathsf{LA}(r_{j},d)$.
-∎
-In what remains of this section we describe how Algorithm˜3 iterates over and performs sliding transformations on non-tree edges.
-We use an array $L$ to represent the adjacency list of a line graph where entry $L[i]$ consists of a linked list of tuples $(j,\omega)$ where $j$ is a neighbor of $i$ and $\omega$ is the weight of the edge $(i,j)$.
-A rooted tree $T_{r}$ is likewise implemented as an array, where the index $i$ represents the vertex $i$, and the $i$-th entry $T_{r}[i]$ stores $(p,\omega)$ where $p$ is the parent of $i$ and $\omega$ is the weight of the edge $(i,p)$.
-The depth table of the tree is also implemented as an array to enable constant time lookup for the depth of a vertex $0pt(T_{r},i)$.
-Appendix˜C shows how to construct the weighted line graph and the hypergraph given the original query.
-Then, using the hypergraph we can construct $T_{r}$ and $0pt$ during the MCS algorithm in linear time [40].
-As we iterate over each edge $(j,\omega)$ of $L[i]$, we check if the current edge $(i,j)$ is a tree edge by performing constant-time query of $T_{r}[i]$ and $T_{r}[j]$ to see if one of them is the parent of the other.
-From Appendix˜B we know that we can find the LCA edges $\lambda(i,j)=\{e_{1},e_{2}\}$ in constant time.
-There’s no general solution to lookup an entry in a linked list in constant time.
-Instead of iterating over the linked-list representation of $L$ to find the edge weights of $e_{1}$ and $e_{2}$, we can lookup their weights from the array representation of $T_{r}$ in constant time.
-Let $e_{1}=(r_{1},p(r_{1}))$ and $e_{2}=(r_{2},p(r_{2}))$ with weights $\omega_{1}$ and $\omega_{2}$.
-Because both $e_{1}$ and $e_{2}$ are tree edges, we can find their weights by querying $T_{r}[r_{1}]=(p(r_{1}),\omega_{1})$ and $T_{r}[r_{2}]=(p(r_{2}),\omega_{2})$.
-Each of these operations takes constant time.
-When we need to delete an edge $e^{*}=(i,j)$ from $L$, we first check if $e^{*}$ is a tree edge by checking $T_{r}[i]$ and $T_{r}[j]$.
-Deleting $e^{*}$ from the linked list takes constant time.
-Sliding an edge involves updating the endpoints of the edge which takes constant time.
-Appendix C From Queries to (Hyper-)Graphs
-Algorithm˜3 assumes the weighted line graph and an initial MCS tree are given as input. In practice, we need to construct these graphs from the original SQL query.
-We model a query involving $m$ relations as a set of join predicates each in the form of $R_{a}.x_{i}=R_{b}.x_{j}$ such that $a\neq b\wedge a,b\in[m]$, namely no self-joins or duplicates, and assume each relation is uniquely identified by an integer in $[m]$.
-For each relation $R_{a}$ with arity $k_{a}$, each attribute $x_{i\in[k_{a}]}$ is uniquely identified by an integer in $k_{a}]$.
-Therefore each predicate is uniquely identified by a tuple $(a,i,b,j)$.
-Each query $Q$ is represented by a set of such tuples.
-We will assume that query $Q$ is transatively closed over join predicates, i.e.
-if $R_{a}.x_{i}=R_{b}.x_{j}$ and $R_{b}.x_{j}=R_{c}.x_{k}$ are in $Q$, then $R_{a}.x_{i}=R_{c}.x_{k}$ is also in $Q$.
-The tuple $(a,i,b,j)$ is present in $Q$ if and only if $R_{a}.x_{i}=R_{b}.x_{j}$ is a join predicate over relations $R_{a},R_{b}$.
-We define the size of a query $|Q|$ as the number of join predicates in $Q$, equivalent to the number of $(a,i,b,j)$ tuples.
-Our goal is to construct the query hypergraph and the weighted line graph from this set of tuples.
-C.1 Construction of Hypergraph
-We construct the hypergraph $H$ by finding connected components of the predicate graph, which is a graph where there is a vertex for each attribute $R_{a}.x_{i}$, and there is an edge between $R_{a}.x_{i}$ and $R_{b}.x_{j}$ for every predicate $R_{a}.x_{i}=R_{b}.x_{j}$.
-Each connected component of the predicate graph corresponds to an “equivalence class” of attributes.
-We then assign a unique variable to each connected component, and map each attribute to the variable of its component.
-This can be done in linear time using standard algorithms for computing connected components [9, 39].
-Algorithm˜4 generates hypergraph $H$ from query $Q$.
-$H$ is represented as an array of size $m$ where each entry is a list of integers.
-The algorithm starts by initializing $H$ to an array of size $m$ and $P$ to an array of size $p$, where $p$ is the number of distinct relation-attribute pairs $R_{a}.x_{i}$ in $Q$ as shown by Lines 4-4.
-If numbers $m$ and $p$ are not known a priori they can be found in a constant number of passes over elements in $Q$.
-Array $H$ will serve as a hyperedge list of the query hypergraph and array $P$ will serve as an adjacency list representation of the query predicate graph.
-The subsequent for-loop fills out adjacency list $P$ at Lines 4-4.
-We assume here that each $(a,i)$ pair is mapped to a unique integer in $[p]$ by a perfect hash function $f$ at Line 4, otherwise such a mapping can be established in a constant number of passes over $Q$.
-Next, a standard algorithm [9, 39] for the enumeration of connected components of a graph is called in Line 4.
-This produces array $\mathcal{C}$ where each entry is a list of vertices that forms a connected component in $P$.
-The subsequent nested for-loop iterates over each relation in each component to fill out hyperedge list $H$ at Lines 4-4.
-Input: Query $Q$
-Output: Query hypergraph $H$
-1
-2for $i\in[1\ldots m]$ do
-3 $H[i]\leftarrow\texttt{linkedList}()$
-4
-5for $i\in[1\ldots p]$ do
-6 $P[i]\leftarrow\texttt{linkedList}()$
-7
-8for $(a,i,b,j)\in Q$ do
-9 $p=f(a,i)$
-10 $P[p].\text{append}(b,j)$
-[…]
-12
-13for $i\in[|\mathcal{C}|]$ do
-14 for $R_{a}\in\mathcal{C}[i]$ do
-15    $H[a].\text{append}(i)$
-16
-17
-return $H$
-Algorithm 4 Generating query hypergraph $H$ from query $Q$.
-Observation 45.
-Algorithm˜4 generates the hypergraph of $Q$ in time $\mathcal{O}(|Q|)$.
-Proof.
-Arrays $H$ and $P$ can be initialized in time $\mathcal{O}(|Q|)$ since $m\leq|Q|$ and $p\leq|Q|$.
-The subsequent for-loop iterates over all elements in $|Q|$ making a single insertion to $P$ thus performing at most $\mathcal{O}(|Q|)$ steps.
-Predicate graph $P$ therefore has size at most $\mathcal{O}(|Q|)$.
-Generating a list of connected components of predicate graph $P$ is done in time linear in $|P|=\mathcal{O}(|Q|)$ by standard algorithms [9, 39].
-The subsequent nested for-loop iterates over each relation in each component performing a single insertion on each iteration.
-The sum of the sizes of each component is exactly number of vertices of $P$ and therefore $\mathcal{O}(|Q|)$ elementary operations are performed.
-∎
-C.2 Construction of Weighted Line Graph
-To construct the weighted line graph, we will ignore the attribute indices $(i,j)$ and focus on the relation indices $(a,b)$ for each join predicate in $Q$.
-The set of predicates can then be seen as representing the edge list of a multigraph.
-Algorithm˜5 generates the weighted line graph of $Q$ by converting this multigraph into a weighted graph $L$ where the weight of each edge $(u,v)$ is the number of parallel edges between $u$ and $v$ in the multigraph.
-The algorithm first initializes arrays $L$ and $M$ both of size $m$, where $M$ will be an adjacency list representation of a multigraph and $L$ will be adjacency list representation of the final weighted line graph in Line 5-5.
-Auxiliary array weight is initialized in Line 5.
-Entries of weight are initialized to be an empty linked list.
-In the subsequent for-loop, "edges" in $Q$ are inserted into multigraph $M$ in Line 5.
-Array weight[j] stores the weight of the edge $(i,j)$ for the current $i$.
-Entries of weight[j] are initialized to 0.
-The subsequent for-loop iterates over adjacency list $M$ in Line 5.
-The first nested for-loop at Line 5 iterates over each neighbor $j$ of $i$ stored in $M[i]$.
-For each neighbor $j$, array weight[j] is incremented by 1.
-Array weight[j] now contains the weights of all edges adjacent to $i$ at its non-zero entries.
-The second nested for-loop from Line 5 to Line 5 iterates over each neighbor $j$ one more time and inserts neighbor and weight pairs into array $L$ if weight[j] is not 0, otherwise skipping to the next neighbor.
-After an insertion is made, weight[j] is set to 0.
-A weight of 0 at weight[j] indicates the edge at $(i,j)$ and its weight has already been outputted.
-This prevents the insertion of duplicate edges and resets array weight[j] to be reused for the next relation in $M$.
-Since each array operation takes constant time and we visit each edge a constant number of times, the overall complexity of this procedure is linear in the size of the input.
-Input: Query $Q$
-Output: Weighted line graph $L$
-1
-2for $i\in[1\ldots m]$ do
-3 $L[i]\leftarrow\texttt{linkedList}()$
-4 $M[i]\leftarrow\texttt{linkedList}()$
-5 $\texttt{weight}[i]\leftarrow 0$
-6
-7for $(a,i,b,j)\in Q$ do
-8 $M[a].\text{append}(b)$
-9
-10for $R_{i}\in M$ do
-11 for $R_{j}\in M[i]$ do
-12    $\texttt{weight}[j]\leftarrow\texttt{weight}[j]+1$
-13 for $R_{j}\in M[i]$ do
-14    if $\texttt{weight}[j]\neq 0$ then
-15       $L[i].\text{append}(R_{j}.\texttt{weight}[j])$
-16       $\texttt{weight}[j]\leftarrow 0$
-17
-18
-19
-return $L$
-Algorithm 5 Generating weighted line graph $L$ from query $Q$.
-Observation 46.
-Algorithm˜5 generates the weighted line graph of $Q$ in time $\mathcal{O}(|Q|)$.
-Proof.
-Array initializations for $L$ and $M$ take $\mathcal{O}(m)$ time where $m\leq|Q|$.
-The first for–loop scans all predicates in $Q$ and inserts the corresponding edges in the multigraph $M$, performing $\mathcal{O}(|Q|)$ insertions, so $|M|=\mathcal{O}(|Q|)$.
-The auxiliary array weight has size $m$ and is initialized in $\mathcal{O}(m)$ time.
-In the subsequent traversal of $M$, each adjacency list is visited once.
-For every neighbor $j$ of a vertex $i$, the first inner loop performs a constant–time update of weight[j]; the second inner loop performs at most one constant–time insertion of $(j,\texttt{weight[j]})$ into $L$ and resets weight[j].
-Thus each stored neighbor (i.e., each edge occurrence in $M$) incurs only $\mathcal{O}(1)$ work overall.
-Since the total number of stored neighbors is $\Theta(|M|)=\mathcal{O}(|Q|)$, the time for these loops is $\mathcal{O}(|Q|)$.
-Summing up, all steps are linear in $|Q|$, hence Algorithm˜5 runs in time $\mathcal{O}(|Q|)$. ∎
-Appendix D Missing Proofs in Section˜4
-Section˜4.2 If $H^{\prime}\twoheadrightarrow H$ then $\mathcal{T}(H)\subseteq\mathcal{T}(H^{\prime})$.
-Proof.
-Let $T$ be a join tree of $H$, and $x^{\prime}$ be any vertex in $H^{\prime}$.
-Consider the neighborhood $H^{\prime}|_{x^{\prime}}$ of $x^{\prime}$ in $H^{\prime}$ comprising the set of hyperedges $\left\{r^{\prime}\in H^{\prime}\mid x^{\prime}\in r^{\prime}\right\}$.
-Since $H^{\prime}\twoheadrightarrow H$, we have that $x^{\prime}\in r^{\prime}$ if and only if $\mathsf{f}(x^{\prime})\in\mathsf{f}(r^{\prime})=r^{\prime}$.
-The neighborhood $H|_{\mathsf{f}(x^{\prime})}$ of $\mathsf{f}(x^{\prime})$ in $H$ comprises of the set of hyperedges $\left\{r\in H\mid\mathsf{f}(x^{\prime})\in r\right\}$.
-Therefore, $r^{\prime}\in H^{\prime}|_{x^{\prime}}\iff\mathsf{f}(r^{\prime})=r^{\prime}\in H|_{\mathsf{f}(x^{\prime})}$.
-Because $T$ is a join tree of $H$, $H|_{\mathsf{f}(x^{\prime})}$ is connected in $T$, therefore $H^{\prime}|_{x^{\prime}}$ is also connected in $T$ and $T$ is also a join tree of $H^{\prime}$.
-∎
-Section˜4.2 Let $T$ be an MCS tree. For any edge $e\in T$ that has a parent, then $e\not\subseteq p(e)$.
-Proof.
-In Algorithm˜1, a child $r_{c}$ is only connected to a parent $r$ if they share some previously unmarked vertex $x$.
-Because every vertex is marked only once, and the parent $r_{p}$ of $r$ was labeled before $r$, $x$ cannot be marked by $r$’s parent $r_{p}$.
-Let the $e$ be edge between $r$ and $r_{c}$, then the edge between $r$ and $r_{p}$ is $p(e)$, and $x\in e$ but $x\notin p(e)$, therefore $e\not\subseteq p(e)$.
-∎
-Section˜4.2 Let $T$ be an MCS tree. For two edges $e,e^{\prime}\in T$ that have parents,
-$(e\setminus p(e))\ \cap\ (e^{\prime}\setminus p(e^{\prime}))\ \neq\ \emptyset\quad\implies\quad p(e)=p(e^{\prime}).$
-Proof.
-We prove by contrapositive, assuming $p(e)\neq p(e^{\prime})$.
-Case 1:
-$p(e^{\prime})=e$, then $(e\setminus p(e))\cap(e^{\prime}\setminus e)\subseteq(e\cap e^{\prime})\setminus e=\emptyset$.
-Case 2:
-$p(e^{\prime})\neq e$.
-Without loss of generality, we suppose $e$ is no further from the root than $e^{\prime}$.
-Let $p(e^{\prime})=e^{\prime\prime}$, then $e^{\prime\prime}$ is on the path between $e$ and $e^{\prime}$ in $T$.
-If $(e\setminus p(e))\cap(e^{\prime}\setminus e^{\prime\prime})\neq\emptyset$, then $\exists\ x\in(e^{\prime}\cap e)\setminus e^{\prime\prime}$, which violates the running intersection property of join trees.
-Section˜4.2 Let $H^{\prime}$ be a duplicated hypergraph of $H$. Then,
-  1. 1.
-    $H^{\prime}\twoheadrightarrow H$ and $H^{\prime}\twoheadleftarrow H$
-  2. 2.
-    $\mathcal{T}(H^{\prime})=\mathcal{T}(H)$
-  3. 3.
-    $L(H^{\prime})=L(H)$
-Proof.
-1.
-We only need to show that $H^{\prime}\twoheadrightarrow H$ ($H^{\prime}\twoheadleftarrow H$) holds for $H^{\prime}$ derived from $H$ by a single vertex duplication.
-$H^{\prime}\twoheadrightarrow H$ ($H^{\prime}\twoheadleftarrow H$) will then hold for any duplicate $H^{\prime}$ derived by a series of vertex duplications since strong homomorphisms are closed under composition.
-Suppose $H^{\prime}$ is derived from $H$ by duplicating $x\in X(H)$ to $x^{\prime}$.
-Consider a pair of functions $(\mathsf{f}_{X}:X(H^{\prime})\mapsto X(H),R(H)_{\texttt{id}})$ where $\mathsf{f}_{X}$ maps $x^{\prime}$ to $x$.
-Otherwise, it is an identity map $X(H)_{\texttt{id}}$.
-This function pair constitutes a strong hypergraph homomorphism $H^{\prime}\twoheadrightarrow H$.
-Likewise, the pair of functions $(X(H)_{\texttt{id}},R(H)_{\texttt{id}})$ constitutes a homomorphism $H\twoheadrightarrow H^{\prime}$.
-2. It follows from Section˜4.2 and the facts above.
-3.
-By the strong homomorphisms, the hyperedge sets have the same cardinality $|R(H)|=|R(H^{\prime})|$.
-So do the vertex sets of the line graphs, $R(L(H))=R(H)$ and $R(L(H^{\prime}))=R(H^{\prime})$.
-By construction, the vertex duplication adds $x^{\prime}$ to each hyperedge containing $x$, therefore does not create any additional edges in the line graph, $E(L(H^{\prime}))=E(L(H))$.
-∎
-Appendix E Missing Proofs in Section˜5
-Section˜5 An $\alpha$-acyclic hypergraph is Berge-acyclic if and only if it is linear.
-Proof.
-Fagin [13] showed that every Berge-acyclic hypergraph is linear.
-We prove the other direction by contradiction, by assuming that $H$ is $\alpha$-acyclic and linear.
-We also assume that there is a Berge cycle $(r_{0},x_{0},\dots,r_{k-1},x_{k-1})$ of length $k\geq 2$.
-Without loss of generality, we suppose $r_{1}$ is the first hyperedge on the cycle to be removed by GYO reduction.
-Then $r_{1}$’s parent $r_{p}$ must satisfy $r_{p}\cap r_{1}\supseteq(r_{1}\cap\bigcup_{i\in[k]\wedge i\neq 1}r_{i})=\{x_{0},x_{1}\}$, contradicting $H$’s linearity.
-∎
-Section˜5.1 The line graph $L$ of a Berge-acyclic hypergraph $H$ is a block graph.
-Proof.
-A Berge-acyclic hypergraph is $\gamma$-acyclic [13].
-Zhu proves that the line graph of a $\gamma$-acyclic hypergraph is chordal [48].
-We now prove that the line graph is diamond-free by showing that every diamond must be part of a $K_{4}$.
-Assuming a diamond $(r_{1},r_{2},r_{3},r_{4})$ in $L$ without the edge $(r_{1},r_{3})$, it can be treated as two triangles $(r_{1},r_{2},r_{4})$ and $(r_{2},r_{4},r_{3})$ sharing an edge $(r_{2},r_{4})$.
-We now consider the possible edge labelings in each triangle:
-Case 1:
-If all edges are labeled with distinct variables, each triangle forms a Berge-cycle, contradicting the Berge-acyclicity.
-Case 2:
-If two edges are labeled with the same variable, all three vertices share that variable. The third edge must be labeled with the same variable.
-Since the two triangles share the edge $(r_{2},r_{4})$, all their edges are labeled with the same variable $x$, namely $x\in r_{1}\cap r_{2}\cap r_{3}\cap r_{4}$.
-There has to be an edge $(r_{1},r_{3})\in L$, contradicting the assumption that $(r_{1},r_{2},r_{3},r_{4})$ is a diamond.
-∎
-Section˜5.2 Let $T_{r}$ be the canonical join tree rooted at $r$ of a Berge-acyclic hypergraph.
-Along its root-to-leaf path $r,r_{1},\ldots,r_{k}$, each pair of adjacent vertices shares a distinct variable.
-Proof.
-Let $r_{0}=r$, Section˜5 requires that $|r_{i-1}\cap r_{i}|=1$ for all $i\in[k]$.
-By Theorem˜33, the path $r_{0},\ldots,r_{k}$ is the shortest path from $r_{0}$ to $r_{k}$.
-The variable shared by two adjacent vertices $r_{i-1},r_{i}$ is equivalent to the edge label $\chi(r_{i-1},r_{i})$.
-The edge labels along the path are all distinct.
-Otherwise, two consecutive edges with the same label allows for shortening the path by removing either, contradicting the shortest path property.
-Two non-consecutive edges with the same label violate the running intersection property.
-∎
-Appendix F Additional Figures, Tables and Algorithms
-Name # Queries # $\alpha$-Acyclic # Composite-Key Joins # Berge-Acyclic
-TPC-H[41] 22 21 2 19
-JOB[24] 113 113 0 113
-STATS-CEB[17] 2603 2603 0 2603
-CE[8] 3004 1839 0 1839
-Spider-NLP[46] 4712 4709 6 4703
-Table 1: Acyclic queries in the benchmarks (all $\alpha$-acyclic queries are also $\gamma$-acyclic.)
-Input: Left-deep plan $r_{1},r_{2},\dots,r_{n}$
-Output: Tree rooted at $r_{1}$
-1
-[…]
-3 for $i\leftarrow 2$ to $n$ do
-4   key $\leftarrow r_{i}\cap\bigcup r_{k<i}$
-5
-6 for $j\leftarrow 1$ to i1i-1 do
-7    if $\text{key}\subseteq r_{j}$ then
-8       $p(r_{i})\leftarrow r_{j}$
-9       break
-10
-11      end if
-12
-13   end for
-14
-15 end for
-Algorithm 6 Converting a binary plan to a join tree
-Figure 5: Possible cases of how $r_{k+1}$ is connected to $G_{k}$
-(a) $H_{3}$
-(b) $T_{S}(H_{3})$
-(c) $H_{5}$
-(d) $T_{W}(H_{5})$
-(e) $T^{\prime}_{W}(H_{5})$
-Figure 6: $H_{3}$ is a non-Berge hypergraph that admits a unique canonical join tree $T_{S}(H_{3})$ with any relation chosen as root, such as $T_{S}(H_{3})$.
-$H_{5}$ is a $\gamma$ hypergraph that does not admit unique canonical join trees at any relation chosen as root.
-For example, $T_{W}(H_{5})$ and $T^{\prime}_{W}(H_{5})$ are MCS trees generated by Algorithm˜1.
-Neither of them is a canonical join tree rooted at $W$.
-Input: Hypergraph $H$, root $r$
-Output: weighted join tree $T_{r}$ and unweighted line graph $L$
-1
-[…]
-[…]
-4while $R\neq\emptyset$ do
-5 $i\leftarrow i+1;r_{i}\leftarrow r$
-6 for $x\in r_{i}\cap X$ do
-7    for distinct $r^{\prime},r^{\prime\prime}\in R:x\in r^{\prime}\cap r^{\prime\prime}$ do
-8       $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}E(L)\leftarrow E(L)\cup\{(r^{\prime},r^{\prime\prime})\}$
-9       $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\omega(r^{\prime},r_{i})\leftarrow|r^{\prime}\cap r_{i}|$
-10       $p(r^{\prime})\leftarrow r_{i}$
-11
-12
-13 $X\leftarrow X\setminus r_{i};R\leftarrow R\setminus\{r_{i}\}$
-14 $r\leftarrow\operatorname*{arg\,max}_{r\in R}|r\setminus X|$
-15
-return $L,p(\cdot),\omega(\cdot)$
-Algorithm 7 MCS for the $\gamma$-acyclic
-Input: weighted MCS tree $T_{r}(H)$ weight function $\omega$ and unweighted union join graph $L(H)$
-1 $G^{\equiv}\leftarrow L(H)$
-2for $e^{*}=(r_{i},r_{j})\in R(L(H))\setminus R(T_{r}(H))$ do
- // where $0pt(T_{r}(H),r_{i})\leq 0pt(T_{r}(H),r_{j})$
-3
-4 $l\leftarrow\mathsf{LCA}(r_{i},r_{j});d\leftarrow 0pt(l)+1$
-5 $e_{1}=(l,r_{1}\leftarrow\mathsf{LA}(r_{i},d));w_{1}=\omega(e_{1})$
-6 $e_{2}=(l,r_{2}\leftarrow\mathsf{LA}(r_{j},d));w_{2}=\omega(e_{2})$
-7 if $e_{1}=e_{2}$ then
-8      slide $r_{j}$ to $r_{1}$
-9
-10 else if $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}w_{1}=w_{2}$ then
-11      slide $r_{i}$ to $r_{1}$ and $r_{j}$ to $r_{2}$
-12
-13 else if $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}w_{1}<w_{2}$ then
-14      slide $r_{i}$ to $l$ and $r_{j}$ to $r_{1}$
-15
-16 else if $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}w_{2}<w_{1}$ then
-17      slide $r_{i}$ to $l$ and $r_{j}$ to $r_{2}$
-18
-19return $G^{\equiv}$
-Algorithm 8 buildEGbuildEG for the $\gamma$-acyclic
+Appendix F Notation Table
+Variable Definition
+$H,X,R$ hypergraph, vertex set, hyperedge set
+$\chi:R\to 2^{X}$ Incidence function mapping each hyperedge to its vertices.
+$X(H),R(H)$ Vertex set and edge set of $H$.
+$|H|$ Size of hypergraph: $|H|=\sum_{r\in R}|r|$.
+$H|_{x}$ Neighborhood of $x$ in $H$: set of hyperedges containing $x$.
+$H^{*},X^{*},\chi^{*}$ Equivalent hypergraph, its vertices, and incidence function
+$\mathsf{f}$ A hypergraph homomorphism (pair $(\mathsf{f}_{X},\mathsf{f}_{R})$).
+$H_{1}\rightarrow H_{2},H_{1}\twoheadrightarrow H_{2}$ Homomorphism from $H_{1}$ to $H_{2}$, strong homomorphism from $H_{1}$ to $H_{2}$.
+$G,R,E$ (Multi-)graph, nodes (corresponding to hyperedges), edges
+$\rho:E\to 2^{X},\omega:E\to\mathbb{N}$ Incidence function mapping each edge to its endpoints, weight function
+$R(G),E(G)$ Vertex set and edge set of $G$.
+$G^{\equiv},\rho^{\equiv},T^{\equiv}$ Equivalent graph (Def. 4.1), its incident function, and a spanning tree
+$\leq_{T}G^{\prime} GG$ can be slid into $G^{\prime}$ (Def. 4.1).
+$L(H),L,\chi(e),\omega(e)$ Line graph of $H$, edge label $\chi(\{r_{1},r_{2}\})=r_{1}\cap r_{2}$, weight $\omega(e)=|\chi(e)|$.
+$|L|$ Size of line graph: $|L|=\sum_{e\in E(L)}\omega(e)$.
+$G|_{x}$ Subgraph of $G\subseteq L(H)$ induced by $H|_{x}$.
+$T^{G},T^{M}$ Join tree generated by GYO and MCS, resp.
+$T,T_{r}$ Unrooted and rooted (at $r$) join tree.
+$\hat{T}$, $\hat{T}_{r}$ A monotonic weight join tree (Def. 4.2), one rooted at $r$.
+$\mathcal{T}(L(H))$, $\mathcal{T}(H)$ Set of (unrooted) join trees of $L(H)$ / $H$.
+$0pt(T_{r},r_{i})$ Depth of node $r_{i}$ in rooted tree $T_{r}$.
+$r_{p}$ The parent of $r_{i}$ in a GYO reduction order.
+$c(r),p(e),s(e)$ Children nodes of $r$, parent edge of $e$, sibling edges of $e$
+$\mathsf{LCA}(r_{i},r_{j}),\mathsf{LA}(\cdot)$ Lowest common ancestor and level-ancestor.
+$\lambda(e)$ LCA edges of a non-tree edge $e=(r_{i},r_{j})$.
+$\Delta$ Number of duplications needed to enforce monotonicity.
+Table 2: Notation introduced in sections 34.